A Complete Bibliography of Publications in Synthese, 1990–1999

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

 E-mail: beebe@math.utah.edu, beebe@acm.org, beebe@computer.org (Internet)
            WWW URL: https://www.math.utah.edu/~beebe/

08 November 2023 Version 1.04

Title word cross-reference

#11 [692]. #12 [697]. #13 [703]. #19 [756]. #3 [632]. #4 [646]. #5 [652]. #6 [660]. #7 [669]. #8 [674]. #9 [680].

S_{5} [555].

14 [77]. 15 [142]. 16 [206]. 17 [517]. 1990s [147]. 1994 [359].

Abnormic [790]. abstract [551]. acceptability [283]. Acceptance [719]. accessible [62]. Accidental [218]. Account [682, 394, 128]. Accounting [408, 373]. accuracy [55, 160, 362]. achromatopsia [491]. acquired [141]. Acquisition [591]. Action [584, 586, 76, 100, 494, 187, 99, 385]. Action-Guiding [584, 586]. actions [76, 457]. Actual [649]. Actualism [87]. Adam [308]. Adams [415]. adaptation [229]. addiction [339]. additive [146]. affair [133]. affect [492]. after [245]. Against [90, 684, 708, 195, 317, 46, 233, 551]. age [339]. Agency [789, 546, 111]. agents [738]. Agnostic [746]. ago [42]. AI [551]. aim [534, 46]. ain’t [184]. Akinetopsia [491]. akrasia [307]. Alan [549, 811]. Algorithmic [747]. Algorithms [799]. alleged [94]. Allport [267]. allure [419]. alternative [390]. Ambiguities [388]. Ambiguity [669, 88, 461]. Amsterdam [359]. analogies [469]. analogy [337]. Analysis [668, 636, 233]. analytical [486, 384]. ancient [60]. Announcement [9, 63, 98, 116, 158, 214, 246, 251, 269, 389, 564, 698]. Announcements [86, 123, 439, 472]. answer [233, 385]. Antarctic [399]. anthropology [333]. Anti [505, 557, 708, 676, 56, 115]. Anti-Realism [708, 676, 505, 56, 115]. Anti-realist [557]. anticipation [549, 811]. Anyway [642]. aphasics [457]. Apollo [401]. applications [365, 432]. Applied [727, 232]. Appraisal [677, 365]. appreciation [255]. Approach [647, 248, 333]. Approaches [702]. Approximate [605]. Approximation [745, 247]. Apriority [232]. Arabs [34]. Arbitrary [631, 787]. Archimedean [574]. architecture [393]. Argument [645, 689, 684, 683, 722, 623, 422, 404, 471, 527, 551]. Argumentation [359]. arguments [84, 291, 541, 425, 553]. Aristotle [218, 642, 319, 316, 315, 318, 641, 303]. Arithmetic [629, 693, 435]. Arithmetization [571]. arrows [528]. Art [278]. article [57]. artifacts [146]. Artificial [157]. aspects [175, 324, 379]. ’Aspekt [679]. assessment [459]. Assistants [760]. Association [803]. Assumption [632, 669]. Assumptions [655]. Asymptotic [446]. atomism [125]. atomistic [394, 171]. attitudes [350]. attribution [414, 134]. Aufbau [236, 238]. Aureoli [304]. authority [460]. automaton [547]. avoid [542]. awareness [738, 491]. Axiomatic [372]. axioms [569].

B [527]. B-Theory [527]. Back [621, 145]. Background [628, 320]. Backward [210]. Backwards [622, 72]. Bare [648]. base [209]. Based [731]. basis [153, 543]. Bateson [201]. be [17, 579, 154, 256, 183]. Beauty [600]. Became [763]. becoming [452]. bee [490]. Begging [387]. Behaviorism [390]. Being [618]. Belief [176, 723, 208, 559, 750, 134, 179, 82, 353, 414, 461, 180, 275]. belief-formation [180]. beliefs [210, 211]. believe [274]. Believing [178, 744]. Bell [671, 426, 425]. Bertrand [581]. best [352]. Between [716, 163, 552, 446, 325, 161]. Beyond [153, 12]. bias [259]. Bibliography [608]. Big [127]. Bildtheorie [762]. Bill [469]. Biological [224, 529]. biology [363, 364]. biomedicine [164]. blindsight [491]. blue [220]. Body [730, 498]. Bohm [671, 713, 651]. Bohmian [523]. Boltzmann [764, 765, 759, 760, 761, 766, 763, 762]. Bonjour [613, 643]. book [46]. Boole [634]. boundaries [224]. boundless [242]. Bourbaki [231]. Branching [234, 620, 296]. Brandon [364]. Brenner [44]. Britain [640, 639]. Bulletin [206, 517, 77, 142]. Burge [742, 436]. Buridan [307]. Butts [609].

Calculus [627]. Call [23, 68, 518]. called [133]. calls [472]. Can [646, 160, 17, 647, 139, 542]. cannot [154]. Canonical [644]. Cantor [637]. capability [397]. cardiology [164]. Carnap [703, 485, 240, 443, 241, 243, 237, 242, 235, 244, 245, 239]. Cartesian [29, 320, 181, 623]. Cartesianism [500]. Cartwright [165]. Carving [682]. Case [580, 604, 583, 288, 339, 164, 399, 323, 28]. Cassirer [140]. Catastrophe [310]. categories [207]. Category [454]. Causal [798, 799, 681, 726, 781, 651, 804, 805, 103, 336, 533, 376, 185, 295]. Causation [622, 689, 706, 802, 803, 797, 620, 71, 408, 72, 108, 502]. Cause [250]. Causes [620, 4, 218, 265]. Center [623]. Center-of-Mass [623]. Century [640, 639, 42, 38, 138]. Certain [702]. certainty [361]. Ceteris [754]. Chains [495]. challenge [416]. Challenged [753]. chance [377]. Change [677, 750, 164, 559, 13, 305, 141]. Changes [723]. changing [302, 349]. Chaos [169, 649]. Character [665]. characteristics [141]. Chauvinism [776]. cheap [502]. Chemistry [607, 608, 604, 606, 427]. Chinese [287, 288]. Choice [268, 224, 562]. Chomsky [784]. Christopher [39]. Circle [115]. cladism [205]. claims [533]. classes [196]. Classical [286, 580, 386, 393, 432]. classification [48, 39]. Clavius [39]. Clinton [469]. Close [294]. Closure [667, 112]. cogito [29, 501]. Cognition [645, 393, 736, 301]. Cognitive [673, 110, 459, 417, 83]. Cohen [370]. Coherence [684, 115, 560, 445, 537]. coherent [69]. Collective [719, 100]. color [506]. colors [391]. Colour [455]. come [66]. Comment [49, 284, 540]. commentaries [38]. Comments [364, 348, 343, 51, 44, 47]. Commercial [640, 639]. commitment [110, 339, 179, 87]. Common [715, 28, 4, 161]. common-sense [161]. commons [399]. Community [595]. comparative [302]. Compatibility [722]. Compensatory [400]. competence [370]. competitive [397]. Complementary [783]. Completenes [574]. Completeness [660, 574, 573]. Complexity [330]. Computation [645, 664, 672, 459, 548]. Computational [492, 529]. Computationalism [493]. Concept [749, 656, 810, 495, 552, 231, 418, 243]. conception [48]. Concepts [605, 13, 557, 61, 289]. Conceptual [203, 677, 196, 661, 14, 676, 324, 499, 290, 13]. concerning [141]. Conditio [140]. Condition [587, 798]. Conditional [306, 539, 268, 566]. Conditionalization [684, 90]. Conditionalizing [753]. Conditionals [701, 409, 208]. Conditions [788, 792, 592]. Conference [359]. Confirmation [556]. Confirming [746]. Confusion [205, 540]. Conjecture [602]. Conjunctions [744]. Connectionism [10, 422, 549, 811, 419]. Connectionist [459, 552, 421, 416, 286]. connexions [3]. conscious [325]. Consciousness [789, 187, 342]. Consequence [689]. conservation [35]. Consistency [188, 558, 62]. Consistent [808]. Constant [391]. constants [386]. Constrained [413]. constraints [376]. Constructing [623]. construction [54, 279, 237]. Constructive [411]. constructivism [167]. contemporaries [304]. contemporary [495]. Content [664, 392, 113, 5, 181, 455, 18]. context [357, 276]. contexts [44]. contingency [392]. contingents [304]. Continua [650]. continuous [451]. contraction [209]. contradiction [138]. contrast [533]. control [392].

Convention [678, 725, 52, 531]. conventionalism [240, 46]. conventions [53]. cooperation [413, 399]. Coordinate [751]. coordination [52]. Copernicus [41, 40]. Cordoning [370]. Correction [811]. correlations [6]. correspondence [169]. cosmology [34]. Count [696]. Counterfactual [802]. Counterfactuals [769, 388]. Counting [631]. Covariation [392]. covering [105]. Creation [685]. crisis [43, 501]. Criteria [631, 716, 488]. Critical [668, 636, 449, 255, 308, 313]. criticism [500, 357]. critics [310, 285]. Critique [581, 441, 413, 170, 354]. Crowe [51]. Cultural [202, 146]. Curie [578]. curious [288]. curriculum [258].

Darwinians [201]. Darwinism [201]. data [552]. data-oriented [552]. Davidson [85, 741, 794, 307]. Debate [779, 239]. deception [414]. deceptive [180]. Decidability [747]. Decision [701, 752, 569, 30, 400]. decision-making [400]. Decision-theoretic [30]. Decoherence [424, 647, 519, 431]. Dedekind [60]. Deduction [535, 440, 636, 421, 320]. Deductive [776, 112]. Default [585]. Defeater [587]. Defence [770]. Defense [622, 694, 693, 296, 209]. defies [105]. definability [452]. definitions [88]. deflationary [459]. Deliberative [537]. democracy [255]. Demonstrative [712, 362, 249]. Demonstratives [662]. denoting [437]. deontic [75]. dependence [55, 568]. dependency [14]. Dependent [749]. Descartes [498, 221, 35]. Description [678, 286]. design [404, 471]. Despite [598]. Determinables [778]. Determination [615, 323]. Determinism [521, 172]. Deterministic [649, 804, 174]. deterrence [398]. Development [700, 614, 203, 126]. Deviancy [681]. Devices [809]. Dewey [484]. dialectic [315]. dialogue [381]. dictum [316]. Did [637, 192]. different [44]. difficulties [94]. Dilemma [100, 373, 413]. Dilemmas [102, 139]. dimensional [456]. direct [71]. discipline [224]. Discontinuity [615, 323]. discourse [384]. Discover [799]. discrete [451]. discrimination [456]. Discussion [652]. Discussions [151]. disease [163]. disjunctive [139]. disputations [303]. dispute [331]. Distinction [632, 696]. distinguished [692]. diversification [562]. Diversity [798]. divine [35]. Division [206, 517, 77, 142, 631]. DLMPS [206]. do [356, 66, 755]. doctrine [301]. Does [547, 441]. Dogma [791]. Doing [420, 356, 642]. domains [62]. dominance [99]. Domingo [36]. Donald [307]. don’t [265]. double [475]. draft [149]. Dretske [113]. Dreyfus [343]. Duhem [42, 47, 50, 40, 45, 48, 46, 43, 49, 165, 44, 34, 36]. Duhemian [2, 37]. Dummett [197]. Dumpty [489]. dynamic [552]. Dynamical [771]. Dynamics [772, 623]. Dysfunctional [704].

each [692]. Early [634, 642, 580, 140, 239]. earth [227]. Ecological [216]. economical [266]. edition [308]. editor [235]. Editorial [101, 511, 198, 215, 464, 64, 610, 653, 434, 327, 603, 360]. Editors [314]. Edmund [637]. education [255, 257]. effective [428]. effects [265, 397]. Einstein [47, 45, 430]. either [184]. elaboration [42, 413]. electrostatic [362]. Element [674]. elementary [311]. Elements [131]. elenctic [219]. Elgin [284]. Eliminativism [461]. Elster [343]. elucidation [239]. embryology [202]. Emerging [544]. emotions [340]. emphasis [295]. empirical [451, 331, 114]. empirically [447, 15]. Empiricism [695, 327, 479, 411]. enemy [396]. Energetics [759]. Energy [93]. engineering [202]. English [379]. enough [267, 294]. entanglement [170]. entia [302, 349]. entirely [245]. entrenchment [559]. Entropy [715]. environment [326]. Epistemic [674, 81, 693, 584, 586, 271, 667, 80, 561, 29, 272, 208, 559, 280, 558]. epistemically [82]. epistemological [233, 156, 186]. epistemologies [475, 140]. Epistemology [697, 764, 236, 743, 611, 735, 694, 450, 79, 21, 284, 354, 257, 30]. Epoché [733]. equant [41]. Equations [805]. equilibrium [368]. equivalent [447]. era [12]. Ergodic [543]. Ernest [415]. Erotetic [790, 553]. Erratum [59, 740]. Essay [691, 212, 523, 504, 97, 570, 463, 516, 96, 309, 213, 407]. essays [313]. essence [250]. Essential [598]. established [322]. Ether [774]. Eudoxos [60]. evaluation [84]. Events [620, 104]. Everett [173]. every [547]. Evidence [615, 611, 445, 362, 323, 114]. Evolution [22, 597, 583, 203, 199, 152, 455]. Evolutionary [223, 363, 364, 513, 81, 168]. Ex [685]. Exam [690]. examination [32]. excavations [454]. Excitation [771]. excluded [304]. exclusion [427]. exclusively [12]. Existence [808]. existential [380]. exists [230]. Exner [761]. expected [194, 268]. Experience [648, 612, 658, 484, 345]. Experiment [742, 363, 364, 365]. experimental [224, 93]. experiments [367, 508, 366]. explain [265]. Explained [777]. Explanation [678, 773, 25, 777, 775, 418, 673, 681, 605, 780, 781, 790, 448, 352, 529, 334]. Explanations [616, 783, 299]. Explanatory [778, 781, 533]. explication [447]. exportation [133]. Expositions [635]. Expressions [710]. extended [398]. Extending [541]. Extension [699]. extensionality [346]. Externalism [587, 584, 586, 460, 156]. externality [181]. extrinsic [545, 568]. extrinsicness [181].

faces [270, 168]. factors [202]. failure [312]. fair [402]. Fall [765]. Fallacies [641, 468]. Fallacy [734, 387]. Falsification [2]. Family [488]. fancy [190]. Fast [456]. Feeling [492, 438]. Feels [579]. feminist [475, 477, 479]. fictions [297]. Field [423, 428]. fill [441]. Finite [453, 547]. finite-state [547]. finitism [487, 197]. First [727, 774, 460, 469]. First-Order [727]. first-person [460]. flesh [404, 471]. focused [538]. Folk [134]. Following [652, 489, 128]. foreign [400]. Form [700, 18, 318, 456, 319, 28]. Formal [710, 702, 350, 432, 429, 372]. Formalism [571]. Formalization [576, 483]. formation [495, 552, 180]. forms [293]. Formulating [496]. Formulations [713, 424]. foundation [186]. foundationalism [275]. Foundations [660, 122, 599, 676, 441, 324, 569, 241, 454]. fourteenth [38]. Fraassen [353, 354]. Fragments [808]. Frame [623, 19]. framework [417, 429].

Framing [19]. Freedom [689]. Frege [84, 632, 669, 809, 482, 631, 435, 436]. frequentism [213]. Friedman [105]. friend [396]. front [145]. function [524]. functional [28]. Functionalism [11, 337]. functionalist [154]. Functionality [705]. Functions [593]. Fundamental [695]. Further [299]. future [304]. fuzzy [92]. fuzzy-set [92].

Gadflies [758]. Game [380, 690, 470, 217]. Game-theoretical [380]. games [377, 381]. gap [441]. Gas [758]. Gauthier [413]. Geertz [333]. Gender [476]. General [646, 108, 410, 504]. generality [88, 435]. generalization [194]. generalizations [329]. generalized [336]. Generation [326]. Geniuses [758]. Geometry [700, 435, 576]. Georg [637]. Gestalt [326]. Getting [190, 375]. Giles [305]. Give [690]. Gödel [654, 655, 659, 656, 151, 430, 299]. Good [478, 502]. Goodman [514, 284, 280]. governance [212]. Grades [673]. Graham [299]. graphs [380, 166]. graphs- [166]. gravitatione [623]. Gravity [615]. greatest [230]. Greek [60]. Grene [223]. Group [211, 428, 494]. Grünbaum [2]. Grundgedanke [795]. Grundgesetze [808]. Gts [378]. Guiding [584, 586]. gym [288].

Haack [612]. Hands [264]. Hans [202]. Hansen [642]. harmony [322, 386]. Hath [654]. head [184, 391]. heat [366]. Hegemon [640, 639]. hegemonic [402]. Heidegger [347]. Helm [759]. Help [647]. hemisphere [457]. Hempel [105]. Hermeneutical [687]. herring [220]. Heyting [490]. hidden [522, 427]. hierarchically [204]. hierarchies [27]. High [93]. High-Energy [93]. higher [89, 196]. higher-order [89, 196]. Hilbert [575, 403, 572, 573, 571, 576]. Hintikka [348, 641]. historical [241]. histories [201]. Historiographical [47, 37]. History [206, 517, 758, 607, 77, 142, 700, 50, 33, 504, 51, 222, 34, 313]. Holding [225]. Holism [42, 556, 331, 125]. home [438]. Honderich [343]. hood [337]. Howard [47]. Howell [636]. Human [710, 26, 550, 131, 27]. Hume [723, 688, 220]. Humphrey [539]. Humpty [489]. Hunting [800]. Hussearle [628]. Husserl [84, 627, 677, 582, 637, 733, 573, 347]. Huygens [623]. Hypercomparatives [594]. Hypotheses [790]. Hypothesis [774, 628].

iconicity [514]. Idea [590]. Ideal [616, 83, 560]. Idealisation [371]. idealism [69, 513, 237, 238]. idealist [442]. Idealization [727, 745]. Ideas [702]. identical [6]. Identification [662, 787, 802, 680]. Identity [666, 291, 157, 71]. II [16, 149]. III [24]. image [510]. imaginary [544]. Impact [649]. Impart [646]. impatient [124]. imperfective [135]. implement [547]. implementation [418]. implies [139]. importance [504]. impossible [178]. inaccuracy [160]. Incidence [576]. Incompatible [447]. Incompleteness [707]. Indeterminacy [114, 485, 291, 88, 155, 506]. Indeterminism [761]. indeterministic [426]. Indexical [665]. Indication [229]. Individualism [489, 331]. individualism-holism [331]. Individuality [779, 429]. Individuation [664, 716, 104]. Indoctrination [595]. Inducing [402]. Induction [712, 731, 73, 538, 210, 535, 362, 335, 280, 106]. Inductive [688, 585, 541]. Inequalities [671]. Inference [448, 699, 457, 352, 228, 386, 320]. Infinity [685]. Influence [637, 148]. information [530, 567, 20, 568, 481]. Informational [382]. inheritance [442, 141]. Initial [792]. innate [289]. Inner [462]. innovation [328]. Inquiry [611]. instant [305]. intangible [264]. integration [92]. Integrity [696]. intelligence [157]. Intension [699]. intent [397]. Intention [110]. Intentional [739, 738]. Intentionality [737, 806]. Intentions [76]. Interaction [163]. Interdisciplinary [79]. Interdiscourse [510]. internal [403]. Internalism [587, 82]. Internalism* [694]. Internalistic [275]. internation [396]. International [359, 399, 397, 206, 517, 77, 142]. Interpretation [774, 713, 733, 710, 651, 810, 805, 94, 67, 539, 173, 433]. Interpretations [802, 322]. Interpreted [293]. Interpreting [385]. interpretive [333]. interrelations [29]. interrogative [378]. intertheoretic [446]. intervention [1]. Intrinsic [545]. Intrinsic/ [545]. Introduction [31, 757, 416, 395, 143, 603, 276, 797, 198, 215, 327, 314, 235]. intuitionism [197]. Intuitionistic [674]. intuitions [61]. intuitive [301]. invariant [153]. inverse [362, 539]. inverted [11]. investigation [372]. investigations [355]. invisible [264]. Irrationals [747]. irreducibility [109]. Irreversibility [645]. Isolation [631]. Israel [261, 255]. ISSA [359]. Issue [797]. Issues [607, 203]. IUHPS [206]. IUHPS/DLMPS [206].

John [307]. Joint [494]. Joints [682]. June [359]. Just [782]. Justification [589, 612, 651, 70, 275, 80]. justifications [372].

Kant [634, 61, 139, 199, 70, 440, 633, 375, 430, 441, 636, 810]. Kantian [238]. Kierkegaard [339]. Kim [741]. kinematics [152]. Klemens [78]. Knower [667, 273]. Knowers [216]. Knowledge [743, 588, 156, 20, 726, 112, 221, 505, 262, 113, 284, 388, 351]. known [216]. Kornblith [791]. Koyré [36]. Kripke [465]. Krister [76].

Labor [484]. Lady [469]. Lamp [704]. Language [627, 25, 591, 443, 590, 728, 725, 700, 785, 784, 786, 243, 716, 242, 726, 702, 55, 514, 14, 484, 438, 111]. Late [583]. Lauraudogoitia [772]. law [105, 482, 362, 538]. Laws [773, 741, 745, 794, 614, 266, 412, 103, 330, 165, 263]. Learning [751, 456]. Lectura [308]. Lectures [765]. Legitimate [580]. Leibniz [699, 137, 322]. Length [657]. lesions [457]. Lessons [801, 371]. level [408]. Levels [286, 10]. Liar [728]. Liberal [595]. light [462, 366]. Like [490, 579, 183]. limiting [446]. limits [329, 328]. Linguistic [809, 763]. linguistically [153]. Linguistics [785]. Living [204]. locality [521]. Logic [727, 206, 517, 669, 674, 381, 77, 142, 642, 723, 688, 242, 585, 790, 508, 196, 403, 554, 29, 241, 75, 567, 56, 553, 188]. Logical [591, 443, 32, 729, 243, 237, 242, 808, 702, 383, 293, 120, 386, 182, 125, 213]. logicians [195]. loveliest [448]. Lübeck [759]. Luders [5]. Ludwig [762].

Mâché [694]. Mach [761]. Machine [710, 550]. Machines [685]. macro [408]. macro-level [408]. mad [129]. made [494]. Magnificent [775]. Main [636]. Maiocchi [47]. Making [146, 469, 437, 400]. man [124]. manifest [510]. Manifestation [683]. manipulability [541]. Manuals [638]. many [173]. Marcus [465, 467]. Margenau [170]. Marginal [803]. Marjorie [223]. Markov [798]. Martin [44]. Mass [696, 623]. Mass-Count [696]. materialism [292]. materiality [366]. Math [149]. mathematic [312]. Mathematical [522, 658, 400, 746, 600, 601, 397, 54, 231, 61, 297, 65]. Mathematics [577, 660, 122, 145, 599, 597, 598, 630, 766, 150, 656, 571, 50, 403, 121, 144, 241, 232, 51, 453, 147, 454, 148, 60, 313]. Matter [317, 458, 319]. matters [224, 458]. maximization [413]. Maximum [715]. McTaggart [525]. Meaning [590, 720, 345, 656, 495, 281, 386]. meaningful [146]. meanings [692]. meant [356]. Measurably [753]. Measurement [695, 581, 647, 651, 433, 95]. measurements [519, 170]. mechanicalness [188]. Mechanics [622, 801, 713, 670, 651, 94, 523, 432, 424, 171, 431, 520, 3, 433]. Mechanisms [778]. mediaeval [302, 349]. Medieval [583, 306, 320, 303]. Melancholic [21]. Memoriam [609]. memory [275]. Mental [499, 730, 502, 394]. Mentalese [643]. mentalism [390]. Merleau [737, 736]. Meta [625]. Meta-Ranking [625]. metalanguages [443]. metaphor [470]. metaphors [499]. Metaphysical [441, 635, 244]. Metaphysics [634, 412]. metatickles [99]. Meter [807]. Method [677]. methodological [248, 155, 114]. Methodology [206, 517, 77, 142, 708]. Metonymic [734]. middle [49, 257, 304, 125]. Mind [730, 498, 514, 550, 298, 173, 325, 188, 299]. mind-body [498]. minds [126]. Minimal [371, 750]. Minkowski [452]. misapplied [415]. modal [67, 555]. modal-structural [67]. Model [709, 495, 432, 105]. model-theoretic [432]. Modeling [396, 397]. modelling [413]. Models [577, 622, 803, 207, 804, 552, 286, 384, 400]. modern [162, 514, 223, 60]. Modernizing [144]. Modification [736]. modularity [492]. modules [27]. Molecular [605, 326]. monotonic [105]. monte [418]. Moore [130]. Moral [589, 721, 75, 139, 722, 227]. most [289]. motion [233, 35]. moves [377]. Much [796]. Multiple [672]. music [226]. my [396]. Myth [695, 149].

Naive [247]. Names [665, 666]. nativism [10]. Natural [421, 590, 290, 441, 583, 267, 66, 48]. Naturalism [791, 270, 371]. Naturalistic [735, 716]. naturalized [477]. Nature [646, 648, 745, 331, 155]. Naturforscherversammlung [759]. Necessary [3]. Necessity [807, 666, 320, 185]. Need [621]. negation [379]. Neurath [442]. neuroscience [29]. Newcomb [100, 99, 102]. Newton [774, 623].

Newtonian [668, 615, 792]. Neyman [192]. Nicolas [231]. night [489]. nihilism [339]. Nihilo [685]. Nineteenth [640, 639]. No [206, 517, 587, 77, 142, 754, 230]. No-Defeater [587]. Noemata [483]. Nomic [682, 193]. Non [616, 631, 99, 808, 174, 105, 429, 140]. non-action [99]. Non-Arbitrary [631]. non-deterministic [174]. Non-dominance [99]. Non-Ideal [616]. non-individuality [429]. Non-Logical [808]. non-monotonic [105]. nonclassical [417]. noncompensatory [400]. nonideal [170]. Noninvariance [804]. nonreductive [292]. nonstandard [233]. normative [212]. normativity [271]. Norms [584, 586, 81, 340, 306]. Note [92, 653, 101, 511, 287, 173]. Notes [765, 155, 563]. Notion [737, 498]. Notions [719, 561]. notorious [133]. Number [810]. Numbers [787, 66, 538].

Obituary [78]. Object [709, 457]. objections [170, 233]. objective [452]. Objectively [565]. Objectivity [675, 475, 257, 474]. Objects [808, 65]. obligation [546]. obligational [303]. Obligations [376, 139]. observability [541]. ocean [242]. old [362, 502]. omni [316]. omniscience [182]. One [632, 173, 356]. One-world [632]. oneself [210]. only [217]. Ontological [713, 87, 217]. Ontologically [746]. Ontology [279, 350, 302, 349, 150]. opening [122]. Operators [273, 503]. Optimization [529]. oracle [401]. orbs [41]. Order [727, 650, 89, 196]. Ordinary [784, 786]. organism [224]. organized [204]. oriented [552, 263]. origin [465]. origins [43]. Oscar [183]. Other [809, 692, 126]. Otto [442]. ought [139]. our [289]. Outline [406]. Outstanding [285]. Overdetermination [706]. Oxford [38].

pain [129]. Paley [200]. pantomimed [457]. papers [23, 68, 472, 518, 124]. Papier [694]. Paradigm [731]. Paradigm-Based [731]. Paradox [753, 667, 89, 273, 103, 555, 272, 135, 130, 539, 525]. Paradoxes [577]. Parallelism [576]. Pareto [750]. Paribus [754]. Part [696, 16, 8, 24, 149]. particle [432]. particles [6]. Partitions [103]. Pascalian [536]. Passion [736]. past [72]. pathways [28]. patients [457]. Patterns [164, 630]. Pavlov [141]. pawn [398]. peace [401]. Pearson [192]. Peirce [512, 380, 513, 514, 634, 633]. Peircean [752]. people [274]. perception [491]. perceptual [262, 455]. Perez [772]. perfect [316]. period [163]. permanence [72]. permanent [283]. pernicious [148]. person [460, 99]. personal [291, 157]. Perspectival [662]. Perspective [751, 54, 241, 479]. Perspectives [486]. Perspectivism [675]. Petrus [304]. Phase [770]. phenomena [166]. Phenomenalism [756]. Phenomenology [582, 736, 600, 601, 756]. Philip [267]. Philosophical [607, 655, 454, 355, 702, 324]. philosophies [514, 121]. Philosophizing [774]. Philosophy [206, 517, 632, 646, 619, 626, 660, 669, 674, 680, 697, 703, 692, 686, 711, 718, 724, 732, 748, 756, 765, 736, 687, 77, 142, 127, 147, 608, 604, 118, 636, 583, 255, 50, 484, 144, 486, 504, 322, 192, 312, 327, 120, 477, 70, 222, 478, 216, 148, 131, 313, 435, 151, 430, 540]. Phylosophy [652]. Physical [768, 175, 32, 33, 46, 5, 325]. Physicalism [714, 496]. physicalists [542]. Physicists [598]. Physics [4, 93, 580, 792, 162, 266, 495, 174, 430]. Piagetian [54]. Picture [770, 282]. Pictures [763]. Pierre [48, 46]. Place [618, 352, 201]. planning [110]. platonist [450]. plural [311]. Pluralism [706, 217, 114]. Poincaré [195]. point [383]. policy [400]. politics [476]. Ponty [737, 736]. Port [321]. possibilities [181, 242]. Possibility [735, 717]. Possible [771, 709, 87]. Postulate [651, 170]. power [259, 156]. Practical [720, 125]. practice [163, 93]. practices [224]. pragmatic [513, 387]. Pragmatism [687, 254]. Prawitz [348]. Pre [737, 322]. pre-established [322]. Pre-Reflective [737]. Prècis [611]. precise [494]. predecessors [35]. Predicates [625, 624, 290]. Predication [663, 316, 318]. predictability [172]. Predicting [191]. prediction [160, 328]. Predictivism [582]. Preemption [71]. Preface [252, 8, 16, 24, 338, 117, 497, 159, 596, 300, 473, 464, 360]. preference [569]. preliminary [149]. premises [388]. present [339]. price [530]. Priest [299]. primacy [510]. prime [230]. Principle [750, 169, 139, 372, 357, 427, 106, 578]. Principles [412]. print [423]. Priori [807]. Prisoner [100, 102, 413, 373]. probabilistic [535, 103, 336]. Probabilities [802, 565, 566]. Probability [749, 445, 670, 194, 190, 193]. Problem [652, 632, 646, 660, 669, 674, 680, 697, 703, 756, 591, 754, 595, 100, 787, 647, 102, 730, 95, 326, 137, 76, 126, 19, 399, 258, 182, 481]. Problematics [222]. Problems [652, 632, 646, 619, 626, 660, 669, 674, 680, 697, 703, 692, 686, 711, 718, 724, 732, 748, 756, 694, 175, 383, 285, 99]. procedure [259]. process [177, 351]. Production [705]. Prof. [526, 299]. progress [153, 374]. progressive [135]. project [37]. projection [170]. promise [156]. Promoting [346]. Proof [597, 598, 601, 644, 51]. Proofs [795, 657, 138]. Propensities [174, 539]. Proper [666]. Properties [621]. Propositional [350, 12]. Propositions [787]. Prospects [147]. protocol [239]. prototype [92]. Provisos [754]. Proxy [593]. Psychological [681, 672, 80]. Psychologism [84]. Psychologists [691]. Psychology [743, 219, 83, 134, 161]. Psychophysical [741, 292, 186]. Puccini [226]. Pure [735, 441]. Pursuing [512]. Putnam [444]. Putting [352]. Pythagoreanism [621].

QFT [712]. qua [140]. Qualia [507, 806, 337]. quantification [296]. quantifier [344]. Quantifying [405]. Quantities [768]. quantization [169]. Quantum [615, 6, 622, 801, 580, 170, 713, 670, 325, 651, 326, 522, 519, 521, 94, 523, 424, 429, 171, 323, 431, 520, 173, 433]. Quasi [334]. Quasi-realist [334]. queerness [227]. Quest [384]. question [451, 474, 387]. Questions [782, 383, 553]. Quine [196, 348, 270, 347]. quoque} [527]. Quotation [588].

Rabbit [800]. Ramifications [43]. Ramsey [409, 208]. Random [522]. Ranking [625]. Ratiocinator [627]. Rational [239, 374, 413, 569]. Rationalising [53]. Rationality [255, 793, 340, 179, 52, 73, 272, 368, 176, 22, 371, 257, 15, 560]. rationis [302, 349]. ratios [60, 351]. re [84, 705]. re-evaluation [84]. Reach [755]. Realism [162, 661, 167, 722, 708, 676, 505, 196, 444, 40, 541, 93, 1, 56, 216, 80, 115, 436]. Realisms [74]. Realist [677, 557, 154, 334]. Reality [682, 168]. Realization [672]. really [228]. reals [405]. Reason [441]. Reasonableness [259]. Reasoning [720, 721, 332, 585, 105, 136]. Reciprocal [589]. Recognition [709]. recollection [238]. reconception [281, 284]. Reconceptions [276]. reconcile [160]. Reconsideration [639]. reconsidered [84, 413]. reconsidered- [413]. Reconsidering [498]. Reconstruction [611, 254, 238, 302, 239]. Recursion [751]. red [220]. Reducibility [728]. Reduction [616, 605, 207]. reductionism [542, 207]. reductionist [83]. reexamined [419]. Reference [680, 661, 593, 623, 632, 311, 468, 152, 294, 249, 465, 467, 466]. References [480]. refined [247]. reflection [353]. Reflections [514, 353, 779, 244, 490]. Reflective [737, 368]. Reflexive [739]. refutation [238]. region [399]. Regularities [792]. regularity [20]. Reid [774]. Reification [568]. Reinventing [416]. Reinvention [585]. reject [192]. rejected [40]. rejection [202]. relates [141]. relation [325, 60]. Relations [713, 446, 137, 510]. Relationship [716, 552, 161]. relationships [396]. Relative [62]. Relativism [722, 162, 75, 167]. relativity [262, 433]. relevance [410, 336]. Relevant [567]. Reliabilism [335]. reliable [565]. Remarks [122, 73, 355]. renormalisation [428]. renormalization [324]. replace [284]. Replies [285]. Reply [613, 642, 436, 267, 369, 415, 467, 466, 370, 299]. Reporting [367]. Representation [699, 701, 738, 459, 514, 499, 394, 136]. Representationalism [734, 628]. Representations [422, 617]. representing [492, 420]. reproducibility [365]. Research [33]. resemblance [514]. resemblances [488]. response [474]. Responses [260]. responsible [82]. Rethinking [742, 200]. Return [691]. reunion [203]. Review [57, 691, 212, 523, 504, 97, 570, 107, 7, 463, 516, 96, 58, 309, 213, 189, 407, 313]. revision [208]. Revisionism [466]. Revisited [633, 130]. revived [227]. Revolution [599, 523]. Riddle [731, 280]. Right [618, 375, 457]. rightness [283]. Rigorous [51]. Risk [562]. Road [621, 257]. Robert [364, 609, 636]. rock [547]. rocks [166]. Role [660, 682, 533, 398, 302, 349]. roles [212]. Rome [305]. Room [753, 287]. roots [212]. routines [76]. Royal [321]. Rubin [343]. Rudolf [235]. Rule [652, 774, 365, 489, 5, 128]. Rule-Following [652, 489, 128]. Rules [720, 721, 386, 722, 726, 422]. Rumford [366]. Russell [669, 581, 357, 106].

said [256]. sampling [538]. Sandu [348]. say [274]. scenario [299]. Scheffler [261, 255, 253, 257]. schemes [384]. Schlick [140]. Science [206, 517, 589, 687, 77, 142, 38, 638, 673, 361, 630, 606, 441, 583, 459, 267, 412, 278, 504, 417, 327, 477, 543, 478, 216, 134, 479, 150]. sciences [61, 329, 38, 332, 131, 39]. Scientific [677, 383, 93, 794, 328, 762, 153, 330, 154, 406, 161, 141]. scope [692, 329]. Scott [467]. Scotus [301]. search [5]. Searle [551]. Second [659]. secunda [308]. Seeing [341]. Segerberg [76]. Selection [781]. Self [743, 414, 588, 771, 180, 440, 705, 738]. self-awareness [738]. Self-deception [414]. Self-deceptive [180]. Self-Excitation [771]. Self-Knowledge [743, 588]. Self-Quotation [588]. Self-re-Production [705]. Self-understanding [440]. Semantic [377, 733, 295, 810, 556]. Semantics [132, 716, 380, 321, 17, 302, 349, 87, 18, 481]. Sense [632, 715, 437, 161]. Sense-reference [632]. Sentence [788, 592, 767, 239]. Sentence-Type [767]. sentences [38, 295]. seriously [297]. Set [575, 311, 92, 67]. Sets [753]. seventeenth [138]. shade [220]. Shape [605, 326]. shaping [166]. Shimony [526]. Shooting [753]. Shooting-Room [753]. Should [284]. Sightings [786]. significance [566]. signs [514]. simple [228]. Simplicity [707, 529, 532, 563]. simplification [532]. Simpson [103]. sine [140]. single [413, 28]. small [662, 538]. Smith [348, 525, 466]. Smythies [337]. Soames [467]. Social [794, 719, 340, 329, 330, 327, 328, 332]. society [156]. Socrates [219]. solipsism [490]. Solution [731, 280]. solutions [95]. Solve [647]. Some [95, 410, 809, 29, 371, 73, 659, 155, 779, 379, 102, 313, 175, 94, 277, 35]. sorites [89]. Soto [36]. Soul [583]. Sources [355]. Space [769, 770, 623, 635, 624, 234, 451, 432]. Space-Time [769, 234, 451, 432]. spaces [499, 290]. spacetime [452]. Spatiality [730]. speaker [505]. Special [797, 399]. Species [779, 318]. spectra [11]. Spectral [617]. Spemann [202]. Spin [671]. Sprigge [343]. square [362]. stability [397]. Stance [746, 202]. Standard [807, 475]. state [547]. Statements [666]. States [672]. statics [43]. Statistical [651, 585]. Statistics [797, 6, 192]. status [138]. Stein [369]. steps [410]. Stit [111]. strategic [697]. Strategizing [401]. stress [134]. strict [197]. Strong [694, 228, 551, 474]. Structural [803, 136, 175, 67, 447]. structuralist [65]. Structure [799, 663, 657, 650, 231, 432, 46]. Structures [696, 625, 290, 204]. student [244]. studies [491]. Study [638, 583, 302, 164, 28, 449]. subalternate [38]. Subjective [519, 194, 565]. substance [317]. Substances [606]. Substantive [717]. Subtle [640]. success [81]. suggestions [496]. superassertibility [557]. superior [228]. Superposition [579]. Supertask [772]. Supertasks [668]. Supervaluational [56]. Supervenience [714, 503, 185, 132, 85, 292, 186, 502, 510]. Suppes [625, 624]. support [535]. surgical [163]. Surprise [723, 690]. surprising [29]. surrogative [136]. syllogisms [316]. Syllogistic [699]. symbolic [177]. Symbols [515, 253]. symmetric [456]. syntactic [422, 17]. Syntax [443, 243, 242]. Syntheticity [633]. System [629, 432, 397, 27]. Systematicity [726, 418]. Systems [739, 710, 805, 421, 521, 279]. Szaniawski [78].

tableaux [378]. Taking [297]. talking [216]. Tarskian [729]. task [351].

Tautology [119]. Taxonomy [672]. Technical [563]. teleology [199]. Temporally [263]. tense [520]. Tensed [788, 592, 525]. Tenseless [767]. Terms [716, 321]. Test [458, 409, 208]. testability [22]. testable [15]. Tests [88]. texture [28]. Their [660, 802, 265, 483]. theme [238]. Themes [636]. theology [38]. Theorem [701, 426]. Theorems [146, 543]. Theoretic [751, 569, 432, 30, 481]. Theoretical [580, 380, 476, 134]. Theories [446, 746, 495, 412, 33, 428, 447, 15, 223]. Theory [739, 749, 627, 648, 589, 615, 701, 363, 752, 758, 793, 527, 575, 590, 580, 152, 786, 581, 1, 717, 767, 690, 326, 163, 552, 522, 310, 364, 324, 160, 311, 492, 410, 514, 32, 223, 76, 374, 92, 194, 67, 468, 336, 534, 171, 164, 46, 567, 454, 20, 323, 525, 193, 532, 562, 357, 406, 563, 168, 465, 467, 173, 325, 60, 213, 553, 433]. There [793, 754, 799, 390, 230]. theses [2, 141]. thesis [42]. things [341, 204]. third [390, 99, 359]. Thomson [704]. Thought [742, 25, 716, 726, 177, 508, 14, 306, 515]. Thoughts [659]. Three [765, 760, 713, 418, 802, 249, 528, 10]. Three-concept [418]. Time [622, 769, 618, 431, 520, 670, 623, 767, 430, 635, 620, 624, 234, 544, 451, 432, 525]. Today [809]. Tolerance [443]. toll [425]. Too [796, 347]. tools [484]. Topics [303]. topological [290]. town [217]. Tractatus [795, 629]. trade [402]. Traditional [581]. tragedy [373]. Transcendental [440, 636, 69, 70]. Transformation [606]. Transient [526]. Transitions [620]. Translation [638]. translational [485]. Transzendenz [658]. treatment [163]. tree [221]. triadic [396]. Triffids [489]. trouble [337]. Troubles [227, 85]. true [388, 225]. Truth [675, 788, 729, 482, 283, 593, 592, 717, 351, 266, 91, 248, 445, 557, 444, 247, 165, 243]. Truth-ratios [351]. Truthlikeness [248, 415]. Truths [646, 554]. tu [527]. Turing [549, 811, 550, 458]. TV [337]. TV-hood [337]. twin [227]. Two [692, 456, 561, 631, 322, 683, 496, 27, 291, 270, 223]. Two-dimensional [456]. Type [767]. Types [788, 592, 357]. Typescript [127].

ulcer [163]. ultrafilters [522]. Uncertainty [398]. Unconfirmed [786]. Undecidability [793]. Undercutting [409]. Underdetermination [712, 580]. Undermind [509]. Understanding [278, 762, 284, 440, 406, 26]. Unification [775, 616, 780, 755]. Unification- [775]. Union [206, 517, 77, 142, 498]. Universal [792]. universe [426, 319]. unlimited [242]. unmaking [469]. unorthodox [424]. unpredictable [191]. untapped [259]. upon [148]. USA [469]. Use [679, 690, 119, 233, 457]. uses [315]. utility [194, 372, 268].

Vacuous [91]. Vagueness [793, 89, 507]. validity [561, 555]. Value [768]. values [479]. variability [382]. variables [522, 427]. Varieties [660]. Verdi [226]. verisimilitude [153, 534]. Version [767]. Versions [683, 102]. versus [196]. Very [590]. Vienna [761, 115]. View [784, 638, 383, 65, 141]. viewpoint [697]. views [126, 249]. Vindicated [572, 703]. virtues [476, 558]. vision [529, 506, 455, 28]. visual [27]. voluntaristic [354]. vs [680, 756, 400].

W [415]. wagers [536]. Wake [623]. Want [755]. war [201]. Was [642, 245, 137]. way [49]. Weinberg [712]. Weyl [582]. Where [66, 27]. whole [445]. wholism [160, 120]. whom [425]. Why-Questions [782]. will [109, 306]. Williams [540]. windowpane [490]. Without [669, 785, 422, 248, 420, 393, 491]. Wittgenstein [652, 678, 795, 122, 121, 126, 129, 487, 355, 679, 128, 500, 747, 357, 125, 358, 763, 151]. Wittgensteinian [131]. Wodeham [308]. Woods [642]. Word [679, 119]. work [241, 358]. working [484]. World [796, 632, 404, 471, 146, 87, 237, 325]. worldly [277]. Worlds [662, 709]. worries [277]. Writings [261, 659]. wrong [394, 245]. Wrought [654].

Yu [298].

Zeno [577, 528, 233]. Zero [768]. Zero-Value [768]. ZiF [285]. Zoöepistemology [217]. Zuordnung [140].

References

Morrison:1990:TIR

[1] Margaret Morrison. Theory, intervention and realism. Synthese, 82(1): 1–22, January 1990. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

     0964 (electronic). URL http://link.springer.com/article/10.1007/
     BF00413667.

Millman:1990:FGD

[2] Arthur B. Millman. Falsification and Grünbaum’s Duhemian theses. Synthese, 82(1):23–52, January 1990. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

     article/10.1007/BF00413668.

Sibelius:1990:NCM

[3] Patrick Sibelius. Necessary connexions in mechanics. Synthese, 82(1): 53–76, January 1990. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

     0964 (electronic). URL http://link.springer.com/article/10.1007/
     BF00413669.

Arntzenius:1990:PCC

[4] Frank Arntzenius. Physics and common causes. Synthese, 82(1):77– 96, January 1990. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

     0964 (electronic). URL http://link.springer.com/article/10.1007/
     BF00413670.

Martinez:1990:SPC

[5] Sergio Martinez. A search for the physical content of Luders’ rule. Synthese, 82(1):97–125, January 1990. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/BF00413671.

Dieks:1990:QSI

[6] Dennis Dieks. Quantum statistics, identical particles and correlations. Synthese, 82(1):127–155, January 1990. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/BF00413672.

Redhead:1990:R

[7] Michael Redhead and Richard M. Burian. Review. Synthese, 82(1):157– 174, January 1990. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF00413673.

F:1990:PP

[8] J. H. F. Preface to Part I. Synthese, 82(2):175, February 1990. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://

link.springer.com/article/10.1007/BF00413659.

Anonymous:1990:Aa

[9] Anonymous. Announcement. Synthese, 82(2):176, February 1990. CO-DEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/BF00413660.

Ramsey:1990:CTL

[10] William Ramsey and Stephen Stich. Connectionism and three levels of nativism. Synthese, 82(2):177–205, February 1990. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.

com/article/10.1007/BF00413661.

Cole:1990:FIS

[11] David Cole. Functionalism and inverted spectra. Synthese, 82(2):207– 222, February 1990. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF00413662.

Bechtel:1990:BEP

[12] William Bechtel and Adele A. Abrahamsen. Beyond the exclusively propositional era. Synthese, 82(2):223–253, February 1990. CODEN SYN-TAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.

springer.com/article/10.1007/BF00413663.

Thagard:1990:CCC

[13] Paul Thagard. Concepts and conceptual change. Synthese, 82(2):255– 274, February 1990. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF00413664.

Dunlop:1990:CDL

[14] Charles E. M. Dunlop. Conceptual dependency as the language of thought. Synthese, 82(2):275–296, February 1990. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/BF00413665.

Smokler:1990:TRE

[15] Howard Smokler. Are theories of rationality empirically testable? Synthese, 82(2):297–306, February 1990. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/BF00413666.

F:1990:PPIa

[16] J. H. F. Preface to part II. Synthese, 82(3):307, March 1990. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://

link.springer.com/article/10.1007/BF00413878.

Jahren:1990:CSS

[17] Neal Jahren. Can semantics be syntactic? Synthese, 82(3):309–328, March 1990. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic).

URL http://link.springer.com/article/10.1007/BF00413879.

Wilks:1990:FCS

[18] Yorick Wilks. Form and content in semantics. Synthese, 82(3):329– 351, March 1990. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF00413880.

Lormand:1990:FFP

[19] Eric Lormand. Framing the frame problem. Synthese, 82(3):353– 374, March 1990. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF00413881.

Morris:1990:KRT

[20] William Edward Morris. Knowledge and the regularity theory of information. Synthese, 82(3):375–398, March 1990. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.

com/article/10.1007/BF00413882.

Graham:1990:ME

[21] George Graham. Melancholic epistemology. Synthese, 82(3):399–422, March 1990. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF00413883.

Fetzer:1990:ERT

[22] James H. Fetzer. Evolution, rationality, and testability. Synthese, 82 (3):423–439, March 1990. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/BF00413884.

Anonymous:1990:CPa

[23] Anonymous. Call for papers. Synthese, 82(3):440, March 1990. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://

link.springer.com/article/10.1007/BF00413885.

F:1990:PPIb

[24] J. H. F. Preface to Part III. Synthese, 83(1):1, April 1990. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://

link.springer.com/article/10.1007/BF00413685.

Braddon-Mitchell:1990:ELT

[25] David Braddon-Mitchell and John Fitzpatrick. Explanation and the language of thought. Synthese, 83(1):3–29, April 1990. CODEN SYN-TAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.

springer.com/article/10.1007/BF00413686.

Zemach:1990:HU

[26] Eddy M. Zemach. Human understanding. Synthese, 83(1):31–48, April 1990. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic).

URL http://link.springer.com/article/10.1007/BF00413687.

Vaina:1990:WWH

[27] Lucia M. Vaina. ‘What’ and ‘where’ in the human visual system: Two hierarchies of visual modules. Synthese, 83(1):49–91, April 1990. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://

link.springer.com/article/10.1007/BF00413688.

Vaina:1990:CFP

[28] Lucia M. Vaina. Common functional pathways for texture and form vision: A single case study. Synthese, 83(1):93–131, April 1990. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://

link.springer.com/article/10.1007/BF00413689.

Hintikka:1990:CCE

[29] Jaakko Hintikka. The Cartesian cogito, epistemic logic and neuroscience: Some surprising interrelations. Synthese, 83(1):133–157, April 1990. CO-DEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/BF00413690.

Weintraub:1990:DTE

[30] Ruth Weintraub. Decision-theoretic epistemology. Synthese, 83(1):159– 177, April 1990. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF00413691.

Ariew:1990:I

[31] Roger Ariew and Peter Barker. Introduction. Synthese, 83(2):179– 182, May 1990. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF00413754.

Duhem:1990:LEP

[32] Pierre Duhem. Logical examination of physical theory. Synthese, 83(2): 183–188, May 1990. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF00413755.

Duhem:1990:RHP

[33] Pierre Duhem. Research on the history of physical theories. Synthese, 83(2):189–200, May 1990. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/BF00413756.

Ragep:1990:DAH

[34] F. Jamil Ragep. Duhem, the Arabs, and the history of cosmology. Synthese, 83(2):201–214, May 1990. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/BF00413757.

Menn:1990:DSP

[35] Stephen Menn. Descartes and some predecessors on the divine conservation of motion. Synthese, 83(2):215–238, May 1990. CODEN SYN-TAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.

springer.com/article/10.1007/BF00413758.

Wallace:1990:DKD

[36] William Wallace. Duhem and Koyré on Domingo de Soto. Synthese, 83(2):239–260, May 1990. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/BF00413759.

Westman:1990:DHP

[37] Robert S. Westman. The Duhemian historiographical project. Synthese, 83(2):261–272, May 1990. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/BF00413760.

Livesey:1990:STF

[38] Steven J. Livesey. Science and theology in the fourteenth century: The subalternate sciences in Oxford commentaries on the sentences. Synthese, 83(2):273–292, May 1990. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/BF00413761.

Wilks:1990:CCC

[39] Yorick Wilks. Christopher Clavius and the classification of sciences. Synthese, 83(2):293–300, May 1990. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/BF00413762.

Goddu:1990:RDR

[40] André Goddu. The realism that Duhem rejected in Copernicus. Synthese, 83(2):301–315, May 1990. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/BF00413763.

Barker:1990:COE

[41] Peter Barker. Copernicus, the orbs, and the equant. Synthese, 83(2): 317–323, May 1990. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF00413764.

Brenner:1990:HCA

[42] Anastasios A. Brenner. Holism a century ago: The elaboration of Duhem’s thesis. Synthese, 83(3):325–335, June 1990. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.

com/article/10.1007/BF00413419. See comment [44].

Martin:1990:DOS

[43] R. N. D. Martin. Duhem and the origins of statics: Ramifications of the crisis of 1903–04. Synthese, 83(3):337–355, June 1990. CODEN SYN-TAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.

springer.com/article/10.1007/BF00413420.

Quinn:1990:DDC

[44] Philip L. Quinn. Duhem in different contexts: Comments on Brenner and Martin. Synthese, 83(3):357–362, June 1990. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.

com/article/10.1007/BF00413421.

Howard:1990:ED

[45] Don Howard. Einstein and Duhem. Synthese, 83(3):363–384, June 1990. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/BF00413422.

Maiocchi:1990:PDA

[46] Roberto Maiocchi. Pierre Duhem’s the aim and structure of physical theory: A book against conventionalism. Synthese, 83(3):385–400, June 1990. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic).

URL http://link.springer.com/article/10.1007/BF00413423.

Burian:1990:MDH

[47] Richard M. Burian. Maiocchi on Duhem, Howard on Duhem and Einstein: Historiographical comments. Synthese, 83(3):401–408, June 1990. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/BF00413424.

Lugg:1990:PDC

[48] Andrew Lugg. Pierre Duhem’s conception of natural classification. Synthese, 83(3):409–420, June 1990. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/BF00413425.

McMullin:1990:CDM

[49] Ernan McMullin. Comment: Duhem’s middle way. Synthese, 83(3): 421–430, June 1990. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF00413426.

Crowe:1990:DHP

[50] Michael J. Crowe. Duhem and history and philosophy of mathematics. Synthese, 83(3):431–447, June 1990. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/BF00413427. See comments [51].

Jesseph:1990:RPH

[51] Douglas Jesseph. Rigorous proof and the history of mathematics: Comments on Crowe. Synthese, 83(3):449–453, June 1990. CODEN SYN-TAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.

springer.com/article/10.1007/BF00413428. See [50].

Gilbert:1990:RCC

[52] Margaret Gilbert. Rationality, coordination, and convention. Synthese, 84(1):1–21, July 1990. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF00485004.

Miller:1990:RC

[53] Seumas Miller. Rationalising conventions. Synthese, 84(1):23–41, July 1990. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic).

URL http://link.springer.com/article/10.1007/BF00485005.

Arbib:1990:PPM

[54] Michael A. Arbib. A Piagetian perspective on mathematical construction. Synthese, 84(1):43–58, July 1990. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/BF00485006.

Barnes:1990:LDA

[55] Eric Barnes. The language dependence of accuracy. Synthese, 84(1): 59–95, July 1990. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF00485007.

Rasmussen:1990:SAR

[56] Stig Alstrup Rasmussen. Supervaluational anti-realism and logic. Synthese, 84(1):97–138, July 1990. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/BF00485008.

Adams:1990:RA

[57] Ernest W. Adams. Review article. Synthese, 84(1):139–152, July 1990. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/BF00485009. See rpely

[415].

Strawson:1990:R

[58] P. F. Strawson. Review. Synthese, 84(1):153–161, July 1990. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://

link.springer.com/article/10.1007/BF00485010.

Anonymous:1990:E

[59] Anonymous. Erratum. Synthese, 84(1):163, July 1990. CODEN SYN-TAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://

link.springer.com/article/10.1007/BF00485011;   http://link.
springer.com/content/pdf/10.1007/BF00485011.pdf.

Stein:1990:EDA

[60] Howard Stein. Eudoxos and Dedekind: On the ancient Greek theory of ratios and its relation to modern mathematics. Synthese, 84(2):163– 211, August 1990. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF00485377.

Friedman:1990:KCI

[61] Michael Friedman. Kant on concepts and intuitions in the mathematical sciences. Synthese, 84(2):213–257, August 1990. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.

com/article/10.1007/BF00485378.

Sieg:1990:RCA

[62] Wilfried Sieg. Relative consistency and accessible domains. Synthese, 84 (2):259–297, August 1990. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/BF00485379.

Anonymous:1990:Ab

[63] Anonymous. Announcement. Synthese, 84(2):299, August 1990. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://

link.springer.com/article/10.1007/BF00485380.

Hintikka:1990:E

[64] Jaakko Hintikka. Editorial. Synthese, 84(3):301–302, September 1990. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/BF00485185.

Parsons:1990:SVM

[65] Charles Parsons. The structuralist view of mathematical objects. Synthese, 84(3):303–346, September 1990. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/BF00485186.

Hodes:1990:WDN

[66] Harold T. Hodes. Where do the natural numbers come from? Synthese, 84 (3):347–407, September 1990. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/BF00485187.

Hellman:1990:TMS

[67] Geoffrey Hellman. Toward a modal-structural interpretation of set theory. Synthese, 84(3):409–443, September 1990. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/BF00485188.

Anonymous:1990:CPb

[68] Anonymous. Call for papers. Synthese, 84(3):445, September 1990. CO-DEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/BF00485189.

Baldner:1990:TIC

[69] Kent Baldner. Is transcendental idealism coherent? Synthese, 85(1): 1–23, October 1990. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF00873192.

Pereboom:1990:KJT

[70] Derk Pereboom. Kant on justification in transcendental philosophy. Synthese, 85(1):25–54, October 1990. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/BF00873193.

Ehring:1990:PDC

[71] Douglas Ehring. Preemption, direct causation, and identity. Synthese, 85(1):55–70, October 1990. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/BF00873194.

Oddie:1990:BCP

[72] Graham Oddie. Backwards causation and the permanence of the past. Synthese, 85(1):71–93, October 1990. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/BF00873195.

Indurkhya:1990:SRR

[73] Bipin Indurkhya. Some remarks on the rationality of induction. Synthese, 85(1):95–114, October 1990. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/BF00873196.

Levin:1990:R

[74] Michael Levin. Realisms. Synthese, 85(1):115–138, October 1990. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://

link.springer.com/article/10.1007/BF00873197.

Hugly:1990:MRD

[75] Philip Hugly and Charles Sayward. Moral relativism and deontic logic. Synthese, 85(1):139–152, October 1990. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/BF00873198.

Elgesem:1990:IAR

[76] Dag Elgesem. Intentions, actions and routines: A problem in Krister Segerberg’s theory of action. Synthese, 85(1):153–177, October 1990. CO-DEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/BF00873199.

Hilpinen:1990:IUH

[77] Risto Hilpinen. International Union of the History and Philosophy of Science Division of Logic, Methodology and Philosophy of Science Bulletin No. 14. Synthese, 85(1):179–183, October 1990. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.

com/article/10.1007/BF00873200.

Wolenski:1990:OKS

[78] Jan Woleński. Obituary of Klemens Szaniawski (1925–1990). Synthese, 85(1):184, October 1990. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/BF00873201.

Boden:1990:IE

[79] Margaret A. Boden. Interdisciplinary epistemology. Synthese, 85(2):185– 197, November 1990. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF00484791.

Taylor:1990:EJP

[80] James E. Taylor. Epistemic justification and psychological realism. Synthese, 85(2):199–230, November 1990. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/BF00484792.

Clarke:1990:ENE

[81] Murray Clarke. Epistemic norms and evolutionary success. Synthese, 85 (2):231–244, November 1990. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/BF00484793.

Greco:1990:IER

[82] John Greco. Internalism and epistemically responsible belief. Synthese, 85 (2):245–277, November 1990. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/BF00484794.

Montgomery:1990:RIC

[83] Richard Montgomery. The reductionist ideal in cognitive psychology. Synthese, 85(2):279–314, November 1990. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/BF00484795.

Aach:1990:PRR

[84] John Aach. Psychologism reconsidered: A re-evaluation of the arguments of Frege and Husserl. Synthese, 85(2):315–338, November 1990. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://

link.springer.com/article/10.1007/BF00484796.

Klagge:1990:DTS

[85] James C. Klagge. Davidson’s troubles with supervenience. Synthese, 85 (2):339–352, November 1990. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/BF00484797.

Anonymous:1990:Ac

[86] Anonymous. Announcements. Synthese, 85(2):353–354, November 1990. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/BF00484798.

Menzel:1990:AOC

[87] Christopher Menzel. Actualism, ontological commitment, and possible world semantics. Synthese, 85(3):355–389, December 1990. CODEN SYN-TAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.

springer.com/article/10.1007/BF00484834.

Gillon:1990:AGI

[88] Brendan S. Gillon. Ambiguity, generality, and indeterminacy: Tests and definitions. Synthese, 85(3):391–416, December 1990. CODEN SYN-TAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.

springer.com/article/10.1007/BF00484835.

Burgess:1990:SPH

[89] J. A. Burgess. The sorites paradox and higher-order vagueness. Synthese, 85(3):417–474, December 1990. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/BF00484836.

Bacchus:1990:AC

[90] F. Bacchus, H. E. Kyburg, Jr., and M. Thalos. Against conditionalization. Synthese, 85(3):475–506, December 1990. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/BF00484837.

Almeder:1990:VT

[91] Robert Almeder. Vacuous truth. Synthese, 85(3):507–524, December 1990. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/BF00484838.

Fuhrmann:1991:NIP

[92] Gy. Fuhrmann. Note on the integration of prototype theory and fuzzy-set theory. Synthese, 86(1):1–27, January 1991. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.

com/article/10.1007/BF00485412.

Hones:1991:SRE

[93] Michael J. Hones. Scientific realism and experimental practice in highenergy physics. Synthese, 86(1):29–60, January 1991. CODEN SYN-TAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.

springer.com/article/10.1007/BF00485413.

Dieks:1991:SAD

[94] Dennis Dieks. On some alleged difficulties in the interpretation of quantum mechanics. Synthese, 86(1):77–86, January 1991. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.

com/article/10.1007/BF00485414.

Albert:1991:MPS

[95] David Z. Albert and Barry Loewer. The measurement problem: Some “solutions”. Synthese, 86(1):87–98, January 1991. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.

com/article/10.1007/BF00485415.

Stairs:1991:RE

[96] Allen Stairs. Review essay. Synthese, 86(1):99–122, January 1991. CO-DEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/BF00485416.

Howard:1991:RE

[97] Don Howard. Review essay. Synthese, 86(1):123–141, January 1991. CO-DEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/BF00485417.

Anonymous:1991:Aa

[98] Anonymous. Announcement. Synthese, 86(1):142, January 1991. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://

link.springer.com/article/10.1007/BF00485418.

Sobel:1991:NDT

[99] Jordan Howard Sobel. Non-dominance, third person and non-action Newcomb problems, and metatickles. Synthese, 86(2):143–172, February 1991. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/BF00485805.

Hurley:1991:NPP

[100] S. L. Hurley. Newcomb’s problem, prisoners’ dilemma, and collective action. Synthese, 86(2):173–196, February 1991. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.

com/article/10.1007/BF00485806.

Anonymous:1991:EN

[101] Anonymous. Editorial note. Synthese, 86(2):196, February 1991. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://

link.springer.com/article/10.1007/BF00485807.

Sobel:1991:SVN

[102] Jordan Howard Sobel. Some versions of Newcomb’s problem are prisoners’ dilemmas. Synthese, 86(2):197–208, February 1991. CODEN SYN-TAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.

springer.com/article/10.1007/BF00485808.

Hardcastle:1991:PPC

[103] Valerie Gray Hardcastle. Partitions, probabilistic causal laws, and Simpson’s paradox. Synthese, 86(2):209–228, February 1991. CODEN SYN-TAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.

springer.com/article/10.1007/BF00485809.

Cleland:1991:IE

[104] Carol Cleland. On the individuation of events. Synthese, 86(2):229– 254, February 1991. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF00485810.

Janssen:1991:WFN

[105] M. C. W. Janssen and Y.-H. Tan. Why Friedman’s non-monotonic reasoning defies Hempel’s covering law model. Synthese, 86(2):255–284, February 1991. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic).

URL http://link.springer.com/article/10.1007/BF00485811.

daCosta:1991:RPI

[106] Newton C. A. da Costa and Steven French. On Russell’s principle of induction. Synthese, 86(2):285–295, February 1991. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.

com/article/10.1007/BF00485812.

Rantala:1991:R

[107] Veikko Rantala. Review. Synthese, 86(2):297–319, February 1991. CO-DEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/BF00485813.

Sapire:1991:GC

[108] David Sapire. General causation. Synthese, 86(3):321–347, March 1991. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/BF00485264.

Garcia:1991:IW

[109] J. L. A. Garcia. On the irreducibility of the will. Synthese, 86(3):349– 360, March 1991. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF00485265.

Audi:1991:ICC

[110] Robert Audi. Intention, cognitive commitment, and planning. Synthese, 86(3):361–378, March 1991. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/BF00485266.

Perloff:1991:SLA

[111] Michael Perloff. Stit and the language of agency. Synthese, 86(3):379– 408, March 1991. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF00485267.

White:1991:KDC

[112] James L. White. Knowledge and deductive closure. Synthese, 86(3):409– 423, March 1991. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF00485268.

Gjelsvik:1991:DKC

[113] Olav Gjelsvik. Dretske on knowledge and content. Synthese, 86(3):425– 441, March 1991. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF00485269.

Rouse:1991:IEE

[114] Joseph Rouse. Indeterminacy, empirical evidence, and methodological pluralism. Synthese, 86(3):443–465, March 1991. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.

com/article/10.1007/BF00485270.

Young:1991:CAR

[115] James O. Young. Coherence, anti-realism and the Vienna circle. Synthese, 86(3):467–482, March 1991. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/BF00485271.

Anonymous:1991:Ab

[116] Anonymous. Announcement. Synthese, 86(3):483–484, March 1991. CO-DEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/BF00485272.

Hintikka:1991:P

[117] Jaakko Hintikka. Preface. Synthese, 87(1):1, April 1991. CODEN SYN-TAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.

springer.com/article/10.1007/BF00485327.

Wittgenstein:1991:P

[118] Ludwig Wittgenstein and Heikki Nyman. Philosophy. Synthese, 87(1): 3–22, April 1991. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF00485328.

Dreben:1991:THU

[119] Burton Dreben and Juliet Floyd. Tautology: How not to use a word. Synthese, 87(1):23–49, April 1991. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/BF00485329.

Mccarty:1991:PLW

[120] David Charles Mccarty. The philosophy of logical wholism. Synthese, 87(1):51–123, April 1991. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/BF00485330.

Gerrard:1991:WPM

[121] Steve Gerrard. Wittgenstein’s philosophies of mathematics. Synthese, 87(1):125–142, April 1991. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/BF00485331.

Floyd:1991:WLO

[122] Juliet Floyd. Wittgenstein on 2, 2, 2 …: The opening of remarks on the foundations of mathematics. Synthese, 87(1):143–180, April 1991. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/BF00485332.

Anonymous:1991:Ac

[123] Anonymous. Announcements. Synthese, 87(1):181–182, April 1991. CO-DEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/BF00485333.

Hintikka:1991:IMH

[124] Jaakko Hintikka. An impatient man and his papers. Synthese, 87(2): 183–201, May 1991. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF00485399.

Stern:1991:MWL

[125] David Stern. The “Middle Wittgenstein”: From logical atomism to practical holism. Synthese, 87(2):203–226, May 1991. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.

com/article/10.1007/BF00485400.

Hark:1991:DWV

[126] M. R. M. Ter Hark. The development of Wittgenstein’s views about the other minds problem. Synthese, 87(2):227–253, May 1991. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://

link.springer.com/article/10.1007/BF00485401.

Luckhardt:1991:PBT

[127] C. Grant Luckhardt. Philosophy in the big typescript. Synthese, 87(2): 255–272, May 1991. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF00485402.

Pears:1991:WAR

[128] David Pears. Wittgenstein’s account of rule-following. Synthese, 87(2): 273–283, May 1991. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF00485403.

Kelly:1991:WMP

[129] Michael Lee Kelly. Wittgenstein and mad pain. Synthese, 87(2):285– 294, May 1991. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF00485404.

Linville:1991:MPR

[130] Kent Linville and Merrill Ring. Moore’s paradox revisited. Synthese, 87(2):295–309, May 1991. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/BF00485405.

Schatzki:1991:EWP

[131] Theodore R. Schatzki. Elements of a Wittgensteinian philosophy of the human sciences. Synthese, 87(2):311–329, May 1991. CODEN SYN-TAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.

springer.com/article/10.1007/BF00485406.

Bonevac:1991:SS

[132] Daniel Bonevac. Semantics and supervenience. Synthese, 87(3):331– 361, June 1991. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF00499816.

Burdick:1991:NAC

[133] Howard Burdick. A notorious affair called exportation. Synthese, 87(3): 363–377, June 1991. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF00499817.

Trout:1991:BAS

[134] J. D. Trout. Belief attribution in science: Folk psychology under theoretical stress. Synthese, 87(3):379–400, June 1991. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.

com/article/10.1007/BF00499818.

Lascarides:1991:PIP

[135] Alex Lascarides. The progressive and the imperfective paradox. Synthese, 87(3):401–447, June 1991. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/BF00499819.

Swoyer:1991:SRS

[136] Chris Swoyer. Structural representation and surrogative reasoning. Synthese, 87(3):449–508, June 1991. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/BF00499820.

Burdick:1991:WWL

[137] Howard Burdick. What was Leibniz’s problem about relations? Synthese, 88(1):1–13, July 1991. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF00540090.

Mancosu:1991:SPC

[138] Paolo Mancosu. On the status of proofs by contradiction in the seventeenth century. Synthese, 88(1):15–41, July 1991. CODEN SYN-TAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.

springer.com/article/10.1007/BF00540091.

Jacquette:1991:MDD

[139] Dale Jacquette. Moral dilemmas, disjunctive obligations, and Kant’s principle that ‘ought’ implies ‘can’. Synthese, 88(1):43–55, July 1991. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://

link.springer.com/article/10.1007/BF00540092.

Ryckman:1991:CSQ

[140] T. A. Ryckman. Conditio sine qua non? Zuordnung in the early epistemologies of Cassirer and Schlick. Synthese, 88(1):57–95, July 1991. CO-DEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/BF00540093.

Windholz:1991:PVI

[141] George Windholz and P. A. Lamal. Pavlov’s view of the inheritance of acquired characteristics as it relates to theses concerning scientific change. Synthese, 88(1):97–111, July 1991. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.

com/article/10.1007/BF00540094.

Hilpinen:1991:IUH

[142] Risto Hilpinen. International Union of the History and Philosophy of Science Division of Logic, Methodology and Philosophy of Science Bulletin No. 15. Synthese, 88(1):113–115, July 1991. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.

com/article/10.1007/BF00540095.

Hersh:1991:I

[143] Reuben Hersh. Introduction. Synthese, 88(2):117–118, August 1991. CO-DEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/BF00567739.

Goodman:1991:MPM

[144] Nicolas D. Goodman. Modernizing the philosophy of mathematics. Synthese, 88(2):119–126, August 1991. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/BF00567740.

Hersh:1991:MFB

[145] Reuben Hersh. Mathematics has a front and a back. Synthese, 88(2): 127–133, August 1991. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF00567741.

Krieger:1991:TMC

[146] Martin H. Krieger. Theorems as meaningful cultural artifacts: Making the world additive. Synthese, 88(2):135–154, August 1991. CODEN SYN-TAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.

springer.com/article/10.1007/BF00567742.

Maddy:1991:PMP

[147] Penelope Maddy. Philosophy of mathematics: Prospects for the 1990s. Synthese, 88(2):155–164, August 1991. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/BF00567743.

Rota:1991:PIM

[148] Gian-Carlo Rota. The pernicious influence of mathematics upon philosophy. Synthese, 88(2):165–178, August 1991. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.

com/article/10.1007/BF00567744.

Tharp:1991:MMP

[149] Leslie H. Tharp. Myth & math, Part II (preliminary draft). Synthese, 88 (2):179–199, August 1991. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/BF00567745.

Tymoczko:1991:MSO

[150] Thomas Tymoczko. Mathematics, science and ontology. Synthese, 88 (2):201–228, August 1991. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/BF00567746.

Wang:1991:PDG

[151] Hao Wang. To and from philosophy — discussions with Gödel and Wittgenstein. Synthese, 88(2):229–277, August 1991. CODEN SYN-TAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.

springer.com/article/10.1007/BF00567747.

Lauth:1991:TER

[152] Bernhard Lauth. Theory evolution and reference kinematics. Synthese, 88 (3):279–307, September 1991. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/BF00413550.

Barnes:1991:BVL

[153] Eric Barnes. Beyond verisimilitude: A linguistically invariant basis for scientific progress. Synthese, 88(3):309–339, September 1991. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://

link.springer.com/article/10.1007/BF00413551.

Pereboom:1991:WSR

[154] Derk Pereboom. Why a scientific realist cannot be a functionalist. Synthese, 88(3):341–358, September 1991. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/BF00413552.

Miller:1991:SNN

[155] Steven I. Miller and Marcel Fredericks. Some notes on the nature of methodological indeterminacy. Synthese, 88(3):359–378, September 1991. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/BF00413553.

Moore:1991:KSP

[156] James A. Moore. Knowledge, society, power, and the promise of epistemological externalism. Synthese, 88(3):379–398, September 1991. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://

link.springer.com/article/10.1007/BF00413554.

Cole:1991:AIP

[157] David Cole. Artificial intelligence and personal identity. Synthese, 88(3): 399–417, September 1991. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/BF00413555.

Anonymous:1991:Ad

[158] Anonymous. Announcement. Synthese, 88(3):418, September 1991. CO-DEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/BF00413556.

Kaiser:1991:P

[159] Matthias Kaiser. Preface. Synthese, 89(1):1–2, October 1991. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://

link.springer.com/article/10.1007/BF00413795.

Cartwright:1991:CWR

[160] Nancy Cartwright. Can wholism reconcile the inaccuracy of theory with the accuracy of prediction? Synthese, 89(1):3–13, October 1991. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://

link.springer.com/article/10.1007/BF00413796.

Wilkes:1991:RBS

[161] Kathleen V. Wilkes. The relationship between scientific psychology and common-sense psychology. Synthese, 89(1):15–39, October 1991. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://

link.springer.com/article/10.1007/BF00413797.

Aaberg:1991:RRM

[162] Claes Åberg. Realism and relativism in modern physics. Synthese, 89 (1):41–61, October 1991. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/BF00413798.

Andersen:1991:IBT

[163] Daniel Andersen. Interaction between theory and practice in the surgical treatment of ulcer disease in the period of 1880–1920. Synthese, 89(1): 63–73, October 1991. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF00413799.

Lie:1991:PTC

[164] Reidar K. Lie. Patterns of theory change in biomedicine: A case study from cardiology. Synthese, 89(1):75–88, October 1991. CODEN SYN-TAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.

springer.com/article/10.1007/BF00413800.

Needham:1991:DCT

[165] Paul Needham. Duhem and Cartwright on the truth of laws. Synthese, 89(1):89–109, October 1991. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/BF00413801.

Kaiser:1991:RGS

[166] Matthias Kaiser. From rocks to graphs- the shaping of phenomena. Synthese, 89(1):111–133, October 1991. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/BF00413802.

Niiniluoto:1991:RRC

[167] Ilkka Niiniluoto. Realism, relativism, and constructivism. Synthese, 89 (1):135–162, October 1991. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/BF00413803.

Sintonen:1991:HET

[168] Matti Sintonen. How evolutionary theory faces the reality. Synthese, 89 (1):163–183, October 1991. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/BF00413804.

Batterman:1991:CQC

[169] Robert W. Batterman. Chaos, quantization, and the correspondence principle. Synthese, 89(2):189–227, November 1991. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.

com/article/10.1007/BF00413905.

Kronz:1991:QEN

[170] Frederick M. Kronz. Quantum entanglement and nonideal measurements: A critique of Margenau’s objections to the projection postulate. Synthese, 89(2):229–251, November 1991. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/BF00413906.

Lavine:1991:QMA

[171] Shaughan Lavine. Is quantum mechanics an atomistic theory? Synthese, 89(2):253–271, November 1991. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/BF00413907.

VanKampen:1991:DP

[172] N. G. Van Kampen. Determinism and predictability. Synthese, 89(2): 273–281, November 1991. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/BF00413908.

Squires:1991:OMM

[173] Euan J. Squires. One mind or many — a note on the Everett interpretation of quantum theory. Synthese, 89(2):283–286, November 1991. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://

link.springer.com/article/10.1007/BF00413909.

Gisin:1991:PND

[174] N. Gisin. Propensities in a non-deterministic physics. Synthese, 89(2): 287–297, November 1991. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/BF00413910.

Baltas:1991:SSA

[175] Aristides Baltas. On some structural aspects of physical problems. Synthese, 89(2):299–320, November 1991. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/BF00413911.

Brown:1991:BR

[176] Curtis Brown and Steven Luper-Foy. Belief and rationality. Synthese, 89 (3):323–329, December 1991. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/BF00413500.

Bonjour:1991:TSP

[177] Laurence Bonjour. Is thought a symbolic process? Synthese, 89(3):331– 352, December 1991. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF00413501.

Brown:1991:BI

[178] Curtis Brown. Believing the impossible. Synthese, 89(3):353–364, December 1991. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF00413502.

Foley:1991:RBC

[179] Richard Foley. Rationality, belief and commitment. Synthese, 89(3):365– 392, December 1991. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF00413503.

Pears:1991:SDB

[180] David Pears. Self-deceptive belief-formation. Synthese, 89(3):393–405, December 1991. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF00413504.

Sainsbury:1991:CPE

[181] R. M. Sainsbury. Cartesian possibilities and the externality and extrinsicness of content. Synthese, 89(3):407–424, December 1991. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://

link.springer.com/article/10.1007/BF00413505.

Stalnaker:1991:PLO

[182] Robert Stalnaker. The problem of logical omniscience, I. Synthese, 89 (3):425–440, December 1991. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/BF00413506.

Stubenberg:1992:WIL

[183] Leopold Stubenberg. What is it like to be Oscar? Synthese, 90(1):1– 26, January 1992. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF00485190.

Koethe:1992:TAO

[184] John Koethe. And they ain’t outside the head either. Synthese, 90(1): 27–53, January 1992. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF00485191.

Shalkowski:1992:SCN

[185] Scott A. Shalkowski. Supervenience and causal necessity. Synthese, 90 (1):55–87, January 1992. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/BF00485192.

Owens:1992:PSE

[186] Joseph Owens. Psychophysical supervenience: Its epistemological foundation. Synthese, 90(1):89–117, January 1992. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.

com/article/10.1007/BF00485193.

Smith:1992:CA

[187] David Woodruff Smith. Consciousness in action. Synthese, 90(1):119– 143, January 1992. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF00485194.

Yu:1992:CML

[188] Qiuen Yu. Consistency, mechanicalness, and the logic of the mind. Synthese, 90(1):145–179, January 1992. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/BF00485195. See [298, 299].

Wolenski:1992:R

[189] Jan Woleński. Review. Synthese, 90(1):181–187, January 1992. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://

link.springer.com/article/10.1007/BF00485196.

Kyburg:1992:GFP

[190] Henry E. Kyburg, Jr. Getting fancy with probability. Synthese, 90(2): 189–203, February 1992. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/BF00485350.

Zabell:1992:PU

[191] S. L. Zabell. Predicting the unpredictable. Synthese, 90(2):205–232, February 1992. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF00485351.

Mayo:1992:DPR

[192] Deborah G. Mayo. Did Pearson reject the Neyman–Pearson philosophy of statistics? Synthese, 90(2):233–262, February 1992. CODEN SYN-TAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.

springer.com/article/10.1007/BF00485352.

Pollock:1992:TNP

[193] John L. Pollock. The theory of nomic probability. Synthese, 90(2):263– 299, February 1992. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF00485353.

Giles:1992:GTS

[194] Robin Giles. A generalization of the theory of subjective probability and expected utility. Synthese, 90(2):301–343, February 1992. CODEN SYN-TAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.

springer.com/article/10.1007/BF00485354.

Detlefsen:1992:PAL

[195] Michael Detlefsen. Poincaré against the logicians. Synthese, 90(3):349– 378, March 1992. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF00500033.

Cocchiarella:1992:CRV

[196] Nino B. Cocchiarella. Conceptual realism versus Quine on classes and higher-order logic. Synthese, 90(3):379–436, March 1992. CODEN SYN-TAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.

springer.com/article/10.1007/BF00500034.

Mitchell:1992:DIS

[197] Samuel William Mitchell. Dummett’s intuitionism is not strict finitism. Synthese, 90(3):437–458, March 1992. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/BF00500035.

Burian:1992:EIa

[198] Richard M. Burian and Marjorie Grene. Editorial introduction. Synthese, 91(1–2):1–7, January 1992. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/BF00484966.

Kolb:1992:KTE

[199] Daniel Kolb. Kant, teleology, and evolution. Synthese, 91(1–2):9– 28, January 1992. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF00484967.

Nuovo:1992:RP

[200] Victor Nuovo. Rethinking Paley. Synthese, 91(1–2):29–51, January 1992. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/BF00484968.

Nordmann:1992:DWB

[201] Alfred Nordmann. Darwinians at war: Bateson’s place in histories of Darwinism. Synthese, 91(1–2):53–72, January 1992. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.

com/article/10.1007/BF00484969.

Rinard:1992:HSC

[202] R. G. Rinard. Hans Spemann: Cultural factors in the rejection of an engineering stance in embryology. Synthese, 91(1–2):73–91, January 1992. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/BF00484970.

Atkinson:1992:CIR

[203] J. W. Atkinson. Conceptual issues in the reunion of development and evolution. Synthese, 91(1–2):93–110, January 1992. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.

com/article/10.1007/BF00484971.

Zylstra:1992:LTH

[204] Uko Zylstra. Living things as hierarchically organized structures. Synthese, 91(1–2):111–133, January 1992. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/BF00484972.

Williams:1992:CC

[205] Patricia A. Williams. Confusion in cladism. Synthese, 91(1–2):135– 152, January 1992. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF00484973.

Anonymous:1992:IUH

[206] Anonymous. International Union of History and Philosophy of Science Division of Logic, Methodology and Philosophy of Science (IUHPS/DLMPS) Bulletin No. 16. Synthese, 91(1–2):153–165, January 1992. CODEN SYN-TAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.

springer.com/article/10.1007/BF00484974.

Sarkar:1992:MRC

[207] Sahotra Sarkar. Models of reduction and categories of reductionism. Synthese, 91(3):167–194, June 1992. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/BF00413566.

Lindstrom:1992:BRE

[208] Sten Lindström and Wlodzimierz Rabinowicz. Belief revision, epistemic conditionals and the Ramsey test. Synthese, 91(3):195–237, June 1992. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/BF00413567.

Hansson:1992:DBC

[209] Sven Ove Hansson. In defense of base contraction. Synthese, 91(3): 239–245, June 1992. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF00413568.

Bacharach:1992:BIB

[210] Michael Bacharach. Backward induction and beliefs about oneself. Synthese, 91(3):247–284, June 1992. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/BF00413569.

Tuomela:1992:GB

[211] Raimo Tuomela. Group beliefs. Synthese, 91(3):285–318, June 1992. CO-DEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/BF00413570.

Gauthier:1992:RER

[212] David Gauthier. Review essay: The roots and roles of normative governance. Synthese, 91(3):319–335, June 1992. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.

com/article/10.1007/BF00413571.

Williams:1992:RET

[213] Peter M. Williams. Review essay: A theory of logical frequentism. Synthese, 91(3):337–346, June 1992. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/BF00413572.

Anonymous:1992:Aa

[214] Anonymous. Announcement. Synthese, 91(3):347, June 1992. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://

link.springer.com/article/10.1007/BF00413573.

Burian:1992:EIb

[215] Richard M. Burian and Joseph C. Pitt. Editorial introduction. Synthese, 92(1):3–7, July 1992. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF00413738.

Reed:1992:KTA

[216] Edward S. Reed. Knowers talking about the known: Ecological realism as a philosophy of science. Synthese, 92(1):9–23, July 1992. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://

link.springer.com/article/10.1007/BF00413739.

Miller:1992:OGT

[217] Harlan B. Miller. Not the only game in town: Zoöepistemology and ontological pluralism. Synthese, 92(1):25–37, July 1992. CODEN SYN-TAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.

springer.com/article/10.1007/BF00413740.

Frede:1992:ACA

[218] Dorothea Frede. Accidental causes in Aristotle. Synthese, 92(1):39– 62, July 1992. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF00413741.

Brickhouse:1992:SEP

[219] Thomas C. Brickhouse and Nicholas D. Smith. Socrates’ elenctic psychology. Synthese, 92(1):63–82, July 1992. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.

com/article/10.1007/BF00413742.

Williams:1992:HSB

[220] William H. Williams. Is Hume’s shade of blue a red herring? Synthese, 92 (1):83–99, July 1992. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF00413743.

Ariew:1992:DTK

[221] Roger Ariew. Descartes and the tree of knowledge. Synthese, 92(1): 101–116, July 1992. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF00413744.

Pitt:1992:PHP

[222] Joseph C. Pitt. Problematics in the history of philosophy. Synthese, 92(1):117–134, July 1992. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/BF00413745.

Eldredge:1992:MGT

[223] Niles Eldredge. Marjorie Grene, ‘Two Evolutionary Theories’ and modern evolutionary theory. Synthese, 92(1):135–149, July 1992. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://

link.springer.com/article/10.1007/BF00413746.

Burian:1992:HCE

[224] Richard M. Burian. How the choice of experimental organism matters: Biological practices and discipline boundaries. Synthese, 92(1):151–166, July 1992. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic).

URL http://link.springer.com/article/10.1007/BF00413747.

Ullmann-Margalit:1992:HTH

[225] Edna Ullmann-Margalit and Avishai Margalit. Holding true and holding as true. Synthese, 92(2):167–187, August 1992. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.

com/article/10.1007/BF00414298.

Woods:1992:VPM

[226] John Woods and Brent Hudak. Verdi is the Puccini of music. Synthese, 92(2):189–220, August 1992. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/BF00414299.

Horgan:1992:TMT

[227] Terence Horgan and Mark Timmons. Troubles on moral twin earth: Moral queerness revived. Synthese, 92(2):221–260, August 1992. CODEN SYN-TAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.

springer.com/article/10.1007/BF00414300.

McDonald:1992:SIR

[228] John McDonald. Is strong inference really superior to simple inference? Synthese, 92(2):261–282, August 1992. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/BF00414301.

Godfrey-Smith:1992:IA

[229] Peter Godfrey-Smith. Indication and adaptation. Synthese, 92(2):283– 312, August 1992. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF00414302.

Shimony:1992:TEN

[230] Abner Shimony. That there exists no greatest prime. Synthese, 92(3): 313–314, September 1992. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/BF00414285.

Corry:1992:NBC

[231] Leo Corry. Nicolas Bourbaki and the concept of mathematical structure. Synthese, 92(3):315–348, September 1992. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/BF00414286.

Holland:1992:AAM

[232] Robert A. Holland. Apriority and applied mathematics. Synthese, 92(3): 349–370, September 1992. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/BF00414287.

McLaughlin:1992:EUN

[233] William I. McLaughlin and Sylvia L. Miller. An epistemological use of nonstandard analysis to answer Zeno’s objections against motion. Synthese, 92 (3):371–384, September 1992. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/BF00414288.

Belnap:1992:BST

[234] Nuel Belnap. Branching space-time. Synthese, 92(3):385–434, September 1992. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic).

URL http://link.springer.com/article/10.1007/BF00414289.

Sarkar:1992:RCE

[235] Sahotra Sarkar. Rudolf Carnap, 1891–1970: The editor’s introduction. Synthese, 93(1–2):1–14, November 1992. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/BF00869419.

Friedman:1992:EA

[236] Michael Friedman. Epistemology in the Aufbau. Synthese, 93(1–2):15– 57, November 1992. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF00869420.

Richardson:1992:LIC

[237] Alan W. Richardson. Logical idealism and Carnap’s construction of the world. Synthese, 93(1–2):59–92, November 1992. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.

com/article/10.1007/BF00869421.

Webb:1992:RRR

[238] Judson Webb. Reconstruction from recollection and the refutation of idealism: A Kantian theme in the Aufbau. Synthese, 93(1–2):93–105, November 1992. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF00869422.

Uebel:1992:RRE

[239] Thomas E. Uebel. Rational reconstruction as elucidation? Carnap in the early protocol sentence debate. Synthese, 93(1–2):107–140, November 1992. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic).

URL http://link.springer.com/article/10.1007/BF00869423.

Creath:1992:CC

[240] Richard Creath. Carnap’s conventionalism. Synthese, 93(1–2):141–165, November 1992. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF00869424.

Hintikka:1992:CWF

[241] Jaakko Hintikka. Carnap’s work in the foundations of logic and mathematics in a historical perspective. Synthese, 93(1–2):167–189, November 1992. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic).

URL http://link.springer.com/article/10.1007/BF00869425.

Sarkar:1992:BOU

[242] Sahotra Sarkar. “The boundless ocean of unlimited possibilities”: Logic in Carnap’s logical syntax of language. Synthese, 93(1–2):191–237, November 1992. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic).

URL http://link.springer.com/article/10.1007/BF00869426.

Oberdan:1992:CTC

[243] Thomas Oberdan. The concept of truth in Carnap’s logical syntax of language. Synthese, 93(1–2):239–260, November 1992. CODEN SYN-TAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.

springer.com/article/10.1007/BF00869427.

Shimony:1992:CRM

[244] Abner Shimony. On Carnap: Reflections of a metaphysical student. Synthese, 93(1–2):261–274, November 1992. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/BF00869428.

Stein:1992:WCE

[245] Howard Stein. Was Carnap entirely wrong, after all? Synthese, 93(1–2): 275–295, November 1992. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/BF00869429.

Anonymous:1992:Ab

[246] Anonymous. Announcement. Synthese, 93(1–2):297, November 1992. CO-DEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/BF00869430.

Kuipers:1992:NRT

[247] Theo A. F. Kuipers. Naive and refined truth approximation. Synthese, 93 (3):299–341, December 1992. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/BF01089273.

Bonilla:1992:TTM

[248] Jesús P. Zamora Bonilla. Truthlikeness without truth: A methodological approach. Synthese, 93(3):343–372, December 1992. CODEN SYN-TAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.

springer.com/article/10.1007/BF01089274.

Reimer:1992:TVD

[249] Marga Reimer. Three views of demonstrative reference. Synthese, 93(3): 373–402, December 1992. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/BF01089275.

Yablo:1992:CE

[250] Stephen Yablo. Cause and essence. Synthese, 93(3):403–449, December 1992. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic).

URL http://link.springer.com/article/10.1007/BF01089276.

Anonymous:1992:Ac

[251] Anonymous. Announcement. Synthese, 93(3):450, December 1992. CO-DEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/BF01089277.

Elgin:1993:P

[252] Catherine Z. Elgin. Preface. Synthese, 94(1):1, January 1993. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://

link.springer.com/article/10.1007/BF01064567.

Elgin:1993:SS

[253] Catherine Z. Elgin. Scheffler’s symbols. Synthese, 94(1):3–12, January 1993. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic).

URL http://link.springer.com/article/10.1007/BF01064568.

Hanson:1993:RP

[254] Karen Hanson. Reconstruction in pragmatism. Synthese, 94(1):13– 23, January 1993. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF01064569.

Arnstine:1993:RDC

[255] Donald Arnstine and Barbara Arnstine. Rationality and democracy: A critical appreciation of Israel Scheffler’s philosophy of education. Synthese, 94(1):25–41, January 1993. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/BF01064570.

Schwartz:1993:WS

[256] Robert Schwartz. On ‘what is said to be’. Synthese, 94(1):43–54, January 1993. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic).

URL http://link.springer.com/article/10.1007/BF01064571.

Neiman:1993:ORE

[257] Alven Neiman and Harvey Siegel. Objectivity and rationality in epistemology and education: Scheffler’s middle road. Synthese, 94(1):55– 83, January 1993. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF01064572.

Martin:1993:NPC

[258] Jane Roland Martin. The new problem of curriculum. Synthese, 94(1): 85–104, January 1993. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF01064573.

Adler:1993:RBU

[259] Jonathan E. Adler. Reasonableness, bias, and the untapped power of procedure. Synthese, 94(1):105–125, January 1993. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.

com/article/10.1007/BF01064574.

Scheffler:1993:R

[260] Israel Scheffler. Responses. Synthese, 94(1):127–137, January 1993. CO-DEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/BF01064575.

Anonymous:1993:WIS

[261] Anonymous. Writings of Israel scheffler. Synthese, 94(1):139–144, January 1993. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic).

URL http://link.springer.com/article/10.1007/BF01064576.

Boardman:1993:RPK

[262] William S. Boardman. The relativity of perceptual knowledge. Synthese, 94(2):145–169, February 1993. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/BF01064336.

Sober:1993:TOL

[263] Elliott Sober. Temporally oriented laws. Synthese, 94(2):171–189, February 1993. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF01064337.

Brennan:1993:HII

[264] Geoffrey Brennan and Philip Pettit. Hands invisible and intangible. Synthese, 94(2):191–225, February 1993. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/BF01064338.

Hausman:1993:WDE

[265] Daniel M. Hausman. Why don’t effects explain their causes? Synthese, 94 (2):227–244, February 1993. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/BF01064339.

Allport:1993:LPE

[266] P. P. Allport. Are the laws of physics ‘economical with the truth’? Synthese, 94(2):245–290, February 1993. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/BF01064340. See reply [267].

Cartwright:1993:NSN

[267] Nancy Cartwright. Is natural science ‘natural’ enough?: A reply to Philip Allport. Synthese, 94(2):291–301, February 1993. CODEN SYN-TAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.

springer.com/article/10.1007/BF01064341. See [266].

Rawling:1993:CCE

[268] Piers Rawling. Choice and conditional expected utility. Synthese, 94(2): 303–328, February 1993. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/BF01064342.

Anonymous:1993:A

[269] Anonymous. Announcement. Synthese, 94(2):329–333, February 1993. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/BF01064343.

Haack:1993:TFQ

[270] Susan Haack. The two faces of Quine’s naturalism. Synthese, 94(3): 335–356, March 1993. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF01064484.

Kornblith:1993:EN

[271] Hilary Kornblith. Epistemic normativity. Synthese, 94(3):357–376, March 1993. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic).

URL http://link.springer.com/article/10.1007/BF01064485.

Kroon:1993:REP

[272] Frederick Kroon. Rationality and epistemic paradox. Synthese, 94(3): 377–408, March 1993. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF01064486.

Grim:1993:OPK

[273] Patrick Grim. Operators in the paradox of the knower. Synthese, 94 (3):409–428, March 1993. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/BF01064487.

Stevenson:1993:WBW

[274] Leslie Stevenson. Why believe what people say? Synthese, 94(3):429– 451, March 1993. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF01064488.

Senor:1993:IFJ

[275] Thomas D. Senor. Internalistic foundationalism and the justification of memory belief. Synthese, 94(3):453–476, March 1993. CODEN SYN-TAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.

springer.com/article/10.1007/BF01064489.

Scholz:1993:IRC

[276] Oliver R. Scholz. Introduction: Reconceptions in context. Synthese, 95 (1):1–7, April 1993. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF01064663.

Goodman:1993:SWW

[277] Nelson Goodman. On some worldly worries. Synthese, 95(1):9–12, April 1993. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic).

URL http://link.springer.com/article/10.1007/BF01064664.

Elgin:1993:UAS

[278] Catherine Z. Elgin. Understanding: Art and science. Synthese, 95(1): 13–28, April 1993. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF01064665.

Kung:1993:OCS

[279] Guido Küng. Ontology and the construction of systems. Synthese, 95 (1):29–53, April 1993. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF01064666.

Rheinwald:1993:ESG

[280] Rosemarie Rheinwald. An epistemic solution to Goodman’s new riddle of induction. Synthese, 95(1):55–76, April 1993. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.

com/article/10.1007/BF01064667.

Heydrich:1993:RM

[281] Wolfgang Heydrich. A reconception of meaning. Synthese, 95(1):77– 94, April 1993. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF01064668.

Scholz:1993:WP

[282] Oliver R. Scholz. When is a picture? Synthese, 95(1):95–106, April 1993. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/BF01064669.

Kunne:1993:TRP

[283] Wolfgang Künne. Truth, rightness, and permanent acceptability. Synthese, 95(1):107–117, April 1993. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/BF01064670.

Koppelberg:1993:SWR

[284] Dirk Koppelberg. Should we replace knowledge by understanding? — a comment on Elgin and Goodman’s reconception of epistemology. Synthese, 95(1):119–128, April 1993. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/BF01064671.

Elgin:1993:OPR

[285] Catherine Z. Elgin. Outstanding problems: Replies to ZiF critics. Synthese, 95(1):129–140, April 1993. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/BF01064672.

Corbi:1993:CCM

[286] Josep E. Corbí. Classical and connectionist models: Levels of description. Synthese, 95(2):141–168, May 1993. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/BF01064585.

Ben-Yami:1993:NCR

[287] Hannóch Ben-Yami. A note on the Chinese room. Synthese, 95(2): 169–172, May 1993. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF01064586.

Copeland:1993:CCC

[288] B. J. Copeland. The curious case of the Chinese gym. Synthese, 95(2): 173–186, May 1993. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF01064587.

Kaye:1993:MOC

[289] Lawrence J. Kaye. Are most of our concepts innate? Synthese, 95(2): 187–217, May 1993. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF01064588.

Mormann:1993:NPT

[290] Thomas Mormann. Natural predicates and topological structures of conceptual spaces. Synthese, 95(2):219–240, May 1993. CODEN SYN-TAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.

springer.com/article/10.1007/BF01064589.

Cartwright:1993:TAI

[291] Helen Morris Cartwright. On two arguments for the indeterminacy of personal identity. Synthese, 95(2):241–273, May 1993. CODEN SYN-TAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.

springer.com/article/10.1007/BF01064590.

Marras:1993:PSN

[292] Ausonio Marras. Psychophysical supervenience and nonreductive materialism. Synthese, 95(2):275–304, May 1993. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.

com/article/10.1007/BF01064591.

Larson:1993:ILF

[293] Richard K. Larson and Peter Ludlow. Interpreted logical forms. Synthese, 95(3):305–355, June 1993. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/BF01063877.

Martens:1993:CER

[294] David B. Martens. Close enough to reference. Synthese, 95(3):357– 377, June 1993. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF01063878.

Stern:1993:SEC

[295] Cindy D. Stern. Semantic emphasis in causal sentences. Synthese, 95(3): 379–418, June 1993. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF01063879.

Hand:1993:DBQ

[296] Michael Hand. A defense of branching quantification. Synthese, 95(3): 419–432, June 1993. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF01063880.

Liston:1993:TMF

[297] Michael Liston. Taking mathematical fictions seriously. Synthese, 95(3): 433–458, June 1993. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF01063881.

Priest:1993:YYM

[298] Graham Priest. Yu and your mind. Synthese, 95(3):459–460, June 1993. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/BF01063882. See [188]

and reply [299].

Yu:1993:FEG

[299] Qiuen Yu. Further explanations of the Gödel scenario of the mind — a reply to Prof. Graham Priest. Synthese, 95(3):461–465, June 1993. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://

link.springer.com/article/10.1007/BF01063883. See [188, 298].

Knuuttila:1993:P

[300] Simo Knuuttila. Preface. Synthese, 96(1):1–2, July 1993. CODEN SYN-TAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.

springer.com/article/10.1007/BF01063799.

Langston:1993:SDI

[301] Douglas C. Langston. Scotus’s doctrine of intuitive cognition. Synthese, 96(1):3–24, July 1993. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF01063800.

Klima:1993:CRE

[302] Gyula Klima. The changing role of entia rationis in mediaeval semantics and ontology: A comparative study with a reconstruction. Synthese, 96 (1):25–58, July 1993. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF01063801.

Yrjonsuuri:1993:ATM

[303] Mikko Yrjönsuuri. Aristotle’s topics and medieval obligational disputations. Synthese, 96(1):59–82, July 1993. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/BF01063802.

Normore:1993:PAH

[304] Calvin G. Normore. Petrus aureoli and his contemporaries on future contingents and excluded middle. Synthese, 96(1):83–92, July 1993. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://

link.springer.com/article/10.1007/BF01063803.

Trifogli:1993:GRI

[305] Cecilia Trifogli. Giles of Rome on the instant of change. Synthese, 96 (1):93–114, July 1993. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF01063804.

Knuuttila:1993:CWC

[306] Simo Knuuttila and Taina Holopainen. Conditional will and conditional norms in medieval thought. Synthese, 96(1):115–132, July 1993. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://

link.springer.com/article/10.1007/BF01063805.

Saarinen:1993:JBD

[307] Risto Saarinen. John Buridan and Donald Davidson on akrasia. Synthese, 96(1):133–153, July 1993. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/BF01063806.

Knuuttila:1993:CEA

[308] Simo Knuuttila. The critical edition of Adam Wodeham’s Lectura secunda. Synthese, 96(1):155–159, July 1993. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.

com/article/10.1007/BF01063807.

Wians:1993:RE

[309] William Wians. Review essay. Synthese, 96(1):161–165, July 1993. CO-DEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/BF01063808.

Boutot:1993:CTC

[310] Alain Boutot. Catastrophe theory and its critics. Synthese, 96(2):167– 200, August 1993. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF01306896.

Cartwright:1993:PRE

[311] Helen Morris Cartwright. On plural reference and elementary set theory. Synthese, 96(2):201–254, August 1993. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/BF01306897.

McCarty:1993:FMP

[312] David Charles McCarty. On the failure of mathematics’ philosophy. Synthese, 96(2):255–291, August 1993. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/BF01306898.

Tait:1993:SRE

[313] William W. Tait. Some recent essays in the history of the philosophy of mathematics: A critical review. Synthese, 96(2):293–331, August 1993. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/BF01306899.

Posy:1993:EI

[314] Carl J. Posy and Michael T. Ferejohn. Editors’ introduction. Synthese, 96 (3):333–334, September 1993. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/BF01064005.

Smith:1993:AUD

[315] Robin Smith. Aristotle on the uses of dialectic. Synthese, 96(3):335– 358, September 1993. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF01064006.

Patterson:1993:APS

[316] Richard Patterson. Aristotle’s perfect syllogisms, predication, and the dictum de omni. Synthese, 96(3):359–378, September 1993. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http:

//link.springer.com/article/10.1007/BF01064007.

Gill:1993:MAS

[317] Mary Louise Gill. Matter against substance. Synthese, 96(3):379–397, September 1993. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF01064008.

Woods:1993:FSP

[318] Michael Woods. Form, species, and predication in Aristotle. Synthese, 96 (3):399–415, September 1993. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/BF01064009.

Matthen:1993:AUF

[319] Mohan Matthen and R. J. Hankinson. Aristotle’s universe: Its form and matter. Synthese, 96(3):417–435, September 1993. CODEN SYN-TAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.

springer.com/article/10.1007/BF01064010.

Normore:1993:NDC

[320] Calvin G. Normore. The necessity in deduction: Cartesian inference and its medieval background. Synthese, 96(3):437–454, September 1993. CO-DEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/BF01064011.

Buroker:1993:PRS

[321] Jill Vance Buroker. The port–Royal semantics of terms. Synthese, 96(3): 455–475, September 1993. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/BF01064012.

Kulstad:1993:TIP

[322] Mark A. Kulstad. Two interpretations of the pre-established harmony in the philosophy of Leibniz. Synthese, 96(3):477–504, September 1993. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/BF01064013.

Norton:1993:DTE

[323] John D. Norton. The determination of theory by evidence: The case for quantum discontinuity, 1900–1915. Synthese, 97(1):1–31, October 1993. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/BF01255831.

Cao:1993:CFP

[324] Tian Yu Cao and Silvan S. Schweber. The conceptual foundations and the philosophical aspects of renormalization theory. Synthese, 97(1):33– 108, October 1993. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF01255832.

Squires:1993:QTR

[325] Euan J. Squires. Quantum theory and the relation between the conscious mind and the physical world. Synthese, 97(1):109–123, October 1993. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/BF01255833.

Amann:1993:GPQ

[326] Anton Amann. The gestalt problem in quantum theory: Generation of molecular shape by the environment. Synthese, 97(1):125–156, October 1993. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic).

URL http://link.springer.com/article/10.1007/BF01255834.

McIntyre:1993:EIE

[327] Lee C. McIntyre. Editorial introduction: Empiricism in the philosophy of social science. Synthese, 97(2):159, November 1993. CODEN SYN-TAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.

springer.com/article/10.1007/BF01064112.

Rosenberg:1993:SIL

[328] Alex Rosenberg. Scientific innovation and the limits of social scientific prediction. Synthese, 97(2):161–181, November 1993. CODEN SYN-TAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.

springer.com/article/10.1007/BF01064113.

Little:1993:SLG

[329] Daniel Little. On the scope and limits of generalizations in the social sciences. Synthese, 97(2):183–207, November 1993. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.

com/article/10.1007/BF01064114.

McIntyre:1993:CSS

[330] Lee C. McIntyre. Complexity and social scientific laws. Synthese, 97(2): 209–227, November 1993. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/BF01064115.

Kincaid:1993:ENI

[331] Harold Kincaid. The empirical nature of the individualism-holism dispute. Synthese, 97(2):229–247, November 1993. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/BF01064116.

Salmon:1993:RSS

[332] Merrilee H. Salmon. Reasoning in the social sciences. Synthese, 97(2): 249–267, November 1993. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/BF01064117.

Martin:1993:GIA

[333] Michael Martin. Geertz and the interpretive approach in anthropology. Synthese, 97(2):269–286, November 1993. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/BF01064118.

Zangwill:1993:QRE

[334] Nick Zangwill. Quasi-realist explanation. Synthese, 97(3):287–296, December 1993. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF01064071.

Levin:1993:RI

[335] Michael Levin. Reliabilism and induction. Synthese, 97(3):297–334, December 1993. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF01064072.

Hitchcock:1993:GPT

[336] Christopher Read Hitchcock. A generalized probabilistic theory of causal relevance. Synthese, 97(3):335–364, December 1993. CODEN SYN-TAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.

springer.com/article/10.1007/BF01064073.

Wright:1993:MQT

[337] Edmond Wright. More qualia trouble for functionalism: The Smythies TV-hood analogy. Synthese, 97(3):365–382, December 1993. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://

link.springer.com/article/10.1007/BF01064074.

Gjelsvik:1994:P

[338] Olav Gjelsvik and Herman Ruge Jervell. Preface. Synthese, 98(1):1– 2, January 1994. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF01064022.

Dreyfus:1994:KNP

[339] Hubert Dreyfus and Jane Rubin. Kierkegaard on the nihilism of the present age: The case of commitment as addiction. Synthese, 98(1):3– 19, January 1994. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/

BF01064023. See comments [343].

Elster:1994:RES

[340] Jon Elster. Rationality, emotions, and social norms. Synthese, 98(1): 21–49, January 1994. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/

BF01064024. See comments [343].

Honderich:1994:ST

[341] Ted Honderich. Seeing things. Synthese, 98(1):51–71, January 1994. CO-DEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/BF01064025. See com-

ments [343].

Sprigge:1994:C

[342] T. L. S. Sprigge. Consciousness. Synthese, 98(1):73–93, January 1994. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/BF01064026. See com-

ments [343].

Hannay:1994:CHS

[343] Alastair Hannay. Comments on Honderich, Sprigge, Dreyfus and Rubin, and Elster. Synthese, 98(1):95–112, January 1994. CODEN SYN-TAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.

springer.com/article/10.1007/BF01064027. See [341, 342, 339, 340].

Hintikka:1994:WQ

[344] Jaakko Hintikka and Gabriel Sandu. What is a quantifier? Synthese, 98 (1):113–129, January 1994. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.

1007/BF01064028. See comments [348].

Prawitz:1994:ME

[345] Dag Prawitz. Meaning and experience. Synthese, 98(1):131–141, January 1994. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic).

URL http://link.springer.com/article/10.1007/BF01064029. See

comments [348].

Quine:1994:PE

[346] W. V. Quine. Promoting extensionality. Synthese, 98(1):143–151, January 1994. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic).

URL http://link.springer.com/article/10.1007/BF01064030. See

comments [348].

Smith:1994:HHQ

[347] David Woodruff Smith. How to Husserl a Quine — and a Heidegger, too. Synthese, 98(1):153–173, January 1994. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.

com/article/10.1007/BF01064031. See comments [348].

Follesdal:1994:CQP

[348] Dagfinn Føllesdal. Comments on Quine, Prawitz, Hintikka and Sandu, and Smith. Synthese, 98(1):175–186, January 1994. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.

com/article/10.1007/BF01064032. See [345, 346, 344, 347].

Klima:1994:CRE

[349] G. Klima. The changing role of entia rationis in mediaeval semantics and ontology. Synthese, 98(1):187, January 1994. CODEN SYN-TAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://

link.springer.com/article/10.1007/BF01064033;   http://link.
springer.com/content/pdf/10.1007/BF01064033.pdf.

Boer:1994:PAF

[350] Steven E. Boër. Propositional attitudes and formal ontology. Synthese, 98 (2):187–242, February 1994. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/BF01063942.

Wallis:1994:TRP

[351] Charles Wallis. Truth-ratios, process, task, and knowledge. Synthese, 98 (2):243–269, February 1994. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/BF01063943.

Day:1994:PIB

[352] Timothy Day and Harold Kincaid. Putting inference to the best explanation in its place. Synthese, 98(2):271–295, February 1994. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://

link.springer.com/article/10.1007/BF01063944.

Green:1994:RRV

[353] Mitchell S. Green and Christopher R. Hitchcock. Reflections on reflection: Van Fraassen on belief. Synthese, 98(2):297–324, February 1994. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://

link.springer.com/article/10.1007/BF01063945.

Kvanvig:1994:CVF

[354] Jonathan L. Kvanvig. A critique of Van Fraassen’s voluntaristic epistemology. Synthese, 98(2):325–348, February 1994. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.

com/article/10.1007/BF01063946.

Maury:1994:SRW

[355] André Maury. Sources of the remarks in Wittgenstein’s philosophical investigations. Synthese, 98(3):349–378, March 1994. CODEN SYN-TAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.

springer.com/article/10.1007/BF01063925.

Falk:1994:DWO

[356] Barrie Falk. Doing what one meant to do. Synthese, 98(3):379–399, March 1994. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic).

URL http://link.springer.com/article/10.1007/BF01063926.

Ruffino:1994:CPW

[357] Marco Antonio Ruffino. The context principle and Wittgenstein’s criticism of Russell’s theory of types. Synthese, 98(3):401–414, March 1994. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/BF01063927.

Stern:1994:RWW

[358] David G. Stern. Recent work on Wittgenstein, 1980–1990. Synthese, 98 (3):415–458, March 1994. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/BF01063928.

Anonymous:1994:TII

[359] Anonymous. Third ISSA International Conference on Argumentation Amsterdam, 21–24 June 1994. Synthese, 98(3):459–463, March 1994. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://

link.springer.com/article/10.1007/BF01063929.

Pitt:1994:EP

[360] Joseph C. Pitt. Editorial preface. Synthese, 99(1):1, April 1994. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://

link.springer.com/article/10.1007/BF01064527.

Norton:1994:SC

[361] John D. Norton. Science and certainty. Synthese, 99(1):3–22, April 1994. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/BF01064528.

Laymon:1994:DIO

[362] Ronald Laymon. Demonstrative induction, old and new evidence and the accuracy of the electrostatic inverse square law. Synthese, 99(1):

23–58, April 1994. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF01064529.

Brandon:1994:TEE

[363] Robert N. Brandon. Theory and experiment in evolutionary biology. Synthese, 99(1):59–73, April 1994. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/BF01064530.

Burian:1994:CRB

[364] Richard M. Burian. Comments on Robert Brandon’s ‘theory and experiment in evolutionary biology’. Synthese, 99(1):75–86, April 1994. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://

link.springer.com/article/10.1007/BF01064531.

Chen:1994:RRA

[365] Xiang Chen. The rule of reproducibility and its applications in experiment appraisal. Synthese, 99(1):87–109, April 1994. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.

com/article/10.1007/BF01064532.

Schagrin:1994:MHT

[366] Morton L. Schagrin. More heat than light: Rumford’s experiments on the materiality of light. Synthese, 99(1):111–121, April 1994. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://

link.springer.com/article/10.1007/BF01064533.

Ackermann:1994:RE

[367] Robert Ackermann. Reporting experiments. Synthese, 99(1):123– 135, April 1994. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF01064534.

Stein:1994:RRE

[368] Edward Stein. Rationality and reflective equilibrium. Synthese, 99(2): 137–172, May 1994. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF01064426.

Cohen:1994:RS

[369] L. Jonathan Cohen. A reply to Stein. Synthese, 99(2):173–176, May 1994. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/BF01064427.

Stein:1994:CCR

[370] Edward Stein. Cordoning competence: A reply to Cohen. Synthese, 99(2): 177–179, May 1994. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF01064428.

Hooker:1994:INR

[371] C. A. Hooker. Idealisation, naturalism, and rationality: Some lessons from minimal rationality. Synthese, 99(2):181–231, May 1994. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://

link.springer.com/article/10.1007/BF01064429.

Malmnas:1994:AJU

[372] Per erik Malmnäs. Axiomatic justifications of the utility principle: A formal investigation. Synthese, 99(2):233–249, May 1994. CODEN SYN-TAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.

springer.com/article/10.1007/BF01064430.

Tilley:1994:ATP

[373] John Tilley. Accounting for the ‘tragedy’ in the Prisoner’s dilemma. Synthese, 99(2):251–276, May 1994. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/BF01064431.

Fisch:1994:TRT

[374] Menachem Fisch. Toward a rational theory of progress. Synthese, 99(2): 277–304, May 1994. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF01064432.

Slater:1994:GKR

[375] B. H. Slater. Getting Kant right. Synthese, 99(2):305–306, May 1994. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/BF01064433.

Jacquette:1994:OUC

[376] Dale Jacquette. Obligations under causal constraints. Synthese, 99(2): 307–310, May 1994. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF01064434.

Blinov:1994:SGC

[377] Arcady Blinov. Semantic games with chance moves. Synthese, 99(3): 311–327, June 1994. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF01063991.

Harris:1994:GIT

[378] Stephen Harris. Gts and interrogative tableaux. Synthese, 99(3):329– 343, June 1994. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF01063992.

Sandu:1994:SAN

[379] Gabriel Sandu. Some aspects of negation in English. Synthese, 99(3): 345–360, June 1994. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF01063993.

Burch:1994:GTS

[380] Robert W. Burch. Game-theoretical semantics for Peirce’s existential graphs. Synthese, 99(3):361–375, June 1994. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.

com/article/10.1007/BF01063994.

Carlson:1994:LDG

[381] Lauri Carlson. Logic for dialogue games. Synthese, 99(3):377–415, June 1994. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic).

URL http://link.springer.com/article/10.1007/BF01063995.

Hand:1994:IV

[382] Michael Hand. Informational variability. Synthese, 99(3):417–456, June 1994. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic).

URL http://link.springer.com/article/10.1007/BF01063996.

Burgin:1994:SPQ

[383] Mark Burgin and Vladimir Kuznetsov. Scientific problems and questions from a logical point of view. Synthese, 100(1):1–28, February 1994. CO-DEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/BF01063918.

Ishmuratov:1994:QSA

[384] Anatoliy Ishmuratov. Quest schemes in analytical models of discourse. Synthese, 100(1):29–38, February 1994. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/BF01063919.

Vasylchenko:1994:IAA

[385] Andriy Vasylchenko. Interpreting action as an answer. Synthese, 100 (1):39–48, February 1994. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/BF01063920.

Milne:1994:CHR

[386] Peter Milne. Classical harmony: Rules of inference and the meaning of the logical constants. Synthese, 100(1):49–94, February 1994. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://

link.springer.com/article/10.1007/BF01063921.

Walton:1994:BQP

[387] Douglas N. Walton. Begging the question as a pragmatic fallacy. Synthese, 100(1):95–131, February 1994. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/BF01063922.

Kvart:1994:CAT

[388] Igal Kvart. Counterfactuals: Ambiguities, true premises, and knowledge. Synthese, 100(1):133–164, February 1994. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/BF01063923.

Anonymous:1994:A

[389] Anonymous. Announcement. Synthese, 100(1):165–166, February 1994. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/BF01063924.

Preston:1994:BMT

[390] Beth Preston. Behaviorism and mentalism: Is there a third alternative? Synthese, 100(2):167–196, August 1994. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/BF01063809.

McGilvray:1994:CCH

[391] James A. McGilvray. Constant colors in the head. Synthese, 100(2): 197–239, August 1994. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF01063810.

Maloney:1994:CCC

[392] J. Christopher Maloney. Content: Covariation, control and contingency. Synthese, 100(2):241–290, August 1994. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/BF01063811.

Garson:1994:CCA

[393] James W. Garson. Cognition without classical architecture. Synthese, 100 (2):291–305, August 1994. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/BF01063812.

Hogan:1994:WWA

[394] Melinda Hogan. What is wrong with an atomistic account of mental representation? Synthese, 100(2):307–327, August 1994. CODEN SYN-TAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.

springer.com/article/10.1007/BF01063813.

Dacey:1994:I

[395] Raymond Dacey. Introduction. Synthese, 100(3):329–332, September 1994. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic).

URL http://link.springer.com/article/10.1007/BF01063906.

Lee:1994:FME

[396] S. C. Lee, R. G. Muncaster, and D. A. Zinnes. ‘The friend of my enemy is my enemy’: Modeling triadic internation relationships. Synthese, 100(3): 333–358, September 1994. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/BF01063907.

Saperstein:1994:MME

[397] Alvin M. Saperstein. Mathematical modeling of the effects of ‘capability’ and ‘intent’ on the stability of a competitive international system. Synthese, 100(3):359–378, September 1994. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.

com/article/10.1007/BF01063908.

Kilgour:1994:URP

[398] D. M. Kilgour and F. C. Zagare. Uncertainty and the role of the pawn in extended deterrence. Synthese, 100(3):379–412, September 1994. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://

link.springer.com/article/10.1007/BF01063909.

Luterbacher:1994:ICP

[399] Urs Luterbacher. International cooperation: The problem of the commons and the special case of the Antarctic region. Synthese, 100(3): 413–440, September 1994. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/BF01063910.

Mintz:1994:MMF

[400] Alex Mintz, Nehemia Geva, and Karl Derouen, Jr. Mathematical models of foreign policy decision-making: Compensatory vs. noncompensatory. Synthese, 100(3):441–460, September 1994. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.

com/article/10.1007/BF01063911.

Seitz:1994:AOS

[401] Steven Thomas Seitz. Apollo’s oracle: Strategizing for peace. Synthese, 100(3):461–495, September 1994. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/BF01063912.

Dacey:1994:IFT

[402] Raymond Dacey. Inducing fair trade out of hegemonic trade. Synthese, 100 (3):497–504, September 1994. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/BF01063913.

Gauthier:1994:HIL

[403] Yvon Gauthier. Hilbert and the internal logic of mathematics. Synthese, 101(1):1–14, October 1994. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/BF01063966.

Boos:1994:WFA

[404] William Boos. The world, the flesh and the argument from design. Synthese, 101(1):15–52, October 1994. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/BF01063967.

Hugly:1994:QR

[405] Philip Hugly and Charles Sayward. Quantifying over the reals. Synthese, 101(1):53–64, October 1994. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/BF01063968.

Schurz:1994:OTS

[406] Gerhard Schurz and Karel Lambert. Outline of a theory of scientific understanding. Synthese, 101(1):65–120, October 1994. CODEN SYN-TAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.

springer.com/article/10.1007/BF01063969.

Wolenski:1994:RE

[407] Jan Wolenski and Gabriel Sandu. Review essay. Synthese, 101(1):121– 127, October 1994. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF01063970.

Henderson:1994:AML

[408] David K. Henderson. Accounting for macro-level causation. Synthese, 101 (2):129–156, November 1994. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/BF01064014.

Fuhrmann:1994:URT

[409] André Fuhrmann and Isaac Levi. Undercutting and the Ramsey test for conditionals. Synthese, 101(2):157–169, November 1994. CODEN SYN-TAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.

springer.com/article/10.1007/BF01064015.

Cohen:1994:SST

[410] L. Jonathan Cohen. Some steps towards a general theory of relevance. Synthese, 101(2):171–185, November 1994. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/BF01064016.

Leeds:1994:CE

[411] Stephen Leeds. Constructive empiricism. Synthese, 101(2):187–221, November 1994. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF01064017.

Dilworth:1994:PLT

[412] Craig Dilworth. Principles, laws, theories and the metaphysics of science. Synthese, 101(2):223–247, November 1994. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/BF01064018.

Franssen:1994:CMR

[413] Maarten Franssen. Constrained maximization reconsidered- an elaboration and critique of Gauthier’s modelling of rational cooperation in a single prisoner’s dilemma. Synthese, 101(2):249–272, November 1994. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://

link.springer.com/article/10.1007/BF01064019.

Hales:1994:SDB

[414] Steven D. Hales. Self-deception and belief attribution. Synthese, 101(2): 273–289, November 1994. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/BF01064020.

Niiniluoto:1994:TMR

[415] Ilkka Niiniluoto. Truthlikeness misapplied: A reply to Ernest W. Adams. Synthese, 101(2):291–300, November 1994. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/BF01064021. See [57].

Clark:1994:IRC

[416] Andy Clark. Introduction: Reinventing the connectionist challenge. Synthese, 101(3):301–303, December 1994. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/BF01063892.

Horgan:1994:NFC

[417] Terence Horgan and John Tienson. A nonclassical framework for cognitive science. Synthese, 101(3):305–345, December 1994. CODEN SYN-TAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.

springer.com/article/10.1007/BF01063893.

Matthews:1994:TCM

[418] Robert J. Matthews. Three-concept monte: Explanation, implementation and systematicity. Synthese, 101(3):347–363, December 1994. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://

link.springer.com/article/10.1007/BF01063894.

McLaughlin:1994:ACR

[419] B. P. McLaughlin and T. A. Warfield. The allure of connectionism reexamined. Synthese, 101(3):365–400, December 1994. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.

com/article/10.1007/BF01063895.

Clark:1994:DR

[420] Andy Clark and Josefa Toribio. Doing without representing? Synthese, 101(3):401–431, December 1994. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/BF01063896.

Bechtel:1994:NDC

[421] William Bechtel. Natural deduction in connectionist systems. Synthese, 101(3):433–463, December 1994. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/BF01063897.

Aizawa:1994:RRC

[422] Kenneth Aizawa. Representations without rules, connectionism and the syntactic argument. Synthese, 101(3):465–492, December 1994. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://

link.springer.com/article/10.1007/BF01063898.

Groenewold:1995:FP

[423] Hip Groenewold. Field or print. Synthese, 102(1):1–59, January 1995. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/BF01063899.

Karakostas:1995:DUF

[424] Vassilios Karakostas and Michael Dickson. Decoherence in unorthodox formulations of quantum mechanics. Synthese, 102(1):61–97, January 1995. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/BF01063900.

Hawthorne:1995:WBA

[425] James Hawthorne and Michael Silberstein. For whom the Bell arguments toll. Synthese, 102(1):99–138, January 1995. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.

com/article/10.1007/BF01063901.

Bedford:1995:BTI

[426] Donald Bedford and Henry P. Stapp. Bell’s Theorem in an indeterministic universe. Synthese, 102(1):139–164, January 1995. CODEN SYN-TAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.

springer.com/article/10.1007/BF01063902.

Scerri:1995:EPC

[427] Eric R. Scerri. The exclusion principle, chemistry and hidden variables. Synthese, 102(1):165–169, January 1995. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/BF01063903.

Huggett:1995:RGE

[428] Nick Huggett and Robert Weingard. The renormalisation group and effective field theories. Synthese, 102(1):171–194, January 1995. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://

link.springer.com/article/10.1007/BF01063904.

Krause:1995:FFQ

[429] Décio Krause and Steven French. A formal framework for quantum nonindividuality. Synthese, 102(1):195–214, January 1995. CODEN SYN-TAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.

springer.com/article/10.1007/BF01063905.

Wang:1995:TPP

[430] Hao Wang. Time in philosophy and in physics: From Kant and Einstein to Gödel. Synthese, 102(2):215–234, February 1995. CODEN SYN-TAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.

springer.com/article/10.1007/BF01089801.

Saunders:1995:TQM

[431] Simon Saunders. Time, quantum mechanics, and decoherence. Synthese, 102(2):235–266, February 1995. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/BF01089802.

Ishigaki:1995:FSC

[432] Toshio Ishigaki. A formal system for classical particle mechanics, its model-theoretic applications and space-time structure. Synthese, 102(2): 267–292, February 1995. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/BF01089803.

deMuynck:1995:MIQ

[433] W. M. de Muynck. Measurement and the interpretation of quantum mechanics and relativity theory. Synthese, 102(2):293–318, February 1995. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/BF01089804.

Kuipers:1995:E

[434] Annie Kuipers and Jaakko Hintikka. Editorial. Synthese, 102(3):i, March 1995. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic).

URL http://link.springer.com/article/10.1007/BF01064119.

Tappenden:1995:GGF

[435] Jamie Tappenden. Geometry and generality in Frege’s philosophy of arithmetic. Synthese, 102(3):319–361, March 1995. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.

com/article/10.1007/BF01064120.

Weiner:1995:RFR

[436] Joan Weiner. Realism bei Frege: Reply to Burge. Synthese, 102(3):363– 382, March 1995. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF01064121.

Makin:1995:MSD

[437] Gideon Makin. Making sense of ‘on denoting’. Synthese, 102(3):383– 412, March 1995. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF01064122.

Minar:1995:FHL

[438] Edward H. Minar. Feeling at home in language. Synthese, 102(3):413– 452, March 1995. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF01064123.

Anonymous:1995:A

[439] Anonymous. Announcements. Synthese, 102(3):453–454, March 1995. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/BF01064124.

Pereboom:1995:SUK

[440] Derk Pereboom. Self-understanding in Kant’s transcendental deduction. Synthese, 103(1):1–42, April 1995. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/BF01063717.

Westphal:1995:DKM

[441] Kenneth R. Westphal. Does Kant’s metaphysical foundations of natural science fill a gap in the critique of pure reason? Synthese, 103(1):

43–86, April 1995. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF01063718.

Uebel:1995:ONI

[442] Thomas E. Uebel. Otto Neurath’s idealist inheritance. Synthese, 103(1): 87–121, April 1995. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF01063719.

Devidi:1995:TMC

[443] David Devidi and Graham Solomon. Tolerance and metalanguages in Carnap’s logical syntax of language. Synthese, 103(1):123–139, April 1995. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/BF01063720.

Folina:1995:PRT

[444] Janet Folina. Putnam, realism and truth. Synthese, 103(2):141–152, May 1995. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic).

URL http://link.springer.com/article/10.1007/BF01090045.

Cross:1995:PEC

[445] Charles B. Cross. Probability, evidence, and the coherence of the whole truth. Synthese, 103(2):153–170, May 1995. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.

com/article/10.1007/BF01090046.

Batterman:1995:TBT

[446] Robert W. Batterman. Theories between theories: Asymptotic limiting intertheoretic relations. Synthese, 103(2):171–201, May 1995. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://

link.springer.com/article/10.1007/BF01090047.

Mormann:1995:IEE

[447] Thomas Mormann. Incompatible empirically equivalent theories: A structural explication. Synthese, 103(2):203–249, May 1995. CODEN SYN-TAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.

springer.com/article/10.1007/BF01090048.

Barnes:1995:ILE

[448] Eric Barnes. Inference to the loveliest explanation. Synthese, 103(2): 251–277, May 1995. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF01090049.

Wright:1995:CS

[449] Crispin Wright. Critical study. Synthese, 103(2):279–302, May 1995. CO-DEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/BF01090050.

Balaguer:1995:PE

[450] Mark Balaguer. A platonist epistemology. Synthese, 103(3):303–325, June 1995. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic).

URL http://link.springer.com/article/10.1007/BF01089731.

Forrest:1995:STD

[451] Peter Forrest. Is space-time discrete or continuous?- an empirical question. Synthese, 103(3):327–354, June 1995. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/BF01089732.

Clifton:1995:DOB

[452] Rob Clifton and Mark Hogarth. The definability of objective becoming in Minkowski spacetime. Synthese, 103(3):355–387, June 1995. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://

link.springer.com/article/10.1007/BF01089733.

Lavine:1995:FM

[453] Shaughan Lavine. Finite mathematics. Synthese, 103(3):389–420, June 1995. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic).

URL http://link.springer.com/article/10.1007/BF01089734.

Marquis:1995:CTF

[454] Jean-Pierre Marquis. Category theory and the foundations of mathematics: Philosophical excavations. Synthese, 103(3):421–447, June 1995. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/BF01089735.

Thompson:1995:CVE

[455] Evan Thompson. Colour vision, evolution, and perceptual content. Synthese, 104(1):1–32, July 1995. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/BF01063672.

Chou:1995:TDS

[456] Ivans Chou and Lucia M. Vaina. Two-dimensional symmetric form discrimination: Fast learning, but not that fast. Synthese, 104(1):33–41, July 1995. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic).

URL http://link.springer.com/article/10.1007/BF01063673.

Vaina:1995:IOU

[457] Lucia M. Vaina, Harold Goodglass, and Lawren Daltroy. Inference of object use from pantomimed actions by aphasics and patients with right hemisphere lesions. Synthese, 104(1):43–57, July 1995. CODEN SYN-TAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.

springer.com/article/10.1007/BF01063674.

Leiber:1995:TTT

[458] Justin Leiber. On Turing’s Turing test and why the matter matters. Synthese, 104(1):59–69, July 1995. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/BF01063675.

Butler:1995:RCD

[459] Keith Butler. Representation and computation in a deflationary assessment of connectionist cognitive science. Synthese, 104(1):71–97, July 1995. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/BF01063676.

Macdonald:1995:EFP

[460] Cynthia Macdonald. Externalism and first-person authority. Synthese, 104(1):99–122, July 1995. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/BF01063677.

Horst:1995:EAB

[461] Steven Horst. Eliminativism and the ambiguity of ‘belief’. Synthese, 104(1):123–145, July 1995. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/BF01063678.

Alroy:1995:IL

[462] Daniel Alroy. Inner light. Synthese, 104(1):147–160, July 1995. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://

link.springer.com/article/10.1007/BF01063679.

Shafer:1995:RE

[463] Glenn Shafer. Review essay. Synthese, 104(1):161–176, July 1995. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://

link.springer.com/article/10.1007/BF01063680.

Fetzer:1995:EP

[464] James H. Fetzer and Paul W. Humphreys. Editorial preface. Synthese, 104(2):177, August 1995. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/BF01063868.

Smith:1995:MKO

[465] Quentin Smith. Marcus, Kripke, and the origin of the new theory of reference. Synthese, 104(2):179–189, August 1995. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.

com/article/10.1007/BF01063869. See reply [466].

Soames:1995:RAR

[466] Scott Soames. Revisionism about reference: A reply to Smith. Synthese, 104(2):191–216, August 1995. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.

1007/BF01063870. See [465] and reply [467].

Smith:1995:MNT

[467] Quentin Smith. Marcus and the new theory of reference: A reply to Scott Soames. Synthese, 104(2):217–244, August 1995. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.

com/article/10.1007/BF01063871. See [465, 466].

Hintikka:1995:FNT

[468] Jaakko Hintikka and Gabriel Sandu. The fallacies of the new theory of reference. Synthese, 104(2):245–283, August 1995. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.

com/article/10.1007/BF01063872.

Lichter:1995:BCF

[469] Tilman Lichter. Bill Clinton is the First Lady of the USA: Making and unmaking analogies. Synthese, 104(2):285–297, August 1995. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://

link.springer.com/article/10.1007/BF01063873.

Margalit:1995:MG

[470] Avishai Margalit and Naomi Goldblum. A metaphor game. Synthese, 104 (2):299–323, August 1995. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/BF01063874.

Boos:1995:WFA

[471] William Boos. The world, the flesh and the argument from design. Synthese, 104(2):325, August 1995. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/BF01063875;   http://link.springer.com/content/
pdf/10.1007/BF01063875.pdf.

Anonymous:1995:ACP

[472] Anonymous. Announcements and calls for papers. Synthese, 104(2):327– 328, August 1995. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF01063876.

Nelson:1995:P

[473] Lynn Hankinson Nelson. Preface. Synthese, 104(3):329–330, September 1995. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic).

URL http://link.springer.com/article/10.1007/BF01064503.

Harding:1995:SOR

[474] Sandra Harding. “Strong objectivity”: A response to the new objectivity question. Synthese, 104(3):331–349, September 1995. CODEN SYN-TAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.

springer.com/article/10.1007/BF01064504.

Lloyd:1995:ODS

[475] Elisabeth A. Lloyd. Objectivity and the double standard for feminist epistemologies. Synthese, 104(3):351–381, September 1995. CODEN SYN-TAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.

springer.com/article/10.1007/BF01064505.

Longino:1995:GPT

[476] Helen E. Longino. Gender, politics, and the theoretical virtues. Synthese, 104(3):383–397, September 1995. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/BF01064506.

Nelson:1995:FNP

[477] Lynn Hankinson Nelson. A feminist naturalized philosophy of science. Synthese, 104(3):399–421, September 1995. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.

com/article/10.1007/BF01064507.

Potter:1995:GSG

[478] Elizabeth Potter. Good science and good philosophy of science. Synthese, 104(3):423–439, September 1995. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/BF01064508.

Tuana:1995:VSE

[479] Nancy Tuana. The values of science: Empiricism from a feminist perspective. Synthese, 104(3):441–461, September 1995. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.

com/article/10.1007/BF01064509.

Anonymous:1995:R

[480] Anonymous. References. Synthese, 104(3):463–477, September 1995. CO-DEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/BF01064510.

Zalabardo:1995:PIT

[481] José L. Zalabardo. A problem for information theoretic semantics. Synthese, 105(1):1–29, October 1995. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/BF01064101.

Kemp:1995:TFL

[482] Gary Kemp. Truth in Frege’s ‘law of truth’. Synthese, 105(1):31– 51, October 1995. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF01064102.

Krysztofiak:1995:NTF

[483] Wojciech Krysztofiak. Noemata and their formalization. Synthese, 105 (1):53–86, October 1995. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/BF01064103.

Garrison:1995:DPE

[484] Jim Garrison. Dewey’s philosophy and the experience of working: Labor, tools and language. Synthese, 105(1):87–114, October 1995. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://

link.springer.com/article/10.1007/BF01064104.

Berge:1995:CTI

[485] William H. Berge. Carnap and translational indeterminacy. Synthese, 105 (1):115–121, October 1995. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/BF01064105.

Haaparanta:1995:PAP

[486] Leila Haaparanta. Perspectives into analytical philosophy. Synthese, 105 (1):123–139, October 1995. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/BF01064106.

Marion:1995:WF

[487] Mathieu Marion. Wittgenstein and finitism. Synthese, 105(2):141–176, November 1995. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF01064216.

Gert:1995:FRC

[488] Heather J. Gert. Family resemblances and criteria. Synthese, 105(2): 177–190, November 1995. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/BF01064217.

Gillett:1995:HDN

[489] Grant Gillett. Humpty Dumpty and the night of the Triffids: Individualism and rule-following. Synthese, 105(2):191–206, November 1995. CO-DEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/BF01064218.

Franchella:1995:LBW

[490] Miriam Franchella. Like a bee on a windowpane: Heyting’s reflections on solipsism. Synthese, 105(2):207–251, November 1995. CODEN SYN-TAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.

springer.com/article/10.1007/BF01064219.

Vaina:1995:AAB

[491] Lucia M. Vaina. Akinetopsia, achromatopsia and blindsight: Recent studies on perception without awareness. Synthese, 105(3):253–271, December 1995. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic).

URL http://link.springer.com/article/10.1007/BF01063559.

Charland:1995:FRC

[492] Louis C. Charland. Feeling and representing: Computational theory and the modularity of affect. Synthese, 105(3):273–301, December 1995. CO-DEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/BF01063560.

Hardcastle:1995:C

[493] Valerie Gray Hardcastle. Computationalism. Synthese, 105(3):303–317, December 1995. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF01063561.

Sandu:1995:JAG

[494] Gabriel Sandu and Raimo Tuomela. Joint action and group action made precise. Synthese, 105(3):319–345, December 1995. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.

springer.com/article/10.1007/BF01063562.

Bartels:1995:CMM

[495] Andreas Bartels. Chains of meaning: A model for concept formation in contemporary physics theories. Synthese, 105(3):347–379, December 1995. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/BF01063563.

Melnyk:1995:FPT

[496] Andrew Melnyk. Formulating physicalism: Two suggestions. Synthese, 105(3):381–407, December 1995. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/BF01063564.

Hintikka:1996:P

[497] Jaakko Hintikka and Dag Prawitz. Preface. Synthese, 106(1):1, January 1996. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic).

URL http://link.springer.com/article/10.1007/BF00413610.

Alanen:1996:RDN

[498] Lilli Alanen. Reconsidering Descartes’s notion of the mind-body union. Synthese, 106(1):3–20, January 1996. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/BF00413611.

Gardenfors:1996:MRC

[499] Peter Gärdenfors. Mental representation, conceptual spaces and metaphors. Synthese, 106(1):21–47, January 1996. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.

springer.com/article/10.1007/BF00413612.

Pears:1996:WCC

[500] David Pears. Wittgenstein’s criticism of Cartesianism. Synthese, 106 (1):49–55, January 1996. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/BF00413613.

Ricoeur:1996:CC

[501] Paul Ricoeur. The crisis of the Cogito. Synthese, 106(1):57–66, January 1996. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic).

URL http://link.springer.com/article/10.1007/BF00413614.

Zangwill:1996:GOS

[502] Nick Zangwill. Good old supervenience: Mental causation on the cheap. Synthese, 106(1):67–101, January 1996. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/BF00413615.

Divers:1996:SO

[503] John Divers. Supervenience for operators. Synthese, 106(1):103–112, January 1996. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF00413616.

Hatfield:1996:REI

[504] Gary Hatfield. Review essay: The importance of the history of science for philosophy in general. Synthese, 106(1):113–138, January 1996. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://

link.springer.com/article/10.1007/BF00413617.

Bar-On:1996:ARS

[505] Dorit Bar-On. Anti-realism and speaker knowledge. Synthese, 106(2): 139–166, February 1996. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/BF00413698.

Montgomery:1996:ICV

[506] Richard Montgomery. The indeterminacy of color vision. Synthese, 106 (2):167–203, February 1996. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/BF00413699.

Everett:1996:QV

[507] Anthony Everett. Qualia and vagueness. Synthese, 106(2):205–226, February 1996. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF00413700.

Bunzl:1996:LTE

[508] Martin Bunzl. The logic of thought experiments. Synthese, 106(2):227– 240, February 1996. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF00413701.

Weinstein:1996:U

[509] Steven Weinstein. Undermind. Synthese, 106(2):241–251, February 1996. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/BF00413702.

vanBrakel:1996:ISR

[510] J. van Brakel. Interdiscourse or supervenience relations: The primacy of the manifest image. Synthese, 106(2):253–297, February 1996. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://

link.springer.com/article/10.1007/BF00413703.

Anonymous:1996:EN

[511] Anonymous. Editorial note. Synthese, 106(3):299, March 1996. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://

link.springer.com/article/10.1007/BF00413588.

Brent:1996:PP

[512] Joseph Brent. Pursuing Peirce. Synthese, 106(3):301–322, March 1996. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/BF00413589.

Burks:1996:PEP

[513] Arthur W. Burks. Peirce’s evolutionary pragmatic idealism. Synthese, 106(3):323–372, March 1996. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/BF00413590.

Dipert:1996:RIR

[514] Randall R. Dipert. Reflections on iconicity, representation, and resemblance: Peirce’s theory of signs, Goodman on resemblance, and modern philosophies of language and mind. Synthese, 106(3):373–397, March 1996. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/BF00413591.

Schwartz:1996:ST

[515] Robert Schwartz. Symbols and thought. Synthese, 106(3):399–407, March 1996. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic).

URL http://link.springer.com/article/10.1007/BF00413592.

Short:1996:RE

[516] T. L. Short. Review essay. Synthese, 106(3):409–430, March 1996. CO-DEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/BF00413593.

Anonymous:1996:IUH

[517] Anonymous. International Union of the History and Philosophy of Science Division of Logic, Methodology and Philosophy of Science Bulletin No.

17. Synthese, 106(3):431–443, March 1996. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.

com/article/10.1007/BF00413594.

Anonymous:1996:CP

[518] Anonymous. Call for papers. Synthese, 106(3):445, March 1996. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://

link.springer.com/article/10.1007/BF00413595.

Breuer:1996:SDQ

[519] Thomas Breuer. Subjective decoherence in quantum measurements. Synthese, 107(1):1–17, April 1996. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/BF00413900.

Saunders:1996:TQM

[520] Simon Saunders. Time, quantum mechanics, and tense. Synthese, 107 (1):19–53, April 1996. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF00413901.

Dickson:1996:DLQ

[521] W. Michael Dickson. Determinism and locality in quantum systems. Synthese, 107(1):55–82, April 1996. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/BF00413902.

Boos:1996:MQT

[522] William Boos. Mathematical quantum theory I: Random ultrafilters as hidden variables. Synthese, 107(1):83–143, April 1996. CODEN SYN-TAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.

springer.com/article/10.1007/BF00413903.

Goldstein:1996:REB

[523] Sheldon Goldstein. Review essay: Bohmian mechanics and the quantum revolution. Synthese, 107(1):145–165, April 1996. CODEN SYN-TAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.

springer.com/article/10.1007/BF00413904.

Antonelli:1996:WF

[524] Gian Aldo Antonelli. What’s in a function? Synthese, 107(2):167– 204, May 1996. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF00413605.

Oaklander:1996:MPS

[525] L. Nathan Oaklander. McTaggart’s paradox and Smith’s tensed theory of time. Synthese, 107(2):205–221, May 1996. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.

com/article/10.1007/BF00413606.

Eilstein:1996:PST

[526] Helena Eilstein. Prof. Shimony on “the transient now”. Synthese, 107(2): 223–247, May 1996. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF00413607.

Craig:1996:NBT

[527] William Lane Craig. The new B-theory’s tu quoque argument. Synthese, 107(2):249–269, May 1996. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/BF00413608.

Harrison:1996:TAZ

[528] Craig Harrison. The three arrows of Zeno. Synthese, 107(2):271–292, May 1996. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic).

URL http://link.springer.com/article/10.1007/BF00413609.

Gilman:1996:OSC

[529] Daniel Gilman. Optimization and simplicity: Computational vision and biological explanation. Synthese, 107(3):293–323, June 1996. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://

link.springer.com/article/10.1007/BF00413839.

Gates:1996:PI

[530] Gary Gates. The price of information. Synthese, 107(3):325–347, June 1996. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic).

URL http://link.springer.com/article/10.1007/BF00413840.

Marmor:1996:C

[531] Andrei Marmor. On convention. Synthese, 107(3):349–371, June 1996. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/BF00413841.

Richmond:1996:STS

[532] Samuel A. Richmond. A simplification of the theory of simplicity. Synthese, 107(3):373–393, June 1996. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/BF00413842.

Hitchcock:1996:RCC

[533] Christopher Read Hitchcock. The role of contrast in causal and explanatory claims. Synthese, 107(3):395–419, June 1996. CODEN SYN-TAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.

springer.com/article/10.1007/BF00413843.

Kieseppa:1996:ATV

[534] I. A. Kieseppä. On the aim of the theory of verisimilitude. Synthese, 107(3):421–438, June 1996. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/BF00413844.

Cussens:1996:DIP

[535] James Cussens. Deduction, induction and probabilistic support. Synthese, 108(1):1–10, July 1996. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF00414003.

Sobel:1996:PW

[536] Jordan Howard Sobel. Pascalian wagers. Synthese, 108(1):11–61, July 1996. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic).

URL http://link.springer.com/article/10.1007/BF00414004.

Millgram:1996:DC

[537] Elijah Millgram and Paul Thagard. Deliberative coherence. Synthese, 108(1):63–88, July 1996. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/BF00414005.

Pust:1996:IFS

[538] Joel Pust. Induction, focused sampling and the law of small numbers. Synthese, 108(1):89–104, July 1996. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/BF00414006.

Mccurdy:1996:HPI

[539] Christopher S. I. Mccurdy. Humphrey’s paradox and the interpretation of inverse conditional propensities. Synthese, 108(1):105–125, July 1996. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/BF00414007.

Williams:1996:CPC

[540] David M. Williams, Robert W. Scotland, Christopher J. Humphries, and Darrell J. Siebert. Confusion in philosophy: A comment on Williams (1992). Synthese, 108(1):127–136, July 1996. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.

com/article/10.1007/BF00414008.

Harre:1996:OME

[541] Rom Harré. From observability to manipulability: Extending the inductive arguments for realism. Synthese, 108(2):137–155, August 1996. CO-DEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/BF00413494.

Kirk:1996:HPC

[542] Robert Kirk. How physicalists can avoid reductionism. Synthese, 108 (2):157–170, August 1996. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/BF00413495.

Petersen:1996:ETB

[543] Karl Petersen. Ergodic theorems and the basis of science. Synthese, 108 (2):171–183, August 1996. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/BF00413496.

Deltete:1996:EIT

[544] Robert J. Deltete and Reed A. Guy. Emerging from imaginary time. Synthese, 108(2):185–203, August 1996. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/BF00413497.

Humberstone:1996:IE

[545] I. L. Humberstone. Intrinsic/ extrinsic. Synthese, 108(2):205–267, August 1996. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic).

URL http://link.springer.com/article/10.1007/BF00413498.

Horty:1996:AO

[546] John F. Horty. Agency and obligation. Synthese, 108(2):269–307, August 1996. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic).

URL http://link.springer.com/article/10.1007/BF00413499.

Chalmers:1996:DRI

[547] David J. Chalmers. Does a rock implement every finite-state automaton? Synthese, 108(3):309–333, September 1996. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.

com/article/10.1007/BF00413692.

Copeland:1996:WC

[548] B. Jack Copeland. What is computation? Synthese, 108(3):335–359, September 1996. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF00413693.

Copeland:1996:ATA

[549] B. Jack Copeland and Diane Proudfoot. On Alan Turing’s anticipation of connectionism. Synthese, 108(3):361–377, September 1996. CODEN SYN-TAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.

springer.com/article/10.1007/BF00413694. See correction [811].

King:1996:HMT

[550] David King. Is the human mind a Turing machine? Synthese, 108(3): 379–389, September 1996. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/BF00413695.

Melnyk:1996:SAA

[551] Andrew Melnyk. Searle’s abstract argument against strong AI. Synthese, 108(3):391–419, September 1996. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/BF00413696.

Bartsch:1996:RBC

[552] Renate Bartsch. The relationship between connectionist models and a dynamic data-oriented theory of concept formation. Synthese, 108(3): 421–454, September 1996. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/BF00413697.

Wisniewski:1996:LQT

[553] Andrzej Wiśniewski. The logic of questions as a theory of erotetic arguments. Synthese, 109(1):1–25, October 1996. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.

com/article/10.1007/BF00413820.

Hammer:1996:TL

[554] Eric M. Hammer. The truths of logic. Synthese, 109(1):27–45, October 1996. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic).

URL http://link.springer.com/article/10.1007/BF00413821.

Jacquette:1996:VPM

[555] Dale Jacquette. The validity paradox in modal S_{5}. Synthese, 109(1): 47–62, October 1996. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF00413822.

Harrell:1996:CHS

[556] Mack Harrell. Confirmation holism and semantic holism. Synthese, 109 (1):63–101, October 1996. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/BF00413823.

Edwards:1996:ART

[557] Jim Edwards. Anti-realist truth and concepts of superassertibility. Synthese, 109(1):103–120, October 1996. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/BF00413824.

Ryan:1996:EVC

[558] Sharon Ryan. The epistemic virtues of consistency. Synthese, 109(2): 121–141, November 1996. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/BF00413765.

Nayak:1996:BCC

[559] Abhaya C. Nayak, Paul Nelson, and Hanan Polansky. Belief change as change in epistemic entrenchment. Synthese, 109(2):143–174, November 1996. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic).

URL http://link.springer.com/article/10.1007/BF00413766.

Zynda:1996:CIR

[560] Lyle Zynda. Coherence as an ideal of rationality. Synthese, 109(2):175– 216, November 1996. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF00413767.

Costa:1996:TNE

[561] Horacio Arló Costa and Isaac Levi. Two notions of epistemic validity. Synthese, 109(2):217–262, November 1996. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/BF00413768.

Rueger:1996:RDT

[562] Alexander Rueger. Risk and diversification in theory choice. Synthese, 109 (2):263–280, November 1996. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/BF00413769.

Scott:1996:TNT

[563] Brian M. Scott. Technical notes on a theory of simplicity. Synthese, 109 (2):281–289, November 1996. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/BF00413770.

Anonymous:1996:A

[564] Anonymous. Announcement. Synthese, 109(2):291, November 1996. CO-DEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/BF00413771.

Juhl:1996:ORS

[565] Cory F. Juhl. Objectively reliable subjective probabilities. Synthese, 109 (3):293–309, December 1996. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/BF00413863.

Sobel:1996:SCP

[566] Jordan Howard Sobel. On the significance of conditional probabilities. Synthese, 109(3):311–344, December 1996. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/BF00413864.

Mares:1996:RLT

[567] Edwin D. Mares. Relevant logic and the theory of information. Synthese, 109(3):345–360, December 1996. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/BF00413865.

Sensat:1996:RDE

[568] Julius Sensat. Reification as dependence on extrinsic information. Synthese, 109(3):361–399, December 1996. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/BF00413866.

Hansson:1996:DTF

[569] Sven Ove Hansson. Decision theoretic foundations for axioms of rational preference. Synthese, 109(3):401–412, December 1996. CODEN SYN-TAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.

springer.com/article/10.1007/BF00413867.

Howell:1996:RE

[570] Robert Howell. Review essay. Synthese, 109(3):413–434, December 1996. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/BF00413868.

Webb:1997:HFA

[571] Judson C. Webb. Hilbert’s formalism and arithmetization of mathematics. Synthese, 110(1):1–14, January 1997. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

content/pdf/10.1023/A%3A1004919107087.pdf.

Hintikka:1997:HV

[572] Jaakko Hintikka. Hilbert vindicated? Synthese, 110(1):15–36, January 1997. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964

(electronic). URL http://link.springer.com/content/pdf/10.1023/
A%3A1004915006179.pdf.

Majer:1997:HHC

[573] Ulrich Majer. Husserl and Hilbert on completeness. Synthese, 110(1):37– 56, January 1997. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964

(electronic). URL http://link.springer.com/content/pdf/10.1023/
A%3A1004962922108.pdf.

Ehrlich:1997:CAC

[574] Philip Ehrlich. From completeness to Archimedean completenes. Synthese, 110(1):57–76, January 1997. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/content/pdf/
10.1023/A%3A1004971123925.pdf.

Dreben:1997:HST

[575] Burton Dreben and Akihiro Kanamori. Hilbert and set theory. Synthese, 110(1):77–125, January 1997. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/content/pdf/
10.1023/A%3A1004908225146.pdf.

vonPlato:1997:FHG

[576] Jan von Plato. Formalization of Hilbert’s geometry of incidence and parallelism. Synthese, 110(1):127–141, January 1997. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.

com/content/pdf/10.1023/A%3A1004959405270.pdf.

Alper:1997:MMZ

[577] Joseph S. Alper and Mark Bridger. Mathematics, models and Zeno’s paradoxes. Synthese, 110(1):143–166, January 1997. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.

com/content/pdf/10.1023/A%3A1004967023017.pdf.

Ismael:1997:CP

[578] Jenann Ismael. Curie’s Principle. Synthese, 110(2):167–190, February 1997. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/content/pdf/10.1023/A%
3A1004929109216.pdf.

Lehner:1997:WIF

[579] Christoph Lehner. What it feels like to be in a superposition. And why. Synthese, 110(2):191–216, February 1997. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

content/pdf/10.1023/A%3A1004981126055.pdf.

Hudson:1997:CPE

[580] Robert G. Hudson. Classical physics and early quantum theory: a legitimate case of theoretical underdetermination. Synthese, 110(2):217–256, February 1997. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964

(electronic). URL http://link.springer.com/content/pdf/10.1023/
A%3A1004933210125.pdf.

Michell:1997:BRC

[581] Joel Michell. Bertrand Russell’s 1897 critique of the traditional theory of measurement. Synthese, 110(2):257–276, February 1997. CODEN SYN-TAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.

springer.com/content/pdf/10.1023/A%3A1004985226963.pdf.

DaSilva:1997:HPW

[582] Jairo José Da Silva. Husserl’s phenomenology and Weyl’s predictivism. Synthese, 110(2):277–296, February 1997. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

content/pdf/10.1023/A%3A1004937311034.pdf.

Zupco:1997:WSS

[583] Jack Zupco. What is the science of the soul? A case study in the evolution of late medieval natural philosophy. Synthese, 110(2):297–334, February 1997. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964

(electronic). URL http://link.springer.com/content/pdf/10.1023/
A%3A1004969404080.pdf.

Jacobson:1997:EAGa

[584] Stephen Jacobson. Externalism and action-guiding epistemic norms. Synthese, 110(3):343–355, March 1997. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

content/pdf/10.1023/A%3A1004945312756.pdf.

Tan:1997:DLR

[585] Yao-Hua Tan. Is default logic a reinvention of inductive–statistical reasoning? Synthese, 110(3):357–379, March 1997. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.

com/content/pdf/10.1023/A%3A1004999920152.pdf.

Jacobson:1997:EAGb

[586] Stephen Jacobson. Externalism and action-guiding epistemic norms. Synthese, 110(3):381–397, March 1997. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

content/pdf/10.1023/A%3A1004941211847.pdf.

Bergmann:1997:IEN

[587] Michael Bergmann. Internalism, externalism and the no-defeater condition. Synthese, 110(3):399–417, March 1997. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.

com/content/pdf/10.1023/A%3A1004993228686.pdf.

Jacobsen:1997:SQS

[588] Rockney Jacobsen. Self-quotation and self-knowledge. Synthese, 110(3): 419–445, March 1997. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/content/pdf/10.
1023/A%3A1004905327777.pdf.

Blachowicz:1997:RJS

[589] James Blachowicz. Reciprocal justification in science and moral theory. Synthese, 110(3):447–468, March 1997. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

content/pdf/10.1023/A%3A1004997329594.pdf.

Fischer:1997:VIT

[590] Eugen Fischer. On the very idea of a theory of meaning for a natural language. Synthese, 111(1):1–8, April 1997. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.

com/content/pdf/10.1023/A%3A1004973407034.pdf.

Cowie:1997:LPL

[591] Fiona Cowie. The logical problem of language acquisition. Synthese, 111 (1):17–51, April 1997. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/content/pdf/10.
1023/A%3A1004975305820.pdf.

Paul:1997:TCT

[592] L. A. Paul. Truth conditions of tensed sentence types. Synthese, 111(1): 53–72, April 1997. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964

(electronic). URL http://link.springer.com/content/pdf/10.1023/
A%3A1004977524781.pdf.

Nelson:1997:PFT

[593] R. J. Nelson. Proxy functions, truth and reference. Synthese, 111(1):73– 96, April 1997. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964

(electronic). URL http://link.springer.com/content/pdf/10.1023/
A%3A1004985725690.pdf.

Morton:1997:H

[594] Adam Morton. Hypercomparatives. Synthese, 111(1):97–114, April 1997. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/content/pdf/10.1023/A%
3A1004933708851.pdf.

Harvey:1997:LIP

[595] Charles W. Harvey. Liberal indoctrination and the problem of community. Synthese, 111(1):115–130, April 1997. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

content/pdf/10.1023/A%3A1004949413665.pdf.

Kanamori:1997:P

[596] Akihiro Kanamori. Preface. Synthese, 111(2):131–132, May 1997. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://

link.springer.com/content/pdf/10.1023/A%3A1004910218600.pdf.

Jaffe:1997:PEM

[597] Arthur Jaffe. Proof and the evolution of mathematics. Synthese, 111(2): 133–146, May 1997. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/content/pdf/10.
1023/A%3A1004903010713.pdf.

MacLane:1997:DPP

[598] Saunders MacLane. Despite physicists, proof is essential in mathematics. Synthese, 111(2):147–154, May 1997. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

content/pdf/10.1023/A%3A1004918402670.pdf.

Hintikka:1997:RFM

[599] Jaakko Hintikka. A revolution in the foundations of mathematics? Synthese, 111(2):155–170, May 1997. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

content/pdf/10.1023/A%3A1004970403579.pdf.

Rota:1997:PMB

[600] Gian-Carlo Rota. The phenomenology of mathematical beauty. Synthese, 111(2):171–182, May 1997. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/content/pdf/
10.1023/A%3A1004930722234.pdf.

Rota:1997:PMP

[601] Gian-Carlo Rota. The phenomenology of mathematical proof. Synthese, 111(2):183–196, May 1997. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/content/pdf/
10.1023/A%3A1004974521326.pdf.

Mazur:1997:C

[602] B. Mazur. Conjecture. Synthese, 111(2):197–210, May 1997. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://

link.springer.com/content/pdf/10.1023/A%3A1004934806305.pdf.

Mcintyre:1997:EI

[603] Lee Mcintyre and Eric Scerri. Editorial introduction. Synthese, 111(3): 211–212, June 1997. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/content/pdf/10.
1023/A%3A1004983130895.pdf.

Scerri:1997:CPC

[604] Eric R. Scerri and Lee McIntyre. The case for the philosophy of chemistry. Synthese, 111(3):213–232, June 1997. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

content/pdf/10.1023/A%3A1004949814965.pdf.

Ramsey:1997:MSR

[605] Jeffry L. Ramsey. Molecular shape, reduction, explanation and approximate concepts. Synthese, 111(3):233–251, June 1997. CODEN SYN-TAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.

springer.com/content/pdf/10.1023/A%3A1004901931804.pdf.

VanBrakel:1997:CST

[606] J. Van Brakel. Chemistry as the science of the transformation of substances. Synthese, 111(3):253–282, June 1997. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.

com/content/pdf/10.1023/A%3A1004953915874.pdf.

Gavroglu:1997:PIH

[607] Kostas Gavroğlu. Philosophical issues in the history of chemistry. Synthese, 111(3):283–304, June 1997. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

content/pdf/10.1023/A%3A1004906132712.pdf.

Scerri:1997:BPC

[608] E. R. Scerri. Bibliography on philosophy of chemistry. Synthese, 111(3): 305–324, June 1997. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/content/pdf/10.
1023/A%3A1004958116783.pdf.

Demopoulos:1997:MRB

[609] William Demopoulos. In memoriam: Robert E. Butts 1928–1997. Synthese, 112(1):1–2, July 1997. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/content/pdf/
10.1023/A%3A1004907530863.pdf.

Hintikka:1997:E

[610] Jaakko Hintikka. Editorial. Synthese, 112(1):3–5, July 1997. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://

link.springer.com/content/pdf/10.1023/A%3A1004972610232.pdf.

Haack:1997:PEI

[611] Susan Haack. Prècis of evidence and inquiry: Towards reconstruction in epistemology. Synthese, 112(1):7–11, July 1997. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.

com/content/pdf/10.1023/A%3A1004924514117.pdf.

Bonjour:1997:HJE

[612] Laurence Bonjour. Haack on justification and experience. Synthese, 112 (1):13–23, July 1997. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/content/pdf/10.
1023/A%3A1004972413209.pdf.

Haack:1997:RB

[613] Susan Haack. Reply to bonjour. Synthese, 112(1):25–35, July 1997. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/content/pdf/10.1023/A%
3A1004920429138.pdf.

Resnik:1997:LD

[614] David Resnik. Laws and development. Synthese, 112(1):37–51, July 1997. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/content/pdf/10.1023/A%
3A1004968312300.pdf.

Bonk:1997:NGQ

[615] Thomas Bonk. Newtonian gravity, quantum discontinuity and the determination of theory by evidence. Synthese, 112(1):53–73, July 1997. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/content/pdf/10.1023/A%
3A1004916328230.pdf.

Jones:1997:URN

[616] Todd Jones. Unification, reduction, and non-ideal explanations. Synthese, 112(1):75–96, July 1997. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/content/pdf/
10.1023/A%3A1004928615026.pdf.

Zangari:1997:SR

[617] Mark Zangari and Dan Censor. Spectral representations. Synthese, 112(1): 97–123, July 1997. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964

(electronic). URL http://link.springer.com/content/pdf/10.1023/
A%3A1004964211391.pdf.

Gaudiano:1997:BRP

[618] Paolo Gaudiano. Being in the right place at the right time. Synthese, 112(1):125–133, July 1997. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/content/pdf/
10.1023/A%3A1005017800922.pdf.

Anonymous:1997:PPa

[619] Anonymous. Problems of philosophy. Synthese, 112(1):135–136, July 1997. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/content/pdf/10.1023/A%
3A1017124709324.pdf.

Xu:1997:CBT

[620] Ming Xu. Causation in branching time (i): Transitions, events and causes. Synthese, 112(2):137–192, August 1997. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

content/pdf/10.1023/A%3A1004948726086.pdf.

Martin:1997:NPR

[621] C. B. Martin. On the need for properties: The road to Pythagoreanism and back. Synthese, 112(2):193–231, August 1997. CODEN SYN-TAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.

springer.com/content/pdf/10.1023/A%3A1004983524345.pdf.

Dowe:1997:DBT

[622] Phil Dowe. A defense of backwards in time causation models in quantum mechanics. Synthese, 112(2):233–246, August 1997. CODEN SYN-TAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.

springer.com/content/pdf/10.1023/A%3A1004932911141.pdf.

Slowik:1997:HCM

[623] Edward Slowik. Huygens’ center-of-mass space–time reference frame: Constructing a Cartesian dynamics in the wake of Newton’s “De gravitatione” argument. Synthese, 112(2):247–269, August 1997. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://

link.springer.com/content/pdf/10.1023/A%3A1004980827071.pdf.

daCosta:1997:SPS

[624] Newton C. A. da Costa, Otávio Bueno, and Steven French. Suppes predicates for space–time. Synthese, 112(2):271–279, August 1997. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://

link.springer.com/content/pdf/10.1023/A%3A1004984927979.pdf.

Tsuji:1997:SPM

[625] Marcelo Tsuji. Suppes predicates for meta-ranking structures. Synthese, 112(2):281–299, August 1997. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/content/pdf/
10.1023/A%3A1004920608415.pdf.

Anonymous:1997:PPb

[626] Anonymous. Problems of philosophy. Synthese, 112(2):301–302, August 1997. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964

(electronic). URL http://link.springer.com/content/pdf/10.1023/
A%3A1017190525254.pdf.

Banchetti-Robino:1997:HTL

[627] Marina Paola Banchetti-Robino. Husserl’s theory of language as calculus ratiocinator. Synthese, 112(3):303–321, September 1997. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://

link.springer.com/content/pdf/10.1023/A%3A1004944525177.pdf.

Beyer:1997:HRH

[628] Christian Beyer. Hussearle’s representationalism and the “hypothesis of the background”. Synthese, 112(3):323–352, September 1997. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://

link.springer.com/content/pdf/10.1023/A%3A1004992424269.pdf.

Frascolla:1997:TSA

[629] Pasquale Frascolla. The Tractatus system of arithmetic. Synthese, 112(3): 353–378, September 1997. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/content/pdf/
10.1023/A%3A1004952810156.pdf.

Oliveri:1997:MSP

[630] Gianluigi Oliveri. Mathematics. A science of patterns? Synthese, 112(3): 379–402, September 1997. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/content/pdf/
10.1023/A%3A1004906107430.pdf.

Koslicki:1997:INA

[631] Kathrin Koslicki. Isolation and non-arbitrary division: Frege’s two criteria for counting. Synthese, 112(3):403–430, September 1997. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://

link.springer.com/content/pdf/10.1023/A%3A1004903625876.pdf.

Anonymous:1997:PPPa

[632] Anonymous. Problems of philosophy problem #3: One-world assumption and Frege’s sense-reference distinction. Synthese, 112(3):431–432, September 1997. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964

(electronic). URL http://link.springer.com/content/pdf/10.1023/
A%3A1017147208339.pdf.

Shin:1997:KSR

[633] Sun joo Shin. Kant’s syntheticity revisited by Peirce. Synthese, 113(1):1– 41, October 1997. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964

(electronic). URL http://link.springer.com/content/pdf/10.1023/
A%3A1005068218051.pdf.

Forster:1997:KBP

[634] Paul Forster. Kant, Boole and Peirce’s early metaphysics. Synthese, 113 (1):43–70, October 1997. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/content/pdf/
10.1023/A%3A1005012100305.pdf.

Wojtowicz:1997:MES

[635] Randy Wojtowicz. The metaphysical expositions of space and time. Synthese, 113(1):71–115, October 1997. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

content/pdf/10.1023/A%3A1005008016234.pdf.

Wolff:1997:RHK

[636] Robert Paul Wolff. Robert Howell, 1992, Kant’s transcendental deduction: An analysis of main themes in his critical philosophy. Synthese, 113(1): 117–144, October 1997. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/content/pdf/10.
1023/A%3A1005064117143.pdf.

Hill:1997:DGC

[637] Claire Ortiz Hill. Did Georg Cantor influence Edmund Husserl? Synthese, 113(1):145–170, October 1997. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

content/pdf/10.1023/A%3A1005099615326.pdf.

Miller:1997:AVT

[638] Steven I. Miller and Marcel Fredericks. Another view of translation manuals and the study of science. Synthese, 113(2):171–193, November 1997. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/content/pdf/10.1023/A%
3A1005043512600.pdf.

Davis:1997:RBC

[639] John R. Davis. A reconsideration of Britain as commercial hegemon in the nineteenth century. Synthese, 113(2):195–203, November 1997. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://

link.springer.com/content/pdf/10.1023/A%3A1004984710348.pdf.

Dacey:1997:NCB

[640] Raymond Dacey and Kevin P. Murrin. Nineteenth century Britain as a subtle commercial hegemon. Synthese, 113(2):205–216, November 1997. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/content/pdf/10.1023/A%
3A1005000926712.pdf.

Woods:1997:HAF

[641] John Woods and Hans V. Hansen. Hintikka on Aristotle’s fallacies. Synthese, 113(2):217–239, November 1997. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

content/pdf/10.1023/A%3A1005035327621.pdf.

Hintikka:1997:WWA

[642] Jaakko Hintikka. What was Aristotle doing in his early logic, anyway? a reply to Woods and Hansen. Synthese, 113(2):241–249, November 1997. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/content/pdf/10.1023/A%
3A1005095513508.pdf.

Weller:1997:BM

[643] Cass Weller. Bonjour and mentalese. Synthese, 113(2):251–263, November 1997. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/content/pdf/10.1023/A%
3A1005039428529.pdf.

Weiss:1997:PCP

[644] Bernhard Weiss. Proof and canonical proof. Synthese, 113(2):265–284, November 1997. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964

(electronic). URL http://link.springer.com/content/pdf/10.1023/
A%3A1005067014400.pdf.

Bringsjord:1997:CCA

[645] Selmer Bringsjord and Michael Zenzen. Cognition is not computation: The argument from irreversibility. Synthese, 113(2):285–320, November 1997. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/content/pdf/10.1023/A%
3A1005019131238.pdf.

Anonymous:1997:PPPb

[646] Anonymous. Problems of philosophy. problem #4: Can nature impart to us general truths? Synthese, 113(2):321–322, November 1997. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://

link.springer.com/content/pdf/10.1023/A%3A1017132503908.pdf.

Pessoa:1997:CDA

[647] Osvaldo Pessoa, Jr. Can the decoherence approach help to solve the measurement problem? Synthese, 113(3):323–346, December 1997. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://

link.springer.com/content/pdf/10.1023/A%3A1004994303863.pdf.

Barrett:1997:NEB

[648] Jeffrey A. Barrett. On the nature of experience in the bare theory. Synthese, 113(3):347–355, December 1997. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

content/pdf/10.1023/A%3A1004983313200.pdf.

Leiber:1997:AID

[649] Theodor Leiber. On the actual impact of deterministic chaos. Synthese, 113(3):357–379, December 1997. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

content/pdf/10.1023/A%3A1004944713074.pdf.

Tzouvaras:1997:OSC

[650] Athanassios Tzouvaras. The order structure of continua. Synthese, 113 (3):381–421, December 1997. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/content/pdf/
10.1023/A%3A1005094430329.pdf.

Subramanyam:1997:MJS

[651] J. Subramanyam. Measurement and the justification of the statistical postulate in Bohm’s causal interpretation of quantum mechanics. Synthese, 113(3):423–445, December 1997. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

content/pdf/10.1023/A%3A1005061321655.pdf.

Anonymous:1997:PPP

[652] Anonymous. Problems of phylosophy; problem #5: Wittgenstein’s rulefollowing discussion. Synthese, 113(3):445, December 1997. CODEN SYN-TAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.

springer.com/content/pdf/10.1023/A%3A1017181804817.pdf.

Hintikka:1998:EN

[653] Jaakko Hintikka. Editorial note. Synthese, 114(1):1, January 1998. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/content/pdf/10.1023/A%
3A1005002526198.pdf.

Dawson:1998:WHG

[654] J. W. Dawson, Jr. What hath Gödel wrought? Synthese, 114(1):3–12, January 1998. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964

(electronic). URL http://link.springer.com/content/pdf/10.1023/
A%3A1005046308451.pdf.

Hintikka:1998:GPA

[655] Jaakko Hintikka. On Gödel’s philosophical assumptions. Synthese, 114 (1):13–23, January 1998. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/content/pdf/
10.1023/A%3A1005042207543.pdf.

Tymoczko:1998:GCM

[656] Thomas Tymoczko. Gödel and the concept of meaning in mathematics. Synthese, 114(1):25–40, January 1998. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

content/pdf/10.1023/A%3A1005090106634.pdf.

Parikh:1998:LSP

[657] Rohit Parikh. Length and structure of proofs. Synthese, 114(1):41–48, January 1998. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964

(electronic). URL http://link.springer.com/content/pdf/10.1023/
A%3A1005050425290.pdf.

Boos:1998:TME

[658] William Boos. The Transzendenz of mathematical ‘experience’. Synthese, 114(1):49–98, January 1998. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/content/pdf/
10.1023/A%3A1005085922564.pdf.

Kreisel:1998:STA

[659] G. Kreisel. Second thoughts around some of Gödel’s writings. Synthese, 114(1):99–160, January 1998. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/content/pdf/
10.1023/A%3A1005006610269.pdf.

Anonymous:1998:PPPa

[660] Anonymous. Problems of philosophy. problem #6: The varieties of completeness and their role in the foundations of mathematics. Synthese, 114 (1):161–162, January 1998. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/content/pdf/
10.1023/A%3A1017133927107.pdf.

Cocchiarella:1998:RCR

[661] Nino B. Cocchiarella. Reference in conceptual realism. Synthese, 114 (2):169–202, February 1998. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/content/pdf/
10.1023/A%3A1005005113229.pdf.

Hintikka:1998:PID

[662] Jaakko Hintikka. Perspectival identification, demonstratives and “small worlds”. Synthese, 114(2):203–232, February 1998. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.

springer.com/content/pdf/10.1023/A%3A1005072129158.pdf.

Lenci:1998:SP

[663] Alessandro Lenci. The structure of predication. Synthese, 114(2):233–276, February 1998. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964

(electronic). URL http://link.springer.com/content/pdf/10.1023/
A%3A1005068021890.pdf.

Butler:1998:CCI

[664] Keith Butler. Content, computation, and individuation. Synthese, 114 (2):277–292, February 1998. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/content/pdf/
10.1023/A%3A1004996729912.pdf.

Pelczar:1998:ICN

[665] M. Pelczar and J. Rainsbury. The indexical character of names. Synthese, 114(2):293–317, February 1998. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

content/pdf/10.1023/A%3A1004992629004.pdf.

Wreen:1998:PNN

[666] Michael Wreen. Proper names and the necessity of identity statements. Synthese, 114(2):319–335, February 1998. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

content/pdf/10.1023/A%3A1005002220981.pdf.

Maitzen:1998:KPE

[667] Stephen Maitzen. The knower paradox and epistemic closure. Synthese, 114(2):337–354, February 1998. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

content/pdf/10.1023/A%3A1005064624642.pdf.

Alper:1998:NSC

[668] Joseph S. Alper and Mark Bridger. Newtonian supertasks: A critical analysis. Synthese, 114(2):355–369, February 1998. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.

com/content/pdf/10.1023/A%3A1005057130067.pdf.

Anonymous:1998:PPPb

[669] Anonymous. Problems of philosophy. problem #7: Logic without the Frege–Russell ambiguity assumption. Synthese, 114(2):371, February 1998. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/content/pdf/10.1023/A%
3A1017129820747.pdf.

Saunders:1998:TQM

[670] Simon Saunders. Time, quantum mechanics, and probability. Synthese, 114(3):373–404, March 1998. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/content/pdf/
10.1023/A%3A1005079904008.pdf.

Bedard:1998:BSB

[671] Katherine Bedard. Bohm, spin, and the Bell inequalities. Synthese, 114 (3):405–444, March 1998. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/content/pdf/
10.1023/A%3A1005097302191.pdf.

Shagrir:1998:MRC

[672] Oron Shagrir. Multiple realization, computation and the taxonomy of psychological states. Synthese, 114(3):445–461, March 1998. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://

link.springer.com/content/pdf/10.1023/A%3A1005072701509.pdf.

Montgomery:1998:GEC

[673] Richard Montgomery. Grades of explanation in cognitive science. Synthese, 114(3):463–495, March 1998. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

content/pdf/10.1023/A%3A1005084615487.pdf.

Anonymous:1998:PPPc

[674] Anonymous. Problems of philosophy. problem #8: The epistemic element in intuitionistic logic. Synthese, 114(3):497, March 1998. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://

link.springer.com/content/pdf/10.1023/A%3A1017148912320.pdf.

Anderson:1998:TOP

[675] R. Lanier Anderson. Truth and objectivity in perspectivism. Synthese, 115(1):1–32, April 1998. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/content/pdf/
10.1023/A%3A1004984312166.pdf.

Shieh:1998:CFA

[676] Sanford Shieh. On the conceptual foundations of anti-realism. Synthese, 115(1):33–70, April 1998. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/content/pdf/
10.1023/A%3A1005020809621.pdf.

Belousek:1998:HSM

[677] Darrin W. Belousek. Husserl on scientific method and conceptual change: A realist appraisal. Synthese, 115(1):71–98, April 1998. CODEN SYN-TAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.

springer.com/content/pdf/10.1023/A%3A1005040800466.pdf.

Ben-Menahem:1998:EDW

[678] Yemima Ben-Menahem. Explanation and description: Wittgenstein on convention. Synthese, 115(1):99–130, April 1998. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.

com/content/pdf/10.1023/A%3A1005016201213.pdf.

Park:1998:WUW

[679] Byong-Chul Park. Wittgenstein’s use of the word ’aspekt’. Synthese, 115(1):131–140, April 1998. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/content/pdf/
10.1023/A%3A1005076802418.pdf.

Anonymous:1998:PPPd

[680] Anonymous. Problems of philosophy problem #9: Reference vs. identification. Synthese, 115(1):141, April 1998. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.

com/content/pdf/10.1023/A%3A1017116818121.pdf.

Owens:1998:PEC

[681] Joseph Owens. Psychological explanation and causal deviancy. Synthese, 115(2):143–169, May 1998. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/content/pdf/
10.1023/A%3A1005036732326.pdf.

Vallentyne:1998:NRA

[682] Peter Vallentyne. The nomic role account of carving reality at the joints. Synthese, 115(2):171–198, May 1998. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

content/pdf/10.1023/A%3A1005024818347.pdf.

Lievers:1998:TVM

[683] Menno Lievers. Two versions of the manifestation argument. Synthese, 115(2):199–227, May 1998. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/content/pdf/
10.1023/A%3A1005056322825.pdf.

Hild:1998:CAA

[684] Matthias Hild. The coherence argument against conditionalization. Synthese, 115(2):229–258, May 1998. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

content/pdf/10.1023/A%3A1005082908147.pdf.

Laraudogoitia:1998:IMC

[685] Jon Perez Laraudogoitia. Infinity machines and creation ex nihilo. Synthese, 115(2):259–265, May 1998. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

content/pdf/10.1023/A%3A1005095006443.pdf.

Anonymous:1998:PPa

[686] Anonymous. Problems of philosophy. Synthese, 115(2):267, May 1998. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/content/pdf/10.1023/A%
3A1017119010325.pdf.

Hellan:1998:HPP

[687] Patrick A. Hellan. Hermeneutical philosophy and pragmatism: A philosophy of science. Synthese, 115(3):269–302, June 1998. CODEN SYN-TAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.

springer.com/content/pdf/10.1023/A%3A1005032631417.pdf.

Mura:1998:HIL

[688] Alberto Mura. Hume’s inductive logic. Synthese, 115(3):303–331, June 1998. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/content/pdf/10.1023/A%
3A1005041012179.pdf.

Ekstrom:1998:FCC

[689] Laura Waddell Ekstrom. Freedom, causation, and the consequence argument. Synthese, 115(3):333–354, June 1998. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.

com/content/pdf/10.1023/A%3A1005009029926.pdf.

Sober:1998:GSE

[690] Elliott Sober. To give a surprise exam, use game theory. Synthese, 115(3): 355–373, June 1998. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/content/pdf/10.
1023/A%3A1005012607804.pdf.

Anonymous:1998:REP

[691] Anonymous. Review essay: The psychologists return. Synthese, 115(3): 375–393, June 1998. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/content/pdf/10.
1023/A%3A1005056929017.pdf.

Anonymous:1998:PPT

[692] Anonymous. Problems of philosophy #11: Two meanings of scope distinguished from each other. Synthese, 115(3):395, June 1998. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://

link.springer.com/content/pdf/10.1023/A%3A1017147118718.pdf.

Horsten:1998:DEA

[693] Leon Horsten. In defense of epistemic arithmetic. Synthese, 116(1):1– 25, July 1998. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964

(electronic). URL http://link.springer.com/content/pdf/10.1023/
A%3A1005016405987.pdf.

Harper:1998:PMP

[694] William Harper. Papier Mâché problems in epistemology: A defense of strong internalism*. Synthese, 116(1):27–49, July 1998. CODEN SYN-TAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.

springer.com/content/pdf/10.1023/A%3A1005014714579.pdf.

Anonymous:1998:EMF

[695] Anonymous. Empiricism and the myth of fundamental measurement. Synthese, 116(1):51–73, July 1998. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

content/pdf/10.1023/A%3A1005016725551.pdf.

Anonymous:1998:PSI

[696] Anonymous. Part structures, integrity, and the mass-count distinction. Synthese, 116(1):75–111, July 1998. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

content/pdf/10.1023/A%3A1005046308299.pdf.

Anonymous:1998:PPPe

[697] Anonymous. Problems of philosophy problem #12: Epistemology from a strategic viewpoint. Synthese, 116(1):113, July 1998. CODEN SYN-TAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.

springer.com/content/pdf/10.1023/A%3A1017169801880.pdf.

Anonymous:1998:A

[698] Anonymous. Announcement. Synthese, 116(1):115–116, July 1998. CO-DEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/content/pdf/10.1023/A%3A1017154622646.
pdf.

Bassler:1998:LIE

[699] O. Bradley Bassler. Leibniz on intension, extension, and the representation of syllogistic inference. Synthese, 116(2):117–139, August 1998. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://

link.springer.com/content/pdf/10.1023/A%3A1005076530958.pdf.

Kvasz:1998:HGD

[700] Ladislav Kvasz. History of geometry and the development of the form of its language. Synthese, 116(2):141–186, August 1998. CODEN SYN-TAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.

springer.com/content/pdf/10.1023/A%3A1005008423734.pdf.

Bradley:1998:RTD

[701] Richard Bradley. A representation theorem for a decision theory with conditionals. Synthese, 116(2):187–229, August 1998. CODEN SYN-TAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.

springer.com/content/pdf/10.1023/A%3A1005030124500.pdf.

Wybraniec-Skardowska:1998:LPI

[702] Urszula Wybraniec-Skardowska. Logical and philosophical ideas in certain formal approaches to language. Synthese, 116(2):231–277, August 1998. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/content/pdf/10.1023/A%
3A1005098325137.pdf.

Anonymous:1998:PPPf

[703] Anonymous. Problems of philosophy. problem #13: Carnap vindicated? Synthese, 116(2):279, August 1998. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

content/pdf/10.1023/A%3A1017115018818.pdf.

McLaughlin:1998:TLD

[704] William I. McLaughlin. Thomson’s lamp is dysfunctional. Synthese, 116 (3):281–301, September 1998. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/content/pdf/
10.1023/A%3A1005045200162.pdf.

Schlosser:1998:SRP

[705] Gerhard Schlosser. Self-re-production and functionality. Synthese, 116(3): 303–354, September 1998. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/content/pdf/
10.1023/A%3A1005073307193.pdf.

Jonathan:1998:PCO

[706] Brian Jonathan. Pluralism, causation and overdetermination. Synthese, 116(3):355–378, September 1998. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

content/pdf/10.1023/A%3A1005018414120.pdf.

Raatikainen:1998:SI

[707] Panu Raatikainen. Simplicity and incompleteness. Synthese, 116(3):355– 362, September 1998. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/content/pdf/10.
1023/A%3A1005033725849.pdf.

Sarkar:1998:ARA

[708] Husain Sarkar. Anti-realism against methodology. Synthese, 116(3):379– 402, September 1998. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/content/pdf/10.
1023/A%3A1005070724228.pdf.

Wilburn:1998:PWM

[709] John Bart Wilburn. A possible worlds model of object recognition. Synthese, 116(3):403–438, September 1998. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

content/pdf/10.1023/A%3A1005066410812.pdf.

Simon:1998:HMI

[710] Herbert A. Simon and Stuart A. Eisenstadt. Human and machine interpretation of expressions in formal systems. Synthese, 116(3):439–461, September 1998. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964

(electronic). URL http://link.springer.com/content/pdf/10.1023/
A%3A1005053105533.pdf.

Anonymous:1998:PPb

[711] Anonymous. Problems of philosophy. Synthese, 116(3):463, September 1998. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/content/pdf/10.1023/A%
3A1017126210100.pdf.

Bain:1998:WQD

[712] Jonathan Bain. Weinberg on QFT: Demonstrative induction and underdetermination. Synthese, 117(1):1–30, October 1998. CODEN SYN-TAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.

springer.com/content/pdf/10.1023/A%3A1005025424031.pdf.

Kronz:1998:BOI

[713] Frederick M. Kronz. Bohm’s ontological interpretation and its relations to three formulations of quantum mechanics. Synthese, 117(1):31–52, October 1998. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964

(electronic). URL http://link.springer.com/content/pdf/10.1023/
A%3A1005098623123.pdf.

Bailey:1998:SP

[714] Andrew Bailey. Supervenience and physicalism. Synthese, 117(1):53–73, October 1998. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964

(electronic). URL http://link.springer.com/content/pdf/10.1023/
A%3A1005080908570.pdf.

Paris:1998:CSM

[715] Jeff Paris. Common sense and maximum entropy. Synthese, 117(1):75– 93, October 1998. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964

(electronic). URL http://link.springer.com/content/pdf/10.1023/
A%3A1005081609010.pdf.

Rupert:1998:RBN

[716] Robert D. Rupert. On the relationship between naturalistic semantics and individuation criteria for terms in a language of thought. Synthese, 117 (1):95–131, October 1998. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/content/pdf/
10.1023/A%3A1005077508102.pdf.

Sher:1998:PST

[717] Gila Sher. On the possibility of a substantive theory of truth. Synthese, 117(1):133–172, October 1998. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

content/pdf/10.1023/A%3A1005068021441.pdf.

Anonymous:1998:PPc

[718] Anonymous. Problems of philosophy. Synthese, 117(1):173, October 1998. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/content/pdf/10.1023/A%
3A1017175824579.pdf.

Tuomela:1998:CAC

[719] Raimo Tuomela and Wolfgang Balzer. Collective acceptance and collective social notions. Synthese, 117(2):175–205, November 1998. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://

link.springer.com/content/pdf/10.1023/A%3A1005000928559.pdf.

Gluer:1998:RMP

[720] Kathrin Glüer and Peter Pagin. Rules of meaning and practical reasoning. Synthese, 117(2):207–227, November 1998. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

content/pdf/10.1023/A%3A1005162503125.pdf.

Goldman:1998:RMR

[721] Alan H. Goldman. Rules and moral reasoning. Synthese, 117(2):229–250, November 1998. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964

(electronic). URL http://link.springer.com/content/pdf/10.1023/
A%3A1005174804034.pdf.

Oddie:1998:MRM

[722] Graham Oddie. Moral realism, moral relativism and moral rules (a compatibility argument). Synthese, 117(2):251–274, November 1998. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://

link.springer.com/content/pdf/10.1023/A%3A1005183121781.pdf.

Johansson:1998:HSL

[723] Ingvar Johansson. Hume’s surprise and the logic of belief changes. Synthese, 117(2):275–291, November 1998. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

content/pdf/10.1023/A%3A1005179004942.pdf.

Anonymous:1998:PPd

[724] Anonymous. Problems of philosophy. Synthese, 117(2):293, November 1998. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/content/pdf/10.1023/A%
3A1017104908649.pdf.

Jackman:1998:CL

[725] Henry Jackman. Convention and language. Synthese, 117(3):295–312, December 1998. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964

(electronic). URL http://link.springer.com/content/pdf/10.1023/
A%3A1005134521663.pdf.

Schroder:1998:KRC

[726] Jürgen Schröder. Knowledge of rules, causal systematicity, and the language of thought. Synthese, 117(3):313–330, December 1998. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://

link.springer.com/content/pdf/10.1023/A%3A1005096727650.pdf.

Adams:1998:IAF

[727] Ernest W. Adams. Idealization in applied first-order logic. Synthese, 117 (3):331–354, December 1998. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/content/pdf/
10.1023/A%3A1005090932292.pdf.

Giaretta:1998:LRL

[728] Pierdaniele Giaretta. Liar, reducibility and language. Synthese, 117(3): 355–374, December 1998. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/content/pdf/
10.1023/A%3A1005076008996.pdf.

Gomez-torrente:1998:LTT

[729] Mario Gómez-torrente. Logical truth and Tarskian logical truth. Synthese, 117(3):375–408, December 1998. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

content/pdf/10.1023/A%3A1005165824990.pdf.

Weintraub:1998:SMM

[730] Ruth Weintraub. The spatiality of the mental and the mind–body problem. Synthese, 117(3):409–417, December 1998. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.

com/content/pdf/10.1023/A%3A1005022524925.pdf.

Changizi:1998:PBS

[731] Mark A. Changizi and Timothy P. Barber. A paradigm-based solution to the riddle of induction. Synthese, 117(3):419–484, December 1998. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/content/pdf/10.1023/A%
3A1005029624940.pdf.

Anonymous:1998:PPe

[732] Anonymous. Problems of philosophy. Synthese, 117(3):485, December 1998. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/content/pdf/10.1023/A%
3A1017156925488.pdf.

Lubcke:1999:SIH

[733] Poul Lübcke. A semantic interpretation of Husserl’s Epoché. Synthese, 118(1):1–12, January 1999. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/content/pdf/
10.1023/A%3A1005101213436.pdf.

Anonymous:1999:RMF

[734] Anonymous. Representationalism and the metonymic fallacy. Synthese, 118(1):13–30, January 1999. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/content/pdf/
10.1023/A%3A1005142908893.pdf.

Haaparanta:1999:PNP

[735] Leila Haaparanta. On the possibility of naturalistic and of pure epistemology. Synthese, 118(1):31–47, January 1999. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.

com/content/pdf/10.1023/A%3A1005192826640.pdf.

Heinamaa:1999:MPM

[736] Sara Heinämaa. Merleau–Ponty’s modification of phenomenology: Cognition, passion and philosophy. Synthese, 118(1):49–68, January 1999. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/content/pdf/10.1023/A%
3A1005140809802.pdf.

Anonymous:1999:MPN

[737] Anonymous. Merleau–Ponty’s notion of pre-reflective intentionality. Synthese, 118(1):69–88, January 1999. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

content/pdf/10.1023/A%3A1005144911619.pdf.

Brinck:1999:RSA

[738] Ingar Brinck and Peter Gärdenfors. Representation and self-awareness in intentional agents. Synthese, 118(1):89–104, January 1999. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://

link.springer.com/content/pdf/10.1023/A%3A1005109414345.pdf.

Anonymous:1999:TTR

[739] Anonymous. Towards a theory of reflexive intentional systems. Synthese, 118(1):105–117, January 1999. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

content/pdf/10.1023/A%3A1005188725731.pdf.

Anonymous:1999:E

[740] Anonymous. Erratum. Synthese, 118(1):119, January 1999. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://

link.springer.com/content/pdf/10.1023/A%3A1017221321696.pdf.

Latham:1999:DKP

[741] Noa Latham. Davidson and Kim on psychophysical laws. Synthese, 118 (2):121–143, February 1999. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/content/pdf/
10.1023/A%3A1005052912629.pdf.

Anonymous:1999:RBT

[742] Anonymous. Rethinking Burge’s thought experiment. Synthese, 118(2): 145–164, February 1999. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/content/pdf/
10.1023/A%3A1017145322572.pdf.

Goldberg:1999:PES

[743] Sanford C. Goldberg. The psychology and epistemology of self-knowledge. Synthese, 118(2):165–199, February 1999. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

content/pdf/10.1023/A%3A1005110319056.pdf.

Evnine:1999:BC

[744] Simon J. Evnine. Believing conjunctions. Synthese, 118(2):201–227, February 1999. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964

(electronic). URL http://link.springer.com/content/pdf/10.1023/
A%3A1005114419965.pdf.

Liu:1999:AIL

[745] Chang Liu. Approximation, idealization, and laws of nature. Synthese, 118 (2):229–256, February 1999. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/content/pdf/
10.1023/A%3A1005186322310.pdf.

Peressini:1999:CMT

[746] Anthony Peressini. Confirming mathematical theories: an ontologically agnostic stance. Synthese, 118(2):257–277, February 1999. CODEN SYN-TAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.

springer.com/content/pdf/10.1023/A%3A1005158202218.pdf.

Rodych:1999:WIA

[747] Victor Rodych. Wittgenstein on irrationals and algorithmic decidability. Synthese, 118(2):279–304, February 1999. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

content/pdf/10.1023/A%3A1005191706419.pdf.

Anonymous:1999:PP

[748] Anonymous. Problems of philosophy. Synthese, 118(2):307, February 1999. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/content/pdf/10.1023/A%
3A1017109109558.pdf.

Atkinson:1999:PTD

[749] David Atkinson and Jeanne Peijnenburg. Probability as a theory dependent concept. Synthese, 118(3):307–328, March 1999. CODEN SYN-TAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.

springer.com/content/pdf/10.1023/A%3A1005242414754.pdf.

Schulte:1999:MBC

[750] Oliver Schulte. Minimal belief change and the Pareto principle. Synthese, 118(3):329–361, March 1999. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/content/pdf/
10.1023/A%3A1005162303126.pdf.

Montagna:1999:LCR

[751] Franco Montagna and Daniel Osherson. Learning to coordinate; a recursion theoretic perspective. Synthese, 118(3):363–382, March 1999. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/content/pdf/10.1023/A%
3A1005182121401.pdf.

Brogaard:1999:PTD

[752] Berit Brogaard. A Peircean theory of decision. Synthese, 118(3):383– 401, March 1999. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964

(electronic). URL http://link.springer.com/content/pdf/10.1023/
A%3A1005154800638.pdf.

Bartha:1999:SRP

[753] Paul Bartha and Christopher Hitchcock. The shooting-room paradox and conditionalizing on measurably challenged sets. Synthese, 118(3):403– 437, March 1999. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964

(electronic). URL http://link.springer.com/content/pdf/10.1023/
A%3A1005100407551.pdf.

Earman:1999:CPT

[754] John Earman and John Roberts. Ceteris paribus, there is no problem of provisos. Synthese, 118(3):439–478, March 1999. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.

com/content/pdf/10.1023/A%3A1005106917477.pdf.

Weber:1999:UWI

[755] Erik Weber. Unification: What is it, how do we reach and why do we want it? Synthese, 118(3):479–499, March 1999. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.

com/content/pdf/10.1023/A%3A1005134205471.pdf.

Anonymous:1999:PPP

[756] Anonymous. Problems of philosophy: Problem #19: Phenomenalism vs. phenomenology. Synthese, 118(3):501, March 1999. CODEN SYN-TAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.

springer.com/content/pdf/10.1023/A%3A1017161126396.pdf.

Blackmore:1999:I

[757] John Blackmore. Introduction. Synthese, 119(1–2):1–9, April 1999. CO-DEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/content/pdf/10.1023/A%3A1005218602230.
pdf.

Brush:1999:GGH

[758] Stephen G. Brush. Gadflies and geniuses in the history of gas theory. Synthese, 119(1–2):11–43, April 1999. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

content/pdf/10.1023/A%3A1005235019068.pdf.

Deltete:1999:HBE

[759] Robert Deltete. Helm and Boltzmann: Energetics at the Lübeck Naturforscherversammlung. Synthese, 119(1–2):45–68, April 1999. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://

link.springer.com/content/pdf/10.1023/A%3A1005287003138.pdf.

Jager:1999:TAB

[760] Gustav Jäger, Josef Nabl, and Stephan Meyer. Three assistants on Boltzmann. Synthese, 119(1–2):69–84, April 1999. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.

com/content/pdf/10.1023/A%3A1005239104047.pdf.

Stoltzner:1999:VIM

[761] Michael Stöltzner. Vienna indeterminism: Mach, Boltzmann, Exner. Synthese, 119(1–2):85–111, April 1999. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

content/pdf/10.1023/A%3A1005243320885.pdf.

deRegt:1999:LBB

[762] Henk W. de Regt. Ludwig Boltzmann’s Bildtheorie and scientific understanding. Synthese, 119(1–2):113–134, April 1999. CODEN SYN-TAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.

springer.com/content/pdf/10.1023/A%3A1005295304955.pdf.

Visser:1999:BWH

[763] Henk Visser. Boltzmann and Wittgenstein or how pictures became linguistic. Synthese, 119(1–2):135–156, April 1999. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.

com/content/pdf/10.1023/A%3A1005247421794.pdf.

Blackmore:1999:BE

[764] John Blackmore. Boltzmann and epistemology. Synthese, 119(1–2):157– 189, April 1999. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964

(electronic). URL http://link.springer.com/content/pdf/10.1023/
A%3A1005299405864.pdf.

Boltzmann:1999:BPN

[765] Ludwig Boltzmann. Boltzmann’s philosophy notes for three lectures (fall 1903). Synthese, 119(1–2):191–202, April 1999. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.

com/content/pdf/10.1023/A%3A1005251522702.pdf.

Tanaka:1999:BM

[766] Setsuko Tanaka. Boltzmann on mathematics. Synthese, 119(1–2):203– 232, April 1999. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964

(electronic). URL http://link.springer.com/content/pdf/10.1023/
A%3A1005203606773.pdf.

Smith:1999:STV

[767] Quentin Smith. The “sentence-type version” of the tenseless theory of time. Synthese, 119(3):233–251, June 1999. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.

com/content/pdf/10.1023/A%3A1005130104563.pdf.

Balashov:1999:ZVP

[768] Yuri Balashov. Zero-value physical quantities. Synthese, 119(3):253– 286, June 1999. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964

(electronic). URL http://link.springer.com/content/pdf/10.1023/
A%3A1005177016330.pdf.

Finkelstein:1999:STC

[769] J. Finkelstein. Space-time counterfactuals. Synthese, 119(3):287–298, June 1999. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964

(electronic). URL http://link.springer.com/content/pdf/10.1023/
A%3A1005286229775.pdf.

Forrest:1999:DPS

[770] Peter Forrest. In defence of the phase space picture. Synthese, 119(3):299– 311, June 1999. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964

(electronic). URL http://link.springer.com/content/pdf/10.1023/
A%3A1005105115011.pdf.

Laraudogoitia:1999:WDS

[771] Jon Pérez Laraudogoitia. Why dynamical self-excitation is possible. Synthese, 119(3):313–323, June 1999. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

content/pdf/10.1023/A%3A1005241129759.pdf.

Bridger:1999:DPL

[772] Mark Bridger and Joseph S. Alper. On the dynamics of Perez Lauraudogoitia’s supertask. Synthese, 119(3):325–337, June 1999. CODEN SYN-TAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.

springer.com/content/pdf/10.1023/A%3A1005222924655.pdf.

Bird:1999:EL

[773] Alexander Bird. Explanation and laws. Synthese, 120(1):1–17, July 1999. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/content/pdf/10.1023/A%
3A1005294018295.pdf.

Callergaard:1999:HER

[774] Robert Callergård. The hypothesis of Ether and Reid’s interpretation of Newton’s first rule of philosophizing. Synthese, 120(1):19–26, July 1999. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/content/pdf/10.1023/A%
3A1005250302365.pdf.

Halonen:1999:UMI

[775] Ilpo Halonen and Jaakko Hintikka. Unification- It’s magnificent but is it explanation? Synthese, 120(1):27–47, July 1999. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.

com/content/pdf/10.1023/A%3A1005202403274.pdf.

Haallsten:1999:DC

[776] Henrik Hållsten. Deductive Chauvinism. Synthese, 120(1):49–59, July 1999. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/content/pdf/10.1023/A%
3A1005206520112.pdf.

Faye:1999:EE

[777] Jan Faye. Explanation explained. Synthese, 120(1):61–75, July 1999. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/content/pdf/10.1023/A%
3A1005258504182.pdf.

Persson:1999:DEM

[778] Johannes Persson. The determinables of explanatory mechanisms. Synthese, 120(1):77–87, July 1999. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

content/pdf/10.1023/A%3A1005210621021.pdf.

Quarfood:1999:ISS

[779] Marcel Quarfood. The individuality of species: Some reflections on the debate. Synthese, 120(1):89–94, July 1999. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.

com/content/pdf/10.1023/A%3A1005262605091.pdf.

Schurz:1999:EU

[780] Gerhard Schurz. Explanation as unification. Synthese, 120(1):95–114, July 1999. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964

(electronic). URL http://link.springer.com/content/pdf/10.1023/
A%3A1005214721929.pdf.

Schweder:1999:CEE

[781] Rebecca Schweder. Causal explanation and explanatory selection. Synthese, 120(1):115–124, July 1999. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

content/pdf/10.1023/A%3A1005266706000.pdf.

Sintonen:1999:WQW

[782] Matti Sintonen. Why questions, and why just why-questions? Synthese, 120(1):125–135, July 1999. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/content/pdf/
10.1023/A%3A1005218822838.pdf.

Urchs:1999:CE

[783] Max Urchs. Complementary explanations. Synthese, 120(1):137–149, July 1999. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964

(electronic). URL http://link.springer.com/content/pdf/10.1023/
A%3A1005270806908.pdf.

Lin:1999:COL

[784] Francis Y. Lin. Chomsky on the ‘ordinary language’ view of language. Synthese, 120(2):151–191, August 1999. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

content/pdf/10.1023/A%3A1005088716396.pdf.

Leiber:1999:LL

[785] Justin Leiber. Language without linguistics. Synthese, 120(2):193–211, August 1999. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964

(electronic). URL http://link.springer.com/content/pdf/10.1023/
A%3A1005211514525.pdf.

McCawley:1999:USO

[786] James D. McCawley. Unconfirmed sightings of an ‘ordinary language’ theory of language. Synthese, 120(2):213–228, August 1999. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://

link.springer.com/content/pdf/10.1023/A%3A1005266531363.pdf.

Moore:1999:PNP

[787] Joseph G. Moore. Propositions, numbers, and the problem of arbitrary identification. Synthese, 120(2):229–263, August 1999. CODEN SYN-TAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.

springer.com/content/pdf/10.1023/A%3A1005231231924.pdf.

Craig:1999:TCT

[788] W. L. Craig. On truth conditions of tensed sentence types. Synthese, 120 (2):265–270, August 1999. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/content/pdf/
10.1023/A%3A1005148323481.pdf.

Cunning:1999:AC

[789] David Cunning. Agency and consciousness. Synthese, 120(2):271–294, August 1999. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964

(electronic). URL http://link.springer.com/content/pdf/10.1023/
A%3A1005192006642.pdf.

Wisniewski:1999:ELE

[790] Andrzej Wiśniewski. Erotetic logic and explanation by abnormic hypotheses. Synthese, 120(3):295–309, September 1999. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.

com/content/pdf/10.1023/A%3A1005221910211.pdf.

Tiel:1999:DKN

[791] Jeffrey R. Tiel. The dogma of Kornblith’s naturalism. Synthese, 120(3): 311–324, September 1999. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/content/pdf/
10.1023/A%3A1005278624115.pdf.

Mcallister:1999:URI

[792] James W. Mcallister. Universal regularities and initial conditions in Newtonian physics. Synthese, 120(3):325–343, September 1999. CODEN SYN-TAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.

springer.com/content/pdf/10.1023/A%3A1005170007328.pdf.

Changizi:1999:VRU

[793] Mark A. Changizi. Vagueness, rationality and undecidability: A theory of why there is vagueness. Synthese, 120(3):345–374, September 1999. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/content/pdf/10.1023/A%
3A1017231303654.pdf.

McIntyre:1999:DSS

[794] Lee McIntyre. Davidson and social scientific laws. Synthese, 120(3):375– 394, September 1999. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/content/pdf/10.
1023/A%3A1005238313845.pdf.

Cheung:1999:PGW

[795] Leo K. C. Cheung. The proofs of the Grundgedanke in Wittgenstein’s Tractatus. Synthese, 120(3):395–410, September 1999. CODEN SYN-TAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.

springer.com/content/pdf/10.1023/A%3A1005209925004.pdf.

Zemac:1999:WTM

[796] Eddy M. Zemac. The world is too much. Synthese, 120(3):411–418, September 1999. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964

(electronic). URL http://link.springer.com/content/pdf/10.1023/
A%3A1005299706715.pdf.

Skyrms:1999:ISI

[797] Brian Skyrms. Introduction to the special issue on statistics and causation. Synthese, 121(1–2):1–2, November 1999. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

content/pdf/10.1023/A%3A1005250012843.pdf.

Cartwright:1999:CDM

[798] Nancy Cartwright. Causal diversity and the Markov condition. Synthese, 121(1–2):3–27, November 1999. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

content/pdf/10.1023/A%3A1005225629681.pdf.

Freedman:1999:TAD

[799] David Freedman and Paul Humphreys. Are there algorithms that discover causal structure? Synthese, 121(1–2):29–54, November 1999. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://

link.springer.com/content/pdf/10.1023/A%3A1005277613752.pdf.

Glymour:1999:RH

[800] Clark Glymour. Rabbit hunting. Synthese, 121(1–2):55–78, November 1999. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/content/pdf/10.1023/A%
3A1005229730590.pdf.

Hausman:1999:LQM

[801] Daniel M. Hausman. Lessons from quantum mechanics. Synthese, 121(1– 2):79–92, November 1999. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/content/pdf/
10.1023/A%3A1005281714660.pdf.

Pearl:1999:PCT

[802] Judea Pearl. Probabilities of causation: Three counterfactual interpretations and their identification. Synthese, 121(1–2):93–149, November 1999. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/content/pdf/10.1023/A%
3A1005233831499.pdf.

Robins:1999:ACM

[803] James M. Robins. Association, causation, and marginal structural models. Synthese, 121(1–2):151–179, November 1999. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.

com/content/pdf/10.1023/A%3A1005285815569.pdf.

Suppes:1999:NDC

[804] Patrick Suppes. The noninvariance of deterministic causal models. Synthese, 121(1–2):181–198, November 1999. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

content/pdf/10.1023/A%3A1005237916478.pdf.

Woodward:1999:CIS

[805] Jim Woodward. Causal interpretation in systems of equations. Synthese, 121(1–2):199–247, November 1999. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

content/pdf/10.1023/A%3A1005242000548.pdf.

Lalor:1999:IQ

[806] Brendan Lalor. Intentionality and qualia. Synthese, 121(3):249–289, December 1999. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964

(electronic). URL http://link.springer.com/content/pdf/10.1023/
A%3A1005209507485.pdf.

Loomis:1999:NPS

[807] Eric Loomis. Necessity, the a priori, and the standard meter. Synthese, 121 (3):291–307, December 1999. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/content/pdf/
10.1023/A%3A1005229326821.pdf.

Wehmeier:1999:CFG

[808] Kai F. Wehmeier. Consistent fragments of grundgesetze and the existence of non-logical objects. Synthese, 121(3):309–328, December 1999. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://

link.springer.com/content/pdf/10.1023/A%3A1005203526185.pdf.

Harcourt:1999:FNT

[809] Edward Harcourt. Frege on ‘i’, ‘now’, ‘today’ and some other linguistic devices. Synthese, 121(3):329–356, December 1999. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.

com/content/pdf/10.1023/A%3A1005161620492.pdf.

Wong:1999:SIK

[810] Wing-Chun Wong. On a semantic interpretation of Kant’s concept of number. Synthese, 121(3):357–383, December 1999. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.

com/content/pdf/10.1023/A%3A1005106218147.pdf.

Copeland:2023:CAT

[811] B. Jack Copeland and Diane Proudfoot. Correction to: On Alan Turing’s anticipation of connectionism. Synthese, 201(2):??, February 2023. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL https://link.springer.com/article/10.1007/s11229-
022-04023-z. See [549].